### v2

# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

Abstract
Recent findings by Cohen et al. (2021) demonstrate that when training neural networks with full-batch gradient descent with step size $\eta$, the largest eigenvalue $\lambda_{\max}$ of the full-batch Hessian consistently stabilizes at $\lambda_{\max}=2/\eta$. These results have significant implications for convergence and generalization. This, however, is not the case of mini-batch stochastic gradient descent (SGD), limiting the broader applicability of its consequences. We show that SGD trains in a different regime we term Edge of Stochastic Stability (EoSS). In this regime, what stabilizes at $2/\eta$ is Batch Sharpness: the expected directional curvature of mini-batch Hessians along their corresponding stochastic gradients. As a consequence, $\lambda_{\max}$–which is generally smaller than Batch Sharpness–is suppressed, aligning with the long-standing empirical observation that smaller batches and larger step sizes favor flatter minima. We further discuss implications for mathematical modeling of SGD trajectories.
Edge of Stability, Optimization for deep learning, SGD
1 Introduction
Training algorithms are a key ingredient to the success of deep learning. Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951), a stochastic variant of Gradient Descent (GD), has been effective in finding parameters that yield good test performance despite the complicated nonlinear nature of neural networks.
In machine learning, full-batch GD and its adaptive versions have been shown to optimize in regime of instability (Xing et al., 2018; Jastrzębski et al., 2020; Cohen et al., 2021, 2022). Precisely, if the step size is $\eta>0$, the highest eigenvalue of the full-batch Hessian–which we will denote $\lambda_{\max}$ for simplicity–grows until $2/\eta$ in a first phase of training called progressive sharpening and hovers right above that value, subject to small oscillations. Cohen et al. (2021) named this regime Edge of Stability (EoS). Surprisingly, unlike for quadratics, where this would imply divergence, this instability does not damage convergence when training neural networks.
The work on EoS is about full-batch methods.
The picture for full-batch algorithms seems (empirically) clear, both for GD (Cohen et al., 2021; Damian et al., 2023) and its adaptive and accelerated version (Cohen et al., 2022, 2024). Unfortunately, neural networks are typically trained with the mini-batch versions of these gradient-based methods, and how Cohen et al. (2021) in Section 6 and Appendices G and H noticed and stated in the introduction and limitations of their work, what they observed is not the case of mini-batch training. Precisely, to quote Cohen et al. (2021):
while the sharpness does not flatline at any value during SGD (as it does during gradient descent), the trajectory of the sharpness is heavily influenced by the step size and batch size (Jastrzębski et al., 2019, 2020), which cannot be explained by existing optimization theory. Indeed, there are indications that the “Edge of Stability” intuition might generalize somehow to SGD, just in a way that does not center around the sharpness. []delimited-[][$\ldots$][ … ] In extending these findings to SGD, the question arises of how to model “stability” of SGD .
We define here a notion of mini-batch sharpness which acclimates to the learning rate and batch size in a way that mirrors $\lambda_{\max}$ in the Edge of Stability.
What was empirically known for mini-batch SGD .
In the case of mini-batch algorithms, (i) Jastrzębski et al. (2020) noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when. Moreover, Cohen et al. (2021) noticed that SGD somehow acclimates to the hyperparameters. However, they did not characterize how, if not negatively, (ii) it seems not to "flatline" at any value of $\lambda_{\max}$ and (iii) if it stabilizes, that always happens at a level they could not quantify which is below the $2/\eta$ threshold, see Figure 1, often without a proper progressive sharpening phase. This scenario leaves the most basic questions open: In what way the location of convergence of SGD acclimates to the choice of hyperparameters? What are the key quantities involved? To be more specific, can we characterize the training phenomena in (i), (ii), (iii) above? What determines them? Does SGD train in an unstable regime?
Edge of Stochastic Stability.
We address the questions above by introducing a general framework for SGD dynamics through the concept of Batch Sharpness: the expectation of the Hessians of mini-batch losses along the direction of the mini-batch gradients, see Section 3.1. This serves as a generalized and direct substitute for $\lambda_{\max}$ in the mini-batch SGD setting. Specifically, we identify a new regime–which we term Edge of Stochastic Stability (EoSS)–where Batch Sharpness stabilizes near $2/\eta$, while maintaining a critical gap with $\lambda_{\max}$. This dynamic explains the plateaus of $\lambda_{\max}$ at lower levels determined by batch size and learning rate, unifying the observed phenomena in mini-batch training.
Organization of the rest of the paper.
Section 2 reviews related work and outlines key open questions. In Section 3, we introduce and empirically validate the phenomenon of Edge of Stochastic Stability. Section 4 discusses what the observed notion of stability entails mathematically. In Section 5, we examine stability mathematically and we study values at which $\lambda_{\max}$ plateaus. In Section 6, we discuss the implications for the structure and modeling of SGD dynamics. Finally, Section 7 summarizes our findings and addresses the limitations of this work.
Figure 1: Mini-batch SGD on CIFAR-10. The full-batch Hessian’s largest eigenvalue $\lambda_{\max}$ typically plateaus below $2/\eta$, here $\eta=1/200$. Smaller batch sizes lead to lower plateau values.
Figure 2: SGD at EoSS under different learning rates and batch sizes. We train a fully-connected network on an 8k subset of CIFAR-10. Batch Sharpness stabilizes at the $2/\eta$ threshold across varying batch sizes and step sizes.
2 Related Work
2.1 Progressive Sharpening and Edge of Stability
Progressive sharpening.
Early work noticed that the local shape of the loss landscape changes rapidly at the beginning of the training (LeCun et al., 2012; Keskar et al., 2016; Achille et al., 2017; Jastrzębski et al., 2018; Fort & Ganguli, 2019). Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training (Keskar et al., 2016; Sagun et al., 2016; Fort & Scherlis, 2019; Jastrzębski et al., 2019). Later, Jastrzębski et al. (2020) and Cohen et al. (2021) managed to make this precise, noticing that, along the trajectories of SGD and GD, $\lambda_{\max}$ increases steadily, often after a small number of steps of decrease. This phenomenon was called progressive sharpening by Cohen et al. (2021).
A phase transition.
Early studies (Goodfellow et al., 2016; Li et al., 2019; Jiang et al., 2019; Lewkowycz et al., 2020) revealed that large initial learning rates often enhance generalization at the cost of initial loss reduction. Jastrzębski et al. (2020) attributed this to a sudden phase transition, termed the break-even point, marking the end of progressive sharpening. This transition slows the convergence by confining the dynamics to a region of lower sharpness. Unlike progressive sharpening, this phenomenon is considered to be induced by the gradient-based algorithm becoming unstable, not by the landscape. Jastrzębski et al. (2020); Cohen et al. (2021, 2022) indeed showed that the phase transition comes at different points for different algorithms on the same landscapes. Jastrzębski et al. (2020) showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it. Cohen et al. (2021, 2022) later showed indeed that it comes at the instability thresholds for the full-batch optimization algorithms.
Full-batch edge of stability.
After the phase transition above, GD and full-batch train in the EoS oscillatory regime (Cohen et al., 2021, 2022), where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value. The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics. Observations from Cohen et al. (2021, 2022) indicate that most training dynamics occur within this regime, effectively determining $\lambda_{\max}$ of the final solution. Lee & Jang (2023) explained why $\lambda_{\max}$ in this regime is generally above $2/\eta$, rarely aligning exactly. The reasons include: (i) nonlinearity of the loss gradient, which shifts the required value depending on higher-order derivatives, and (ii) the EoS being governed by the Hessian along the gradient direction, rather than $\lambda_{\max}$ alone.
EoS and convergence.
There is a growing body of work analyzing the surprising mechanism of EoS in training dynamics with GD. Classically, when gradients depend linearly on parameters, divergence occurs locally if $\eta>\frac{2}{\lambda}$, as demonstrated with one-dimensional parabolas (Cohen et al., 2021). Yet, neural networks often converge even when $\eta\geq\frac{2}{\lambda}$, likely due to the problem’s non-standard geometry. Damian et al. (2023) proposed an explanation under some, empirically tested, assumptions of alignment of third derivatives and gradients. Wang et al. (2022) showed that the product-like structure of shallow networks guarantees convergence with learning rates exceeding the EoS threshold. Beneventano & Woodworth (2025) showed that although convergence is guaranteed, it may be extremely slow, this may, however, benefit implicit regularization effects. For further exploration of convergence and implicit regularizarion in the EoS regime, we direct the reader to Arora et al. (2022); Ahn et al. (2022, 2023); Zhu et al. (2023); Lyu et al. (2023).
2.2 SGD, Hessian, and Generalization
SGD finding flatter minima.
Our result is inherently a result about mini-batch training improving flatness. Specifically, we explain why: {mdframed}
Training with smaller batches constraints the dynamics in areas with smaller eigenvalues for the full-batch Hessian.
This quantifies and characterizes prior observations that SGD tends to locate flat minima (Hochreiter & Schmidhuber, 1994, 1997) and that smaller batch sizes result in reduced Hessian sharpness (Keskar et al., 2016). Analogously, recent work strongly aligns with our result: Jastrzębski et al. (2021) demonstrates that large learning rates in SGD act similarly to penalizing the trace of the Fisher matrix–which approximates the Hessian during training and shares its top eigenspaces (Jastrzębski et al., 2018; Martens, 2020; Thomas et al., 2020)–in image classification tasks.
Sharpness and Generalization.
SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance (LeCun et al., 2012; Keskar et al., 2016; Goyal et al., 2017; Jastrzębski et al., 2018; Masters & Luschi, 2018; Smith et al., 2021). This advantage has been widely attributed to the flatness of the minima (Neyshabur et al., 2017; Wu et al., 2017; Dinh et al., 2017; Kleinberg et al., 2018; Jiang et al., 2019; Xie et al., 2020). Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks (Izmailov et al., 2019; Foret et al., 2021). Moreover, Jastrzębski et al. (2021) shows that, in practice, penalizing the trace of the Fisher information matrix empirically improves generalization and reduces memorization—especially when training with noisy labels. We believe that our quantification and the introduction of a theoretical framework for these improvements in sharpness may be a stepping stone towards understanding these generalization benefits theoretically.
3 SGD Typically Occurs at the EoSS
Extending Cohen et al. (2021), we identify a stable quantity in mini-batch SGD: Batch Sharpness. We empirically demonstrate that SGD operates in a regime we term Edge of Stochastic Stability, observed across multiple datasets and architectures. Batch Sharpness emerges as the key quantity governing this regime, serving as the analogous for mini-batch training of the sharpness.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Comparing different sharpness measures. Red: observed Batch Sharpness on the batch used for the step; Green: average Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$. Top row: MLP (2 hidden layers of width 512); middle: 5-layer CNN; bottom: ResNet-14; all trained on an 8k subset of CIFAR-10.
3.1 Mini-Batch Sharpness
Let $L_{B}(\theta)\!=\!\frac{1}{b}\!\sum_{(x_{i},y_{i})\in B}\!\tilde{L}\bigl{(}f_{\theta}(x_{i}),y_{i}\bigr{)}$ be the empirical loss on a randomly sampled batch $B$ of size $b$. Denote by $\mathcal{H}(L_{B})$ the Hessian w.r.t. $\theta$ of $L_{B}$, and let $\nabla L_{B}(\theta)$ be its gradient. We define:
Definition 1 (Batch Sharpness).
Let $\mathcal{P}_{b}$ be the distribution of random batches of size $b$. The Batch Sharpness is
Batch $\text{Batch Sharpness}\;=\;\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\nabla L_{B}^{\top}\,\mathcal{H}(L_{B})\,\nabla L_{B}\bigr{]}}{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\|\nabla L_{B}\|^{2}\bigr{]}}\,.$
Denote by $\lambda_{\max}:=\lambda_{\max}\left[\mathcal{H}(L)\right]$ the largest eigenvalue of the Hessian of the full-batch loss. While Batch Sharpness behaves as $\lambda_{\max}$ in EoS, the two quantities differ fundamentally. We explore this distinction further in Section 4.
3.2 The Edge of Stochastic Stability
1. Stabilization of Batch Sharpness:
We establish that SGD operates in a stability regime analogous to EoS, which we term Edge of Stochastic Stability. In this regime Batch Sharpness stabilizes at $2/\eta$: {mdframed}
SGD tends to train in an instability regime we call Edge of Stochastic Stability. Precisely, after a phase of progressive sharpening, Batch Sharpness reaches a stability level of $2/\eta$, and hovers there.
2. Stabilization of $\lambda_{\max}$:
Also $\lambda_{\max}$ typically stabilizes. However, it does so at lower values that depends on the trajectory taken (Figures 3 and 5). We discuss the stabilization level of $\lambda_{\max}$ in Section 3.3 and 5. Consequently, standard EoS analysis that relies solely on $\lambda_{\max}$ cannot explain SGD’s actual training behavior in the small-batch setting.
3. Catapults:
Unlike EoS, EoSS stabilizes the expectation of a quantity which the algorithm sees one observation at time. Occasionally, a series of batches sampled have outlying sharpness–that is too high for the stable regime–and steps overshoot, triggering a catapult effect, where Batch Sharpness spikes before rapidly decreasing (Figure 4). This is typically followed by renewed progressive sharpening, returning to the EoSS regime. This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, see, e.g., (Lewkowycz et al., 2020; Zhu et al., 2024).
Figure 4: Catapults during EoSS. During EoSS, Batch Sharpness may through cycles of progressive sharpening and stabilizations. Red dots are the observed sharpnesses of the batches used for the step, the green line is Batch Sharpness.
(a) Increasing step size
(b) Decreasing step size
(c) Increasing batch size
(d) Decreasing batch size
Figure 5: Effects of changing step size or batch size in EoSS. (a) Increasing the learning rate causes a catapult spike before Batch Sharpness settles at the new $2/\eta$. (b) Decreasing the learning rate prompts renewed progressive sharpening. (c) Increasing the batch size narrows the gap between Batch Sharpness and $\lambda_{\max}$. (d) Decreasing the batch size. Triangles mark the steps at which the hyperparameters were changed.
3.3 Batch Sharpness Governs EoSS
Following Cohen et al. (2021) we show here how the dynamics of Batch Sharpness changes when hyper-parameters shift along the dynamics. Interestingly, while $\lambda_{\max}$ and other full-batch Hessian metrics may stabilize, we show here their specific plateaus or trajectories depend on initial conditions, not only on the hyperparameter. By contrast, $\lambda_{\max}(\mathcal{H}(L_{B}))$—the top eigenvalue of the mini-batch Hessian—does consistently stabilizes and shift with hyperparameters.
Changing the step size.
Increasing the step size $\eta$ triggers a catapult spike in all the quantities in considerations and the training loss, before Batch Sharpness re-stabilizes near the updated threshold $2/\eta$ (Fig. 5(a)). This therefore pushes $\lambda_{\max}$ lower. Crucially, if stability was governed solely by $\lambda_{\max}$, this step-size adjustment would have left the dynamics unchanged, as $\lambda_{\max}$ remains below both the original and revised $2/\eta$ thresholds. Conversely, reducing $\eta$ (Fig. 5(b) raises the $2/\eta$ threshold and prompts a new phase of progressive sharpening of Batch Sharpness. Here, $\lambda_{\max}$ also rises, but ultimately stabilizes at a lower value than if the entire training had run with the smaller step size. Again, if stability was governed solely by $\lambda_{\max}$, this step-size adjustment would have had the same effect as starting from scratch with the new step size. Notably, Batch Sharpness initially decreases—a counterintuitive result potentially linked to shifting alignments between gradients and Hessian eigenvectors. This behavior warrants further investigation.
Changing the batch size.
Increasing the batch size reduces the gap between Batch Sharpness and $\lambda_{\max}$, therefore making Batch Sharpness lower,(Fig. 5(c)), and thus below the $2/\eta$ threshold. This causes a new progressive sharpening phase until Batch Sharpness again reaches the $2/\eta$ level, making $\lambda_{\max}$ stabilize at a new, higher level, again lower than the one of the training started from scratch with the new batch size. This shows that it is indeed the Batch Sharpness that is the "limiting" quantity in EoSS. Conversely, decreasing the batch size induces a brief catapult phase, after which Batch Sharpness returns to its previous plateau, as expected. This forces full-batch $\lambda_{\max}$ to adjust to a lower level.
Overall, we find that Batch Sharpness governs EoSS behavior—mirroring how $\lambda_{\max}$ operates in the full-batch EoS—while the actual full-batch $\lambda_{\max}$ lags behind or settles inconsistently, underlining the mini-batch nature of SGD stability.
4 Mathematics of Stability
It may seem surprising at first that Batch Sharpness is not a quantity of the same nature as $\lambda_{\max}$ or is not a function of the full-batch Hessian. The reason is deep into the nature of mini-batch SGD. After recalling the computations of Cohen et al. (2021) in Section 4.1, we show what is the setting of mini-batch SGD and what is the notion of stability which depends on Batch Sharpness in Section 4.2. Later in Section 4.3 we motivate why the stability of mini-batch SGD, surprisingly, can not depend on full-batch quantities.
4.1 A Quadratic Objective: Full-Batch GD
We recall here the arguments of Cohen et al. (2021) and Lee & Jang (2023) for quadratic objective functions to later show how stability differs between the two settings.
Assume we fix a learning rate $\eta$ for the whole training. Consider the example of the quadratic Taylor approximation of a loss function:
$f(x)=\frac{1}{2}x^{\top}\mathbf{A}x+\mathbf{b}^{\top}x+\mathbf{c},$
where $\mathbf{A}$ is the Hessian matrix (symmetric and positive semidefinite), $\mathbf{b}$ is the gradient vector, and $\mathbf{c}$ is a constant. Here, as Cohen et al. (2021) observed, when $\eta\lambda_{\max}(\mathbf{A})>2$ the component of $x_{t}-x^{*}$ parallel to the eigenvector relative to $\lambda_{\max}(\mathbf{A})$ explodes with a power law. This is equivalent to imposing the inequality (1) at every step.
[…]
Indeed, eventually, we have $f(x_{t+1})>f(x_{t})$, i.e., if and only if $\lambda_{\max}\leq 2/\eta$.
4.2 Quadratic Objectives: Mini-Batch GD
Empirically, progressive sharpening halts when the mini-batch analog of (1) reaches zero in expectation. Concretely, EoSS arises if, at each step,
$\eta\,\|\nabla f_{B}(x_{t})\|^{2}\;\sim\;\frac{\eta^{2}}{2}\,\nabla f_{B}(x_{t})^{\top}\mathbf{A}_{B}\nabla f_{B}(x_{t}),$ (2)
preventing the single-batch loss from diverging. This makes sense, as the gradient step is on the mini-batch loss, thus the training is stuck in an oscillatory regime when the single batch loss oscillates (on average) [^1] After completion of this manuscript we found Lee & Jang (2023). We compare this with their work in Appendix C..
One might expect this to imply $\lambda_{\max}(\mathcal{H}(L_{B}))\leq 2/\eta$, as in full-batch (1), but that is typically not the case. What is inducing (1) to rely on $\lambda_{\max}$ is that the Hessians at every step share the eigenvectors relative to the highest eigenvalue. In mini-batch SGD, the sampled Hessians $\mathbf{A}_{B}$ generally have completely different eigenspaces than $\mathbf{A}$ and between each others. Indeed, the case of mini-batch SGD, we are working with the following quadratic Taylor approximation:
$f_{B}(x)=\frac{1}{b}\sum_{i\in B}\left(\frac{1}{2}x^{\top}\mathbf{A}_{i}x+\mathbf{b}_{i}^{\top}x+\mathbf{c}_{i}\right),$ (3)
where $B\subset\{1,2,\ldots,n\}$ is a randomly sampled mini batch of size $b$, and $\mathbf{A}_{i}$, $\mathbf{b}_{i}$, and $\mathbf{c}_{i}$ represent coefficients on the quadratic losses of individual data points. Here the full-batch Hessian $\mathbf{A}$ is given by $\mathbf{A}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i}$. Even if $\lambda_{\max}(\mathbf{A})>2/\eta$ with eigenvector $v_{\max}$, certain $\mathbf{A}_{i}$ may map $v_{\max}$ to 0 or a completely different vector, avoiding the blow-up seen in the full-batch regime.
Enforcing (2) in this setting (with $\mathbf{b}_{i}=0$) thus results in
$\frac{2}{\eta}\ \geq\ \sup_{v\in\mathbb{S}^{d}}\ \frac{v^{\top}\mathbb{E}_{i}\left[(\mathbf{A}_{i})^{3}\right]v}{v^{\top}\mathbb{E}_{i}\left[(\mathbf{A}_{i})^{2}\right]v}\ \begin{subarray}{c}\text{ which does not}\\ \text{simplify to}\end{subarray}\ \lambda_{\max}(\mathbf{A}).$ (4)
4.3 Why Stability cannot Depend on Full-Batch
We show here with an example that in the mini-batch setting the stability thresholds can not depend only on full-batch Hessian or gradients, but it has to depend on the higher moments of them over batch sampling, as we can always construct a counterexample which diverges otherwise.
Imagine on two data points we have $\mathbf{A}_{1}=\alpha\cdot I+M$ and $\mathbf{A}_{2}=\alpha\cdot I-M$, with $\alpha,\gamma>0$ and $M=\begin{pmatrix}0&0\\ 0&\gamma\end{pmatrix}$, and $\mathbf{b}_{1}=\mathbf{b}_{2}=0$. Here $\mathbf{A}$ is $\alpha I$. $\mathbf{A}_{1}$ has eigenvalues $\alpha,\alpha+\gamma$ and $\mathbf{A}_{2}$ has eigenvalues $\alpha,\alpha-\gamma$.
Let us now look at possible full-batch stability results, as developed by Cohen et al. (2021). If the right stability notion for mini-batch SGD depended on the full-batch Hessian or gradients, then it would be independent on $\gamma$, and this can not be the case. For instance, if a notion as the one in Equation (1) would explain stability, then $\alpha$ could grow until $\alpha=2/\eta$. However, no matter the size of $\alpha$, if $\gamma>\max\big{\{}\alpha,\tfrac{2}{\eta}-\alpha\big{\}}$ a mini-batch iteration $x_{t+1}=(1-\eta\mathbf{A}_{i})x_{t}=$ increases the size of the second component of $x$ by a multiplicative factor bigger than 1.
Thus a study of the stability of the system can not depend on the full-batch Hessian but has to depend on how big the oscillations due to the size of $\gamma$ are, i.e., on the higher moments of the distribution of the mini-batch Hessian. Note that this situation would be even more extreme if we had (as in the practice of deep learning) the top eigenvectors of the mini-batch Hessians to point in completely different directions, not just to have high variance. As a sanity check, any stability threshold dependent on the higher moments of the mini-batch Hessian or gradients–as Batch Sharpness or $\mathbb{E}_{B}[\lambda_{\max}(\mathcal{H}(L_{B})]$–would induce $\eta_{\max}$ to depend on $\gamma$ too.
5 The Geometry of the Landscape at EoSS
We devote this section to the discussion of the sizes during training of different quantities which may be or interest of practitioners of theoreticians, as the $\lambda_{\max}$ of full-batch and mini-batch Hessians in terms of learning rate, batch size, and critical batch size. For a batch size $b$, denote by
$\lambda_{\max}^{b}:=\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\lambda_{\max}(\mathcal{H}(L_{B}))\big{]}.$
{mdframed}
Then, we establish here that:
  • Also $\lambda_{\max}^{b}$ stabilizes.
  • $\lambda_{\max}^{b}$ stabilizes at a level that ranges between $4/\eta$ and $2/\eta$. The level is lower for very small and very large batch sizes.
  • $\lambda_{\max}^{b}$ is always greater than $\lambda_{\max}$.
  • We find that the level at which $\lambda_{\max}$ stabilizes is characterized by two different regimes. The threshold is what we call critical batch size. This critical batch size depends on the complexity of the data-model.
  • However, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ typically goes as $1/b^{\alpha}$ generally with $\alpha=0.70$. This is surprising, as when fixing the network, the gap goes as $1/b$ or $1/b^{1/2}$ both empirically and mathematically.
5.1 Average $\lambda_{\max}^{b}$ for the mini-batch Hessians
We establish that $\lambda_{\max}^{b}$–the average over the possible batches sampled of the highest eigenvalue of the mini-batch Hessian–generally stabilizes at a value just bigger than $2/\eta$. This happens on a wide range or models and datasets.
We observe that its stabilization levels is generally very close to $2/\eta$ for very small batch sizes, it increases until the critical batch size, then it decreases again.
Interestingly, we observe that $\lambda_{\max}^{b}$ responds to the change of parameters exactly as Cohen et al. (2021) observed for $\lambda_{\max}$ in the full-batch setting.
5.2 Size of $\lambda_{\max}$
Although we show that the quantity of interest is the mini-batch Hessian, in case of mini-batch SGD, there are both historical and generalization-related reasons to investigate what the spectrum of the full-batch Hessian is.
We establish that also $\lambda_{\max}$ generally stabilizes in the EoSS at a level which is strictly smaller than $2/\eta$. We show that the exact level at which $\lambda_{\max}$ does not only depend on learning rate and batch size but on the full trajectory taken, unlike for EoS (Cohen et al., 2021). Indeed, in Section 3.3 we showed that it does not fully adapt to changes in hyper parameters. However, we observed that when we do not induce any shock in the system by changing hyper parameters during training, we can characterize two regimes for the stabilization levels, see Figure 6. Precisely
  • (i) $b\leq$ critical batch size: $\lambda_{\max}$ grows with $b$ with a power law.
  • (ii) $b\geq$ critical batch size: $\lambda_{\max}$ grows with $b$ much slower while approaching $2/\eta$ from below.
Unfortunately, we could not find a way to relate the size of $\lambda_{\max}$ to learning rate and batch size–see Figure 9–if not through the gap it has with the average $\lambda_{\max}$ of the mini-batch Hessian $\lambda_{\max}^{b}$. Notably, the critical batch size we mention is the point beyond which training enters the full-batch regime. Works as (Zhang et al., 2024) study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation". We conjecture there may be a relation between the two quantities and we leave it to future work.
5.3 The Gap Between $\lambda_{\max}$ and $\lambda_{\max}^{b}$
Note that conceptually $\lambda_{\max}^{b}$ is inherently of the same nature as $\lambda_{\max}$, with the only difference being the "part" of the loss it is computed on-the mini-batch loss for the former, and full-batch loss for the latter. Yet, this difference proves to be crucial in being descriptive of the SGD training dynamics. Indeed, e.g.: {mdframed}
While the mini-batch gradients are an unbiased estimator of the full-batch gradient, $\lambda_{\max}\big{(}\mathcal{H}(L_{B})\big{)}$ is a positively biased estimator of $\lambda_{\max}$.
We indeed show in Lemma 4, that $\lambda_{\max}^{b}$ is strictly bigger than $\lambda_{\max}$ and, as a consequence, decreasing the batch size the gap increases. In particular,
$\lambda_{\max}^{1}\ <\ \lambda_{\max}^{4}\ <\ \lambda_{\max}^{16}\ <\ \lambda_{\max}^{64}\ <\ \lambda_{\max}^{256}\ <\ \ldots$ (5)
Figure 6: Batch Sharpness vs. $\lambda_{\max}^{b}$ vs. $\lambda_{\max}$ at EoSS. The values of Batch Sharpness (green), mini-batch $\lambda_{\max}^{b}$ (orange), and full-batch $\lambda_{\max}$ (blue) against varying batch sizes.
We now empirically show what the gaps $\lambda_{\max}^{b}-\lambda_{\max}$ are, and we discuss why that is surprising. As shown in Appendix G and Figure 9, the log-log plot of the gap against batch size reveals three distinct regimes:
  • Very small $b$: for extremely small batch sizes, the gap plateaus. This is unsurprising, as we are in the (i) regime for $\lambda_{\max}$.
  • Typical/moderate $b$: the gap consistently scales as $1/b^{0.7}$ in the experiment. This trend dominates for a wide range of practical batch sizes, highlighting the systematic scaling behavior of the $\lambda_{\max}^{b}-\lambda_{\max}$ gap. It is important to note that we establish in Appendix G the exponent to be $0.7$ independently from the depth, the seeds, the subsets of Cifar10 considered, and the architecture (MLP, CNN, ResNets). Also note that the critical batch size falls within this range. The limitation of this observation in describing $\lambda_{\max}$ is that this consistent trend is consequence both of $\lambda_{\max}$ growing and $\lambda_{\max}^{b}$ decreasing, not just one of the two.
  • Very big $b$: here generally $\lambda_{\max}^{b}$ is essentially $\lambda_{\max}$ and the gap is converging to zero, we conjecture that some form of scaling limit from random matrix theory applies here. The scaling law is generally of the form: $1/b^{\alpha}$ with $\alpha\geq 1$. The exact value of $\alpha$ may vary with the setting, but the overall trend is robust.
While the behavior for very small batches is not surprising, for bigger batch sizes it is. In computer vision tasks where there are way more parameters than datapoints, we observe that the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ decreases linearly with $1/b$ when evaluating the $\lambda_{\max}^{b}$ of any fixed model with different batch sizes, see Figure 8. This $1/b$ scaling is also what one expects from the analysis in Appendix F.
6 Implications: Why Noise-Injected GD Differs from Mini-Batch SGD
6.1 SGD vs. Noisy Gradient Descent
A common belief is that SGD’s regularization stems from its “noisy” gradients, which find flatter minima. However, our analysis points to the directional structure of mini-batch noise as crucial. To test this, we compare mini-batch SGD (batch size 16) against three noisy GD variants:
  • Directional Noise: Gaussian noise matching the directions and signs of mini-batch SGD noise (Simsekli et al., 2019).
  • Diagonal Noise: Gaussian noise restricted to the diagonal part of the SGD noise covariance.
  • Isotropic Noise: Gaussian noise with isotropic covariance.
As shown in Fig. 7, only directionally accurate noise leads to an EoSS-like regime with $\lambda_{\max}$ stabilizing well below $2/\eta$. More generic (diagonal or isotropic) noise fails to reproduce this behavior, indicating that standard SDE approximations—which typically ignore directional correlations—cannot fully capture the observed dynamics.
Figure 7: SGD vs. Noisy GD. Only noise preserving the covariance structure and direction of SGD leads to $\lambda_{\max}$ plateauing below $2/\eta$ (akin to EoSS). Diagonal or isotropic noise fails to reproduce this behavior. Note that batch sharpness is still plotted for comparison, though for noisy GD it loses its meaning.
From a theoretical perspective (Appendix D), these experiments suggest that stability thresholds differ fundamentally between mini-batch SGD (governed by Batch Sharpness) and noise-injected GD (governed by $\lambda_{\max}$).
6.2 The Mini-Batch Landscape Matters
Classical analyses of neural network optimization often assume a single, static landscape: (i) Online perspective, modeling each sample’s gradient as a noisy unbiased estimator of the expected risk, or (ii) Offline perspective, treating the dataset as fixed and SGD as noisy GD on the empirical loss. In both views, it is the full-batch Hessian that supposedly drives curvature. Our results instead highlight the mini-batch landscape: each update sees a Hessian $\mathcal{H}(L_{B})$ that can differ significantly from $\mathcal{H}(L)$, leading to Batch Sharpness stabilizing at $2/\eta$ even when full-batch $\lambda_{\max}$ is much smaller.
6.3 Challenges for SDE Modeling
Prior works already note limitations of SDE-based approaches for SGD implicit regularization: they may be mathematically ill-posed (Yaida, 2018), fail except under restrictive conditions (Li et al., 2021), converge to qualitatively different minima (HaoChen et al., 2020), or miss higher-order effects (Damian et al., 2021; Li et al., 2022). Recent discrete analyses (Roberts, 2021; Smith et al., 2021; Beneventano, 2023) attempt to address some of these issues. Nonetheless, our findings expose a deeper gap: when batch sizes are small, the geometry of the mini-batch Hessian differs markedly from that of the full-batch, altering both eigenvalues and eigenvector alignments. Conventional SDE models, which assume a static or average Hessian, cannot easily capture these rapid fluctuations. Fully accounting for the distribution of $\{\mathbf{A}_{i}\}$ remains an open challenge, emphasizing the need for further research into truly batch-specific dynamics.
7 Conclusions, Limitations, and Future Work
Conclusions.
We introduced Batch Sharpness as a key descriptor of mini-batch SGD dynamics, showing that it stabilizes near $2/\eta$. This defines an Edge of Stochastic Stability, analogous to the classic EoS (Xing et al., 2018; Cohen et al., 2021, 2022), but now in the small-batch regime most commonly used in practice. Empirically, Batch Sharpness remains at $2/\eta$ while the full-batch $\lambda_{\max}$ hovers below, at a level that systematically depends on batch size; indeed, the gap $\lambda_{\max}^{b}-\lambda_{\max}$ follows a power law $(\text{batch size})^{-0.7}$ across various models and datasets.
Limitations.
(i) We have tested only image-classification tasks, leaving open whether similar phenomena arise in NLP, RL, or other domains. (ii) Although we observe empirical batch-size scaling laws, we lack a full theoretical account, especially at moderate batch sizes. (iii) Our experiments mainly use fixed learning rates and standard architectures, so very large-scale or large-batch settings remain less explored. (iv) We have not analyzed momentum-based or adaptive methods (e.g. Adam), even though full-batch EoS has been seen there (Cohen et al., 2022).
Future Work.
Beyond addressing these limitations, several directions remain: (i) Extending higher-order analyses (Damian et al., 2023; Cohen et al., 2021; Ahn et al., 2022; Lyu et al., 2023) to mini-batch SGD may clarify why $\textit{Batch Sharpness}\sim 2/\eta$, (ii) Refining our batch-size scaling argument via random matrix theory might explain the “typical” exponent (around 0.7) and its possible crossovers for very large or very small batches.
In summary, Batch Sharpness provides a more faithful perspective on SGD than the full-batch $\lambda_{\max}$. Our findings highlight how mini-batch Hessians drive training instabilities and ultimately shape learned solutions, hopefully inspiring further research on how step size and batch size jointly govern deep network training.
Acknowledgement
A special thanks to Stanisław Jastrzębski, Alex Damian, Jeremy Cohen, Afonso Bandeira, Boris Hanin, Mark Lowell, and Renée Carmona for invaluable discussions, which were crucial to the development of this project. We acknowledge the use of DeepSeek, Claude, and ChatGPT for providing code assistance, debugging support, and editing suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A On the Stability
A.1 On Equivalent Formulations of Stability
Consider as working example the quadratic Taylor approximation of a loss function:
$f(x)=\frac{1}{2}x^{\top}\mathbf{A}x+\mathbf{b}^{\top}x+\mathbf{c},$
where $\mathbf{A}$ is the Hessian matrix (symmetric and positive semidefinite), $\mathbf{b}$ is the gradient vector, and $\mathbf{c}$ is a constant.
The question of stability is very nuanced, and we decided to define stability and $\eta_{\max}$ as follows
Definition 2 (Stability).
We say that the optimization algorithm $\mathcal{A}$ on the objective $f$ is stable if it does not diverge.
By Folklore, in the quadratic objective $f$ above, enforcing stability is equivalent to choosing a step size smaller than $\eta_{\max}$ as follows.
Definition 3 (Step size at the edge for an algorithm).
Specify a discrete-time optimization algorithm of choice $\mathcal{A}$. We define "step size at the edge for $\mathcal{A}$" the maximum step size $\eta_{\max}(\mathcal{A})>0$ under which the dynamics do not diverge along the whole trajectory in the quadratic objective case, meaning there does not exist a $t$ such that $|x_{t+1}-x^{*}|>|x_{t}-x^{*}|$. In case of a stochastic update, we define $\eta_{\max}(\mathcal{A})$ analogously, but in expectation.
This is a good definition in the quadratic setting above, no matter how we define divergence, indeed
Lemma 1.
Let $f(x)$ be quadratic, specify a discrete-time optimization algorithm of choice $\mathcal{A}$. Then $\eta_{\max}(\mathcal{A})>0$ is the maximum step size under which the dynamics does not diverge to infinity. Both in terms of $f(x_{t})$ and in terms of $\|x_{t}-x^{*}\|$.
And we have, analogously, stability in terms of non-increasing loss.
Lemma 2.
Let $f(x)$ be quadratic, specify a discrete-time optimization algorithm of choice $\mathcal{A}$. Then $\eta_{\max}(\mathcal{A})>0$ is the maximum step size under which there does not exist a $t$ such that $f(x_{t+1})>f(x_{t})$.
A.2 Specialization of the Setting to Mini-Batch SGD
In the case of mini-batch SGD, we are working, instead, with the following 2-order Taylor approximation:
$f_{B}(x)=\frac{1}{b}\sum_{i\in B}\left(\frac{1}{2}x^{\top}\mathbf{A}_{i}x+\mathbf{b}_{i}^{\top}x+\mathbf{c}_{i}\right),$ (6)
where $B\subset\{1,2,\ldots,n\}$ is a randomly sampled mini batch of size $b$, and $\mathbf{A}_{i}$, $\mathbf{b}_{i}$, and $\mathbf{c}_{i}$ represent contributions from individual data points. The full objective matrices are instead given by:
$\mathbf{A}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i},\quad\mathbf{b}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{b}_{i}.$
Update rule for mini-batch SGD.
Here, the mini-batch SGD update rule is:
$x_{t+1}=x_{t}-\eta\nabla f_{B}(x_{t}),$
where the gradient for mini batch $B$ is:
$\nabla f_{B}(x_{t})\ =\ \frac{1}{b}\sum_{i\in B}\left(\mathbf{A}_{i}x_{t}+\mathbf{b}_{i}\right)\ =\ \left[\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}\right](x_{t}-x^{*})\,+\,\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x^{*}+\mathbf{b}_{i}\right).$ (7)
Substituting this gradient, the update becomes:
$x_{t+1}=x_{t}-\eta\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x_{t}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
To analyze stability, one generally focuses on the evolution of the error $x_{t}-x^{*}$, where $x^{*}=-\mathbf{A}^{-1}\mathbf{b}$ is the optimum. By adding and subtracting $x^{*}$, we obtain that error evolves as:
$x_{t+1}-x^{*}=\left(I-\eta\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}\right)\big{(}x_{t}-x^{*}\big{)}\ -\ \eta\left(\sum_{i\in B}\mathbf{A}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
We can notice here the first difference with the full-batch case:
  1. 1.
    The rest $\sum_{i\in B}\mathbf{A}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}$ it is different than 0 and it is different than 0 even in expectation when sampling is not uniform and independent.
  2. 2.
    The multiplier $\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}$ is not $\mathbf{A}$.
Next note that if we work with an full-batch GD with noise injected with unbiased and independent steps in the fashion of Robbins & Monro (1951), their expectation evens these terms out. We refer the reader to Appendix D for that. If we work with a biased algorithm, as Random Reshuffling (or SGD without replacement) or Shuffle Once, this is not the case (Mishchenko et al., 2020; Smith et al., 2021; Beneventano, 2023). We refer the reader to Appendix E for what changes in this setting.
Furthermore, while in the deterministic case working with $x_{t+1}-x^{*}$ or with $\|x_{t+1}-x^{*}\|^{2}$ is the same, this is not the case for mini-batch algorithms. It is thus non-obvious how to generalize stability to start with. We will see in the next section how to do it.
A.3 Formulating Stability for Mini-Batch SGD
Unlike in the deterministic case, we have to be a little more careful with the formulation here. Let $(\lambda_{i},v_{i})$ be the ordered eigenvalue-eigenvector pairs of $\mathbf{A}$. With a deterministic Hessian matrix $\mathbf{A}$ we have that $x_{t+1}-x^{*}$ does not diverge if and only if the highest eigenvalue $\alpha_{i}$ such that $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ is smaller than $2/\eta$, see (Cohen et al., 2021). The idea is that if a component $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ exists, after $k$ steps it gets multiplied by $(1-\eta\alpha_{i})^{k}$. This implies that the system will diverge if and only if $|1-\eta\alpha_{i}|>1$. When our step is a noise-injected GD similar stability phenomena happens, $\mathbf{A}$ is not impacted by the noise, see Appendix D.
However, in the case of mini-batch SGD we have to be much more precise, and note that at every step, stability does not depend on $\mathbf{A}$, but on the "$\mathbf{A}$" of the mini batch. We have an important changes here:
We do not have persistence of a Hessian matrix "$\mathbf{A}$". In particular, there is no persistence in the eigenspaces of the Hessian from step to step. This implies that we can not make an argument as the one above tracking the components of $x_{t}-x^{*}$. This will have an important implication, that at the step $t$ we care about the direction we are moving towards, not about the eigenvector of the highest eigenvalue. We are now dealing with objects of the form $\nabla f^{\top}\cdot\nabla^{2}f\cdot\nabla f$, not of the form $\lambda_{\max}\nabla^{2}f$.
This imply that we may not really work on the eigenspaces of $\mathbf{A}$ as before, we decide to work with the analogous formulation of stability given by Lemma 2: For quadratic objectives, $|x_{t+1}-x^{*}|\leq|x_{t}-x^{*}|$ reduces in norm if and only if $f(x_{t+1})\leq f(x_{t})$. We will work from now on with the formulation of stability of $f(x_{t+1})\leq f(x_{t})$.
At the step $t$ we can write the following thing update for a quadratic objective $f$.
[…]
which is equivalent to
$2\|\nabla f(x_{t})\|\quad\geq\quad\eta\nabla f(x_{t})^{\top}\cdot\nabla^{2}f(x_{t})\cdot\nabla f(x_{t}).$ (9)
Thus the stability of (Cohen et al., 2021), from this perspective, is equivalent to enforcing for all $t$ the inequality Equation (9), or analogously enforcing it in expectation.
Here, for full-batch GD we could argue that if $\langle x_{t}-x^{*},v_{1}\rangle$ and $\eta\lambda_{1}>2$ but $\eta\lambda_{i}<2$ for all $i\in\{2,3,\ldots,\}$ then there exists a time $T$, logarithmic in the initial values of $x_{0}$, such that $f(x_{T+1})>f(x_{T})$. This time $T$ is defined by the components $2+$ being small enough and the first one big enough that the $|1-\eta\lambda_{1}|>1$ update on the first components outweigh the decrease in the others.
Appendix B Stability for Mini-Batch SGD
In the case of mini-batch SGD we have an update on the mini-batch loss. To get to the loss from the mini-batch loss, we can just take an expectation over the batch-sampling process. Thus, by taking the expectation at every algorithm step we can write the stability requirement as
$\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[f_{B}(x_{t+1})-f_{B}(x_{t})\right]\ =\ -\eta\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\nabla f_{B}(x_{t})\right]\,+\,\frac{\eta^{2}}{2}\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\nabla^{2}f_{B}(x_{t})\nabla f_{B}(x_{t})\right]\ \leq\ 0.$ (10)
This leads to our lemma
Lemma 3.
$\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]$ if and only if the step size $\eta$ is smaller than $2$ divided by
Batch $\textit{Batch Sharpness}\ =\ \frac{\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\cdot\nabla^{2}f_{B}(x_{t})\cdot\nabla f_{B}(x_{t})\right]}{\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\|\nabla f_{B}(x_{t})^{2}\|\right]}.$
Proof.
Note that the Taylor expansion as in Equation 10 ensures that
$\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]\ =\ \mathbb{E}\left[-\eta\nabla f_{B}(x_{t})^{\top}\nabla f_{B}(x_{t})\,+\,\frac{\eta^{2}}{2}\nabla f_{B}(x_{t})^{\top}\cdot\nabla^{2}f_{B}(x_{t})\nabla f_{B}(x_{t})\right]$
Thus, by enforcing $\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]$ is equivalent to enforcing
$2\ \geq\ \eta\cdot\textit{Batch Sharpness}.$
Taking this expectation over the ways in which we can sample the batches we conclude the proof. ∎
Appendix C Comparison with (Lee & Jang, 2023)
After we wrote the first version of this article, we realized that a (Lee & Jang, 2023) already dealt with this problem. We draw here a comparison and show that they come to a different conclusion because they deal with a different notion of stability. While our batch sharpness and their Interaction-Aware Sharpness (IAS) are similar in the experiments–and about the same in case of bigger batch sizes, we discuss how our two notions of stability relate to the one of (Cohen et al., 2021).
They propose that the quantity which stabilizes around $2/\eta$ is, in our notation,
$\mathrm{IAS}\quad:=\quad\frac{\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}^{\top}\cdot\nabla^{2}f\cdot\nabla f_{B}\right]}{\|\nabla f\|^{2}}\quad=\quad\frac{\mathrm{tr}\left(\nabla^{2}f\cdot\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}\nabla f_{B}^{\top}\right]\right)}{\mathrm{tr}\left(\nabla f\nabla f^{\top}\right)}.$ (11)
The very natural way they come up with this quantity is by imposing the following inequality, similar to our Equation (10).
$\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\underbrace{f(x_{t+1})-f(x_{t})}\right]\ =\ -\eta\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla\underbrace{f(x_{t})^{\top}}\nabla f_{B}(x_{t})\right]\,+\,\frac{\eta^{2}}{2}\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\cdot\underbrace{\nabla^{2}f(x_{t})}\nabla f_{B}(x_{t})\right]\ \leq\ 0.$ (12)
The difference between our Equation (10) and this equation above if that the terms in which we put a under brace are here evaluated on the whole dataset and not on the mini batch considered.
While the inequalities seem analogous and both equally motivated at first glance, we claim they carry an important difference. Enforcing Equation (12) does imply that in 1 step the loss does not go up. However, it does not imply that the loss will not go up eventually. The instability threshold discussed in (Cohen et al., 2021), in the following work, and in our is stability about the whole trajectory, not about one precise step of the optimizer in consideration.
[…]
$\begin{split}2/\eta\ \geq\ \mathrm{IAS}\ &=\ \frac{1}{\|\nabla f\|^{2}}\left(\nabla f^{\top}\cdot\nabla^{2}f\cdot\nabla f\mathbb{E}\left[\zeta_{B}^{\top}\cdot\nabla^{2}f\cdot\zeta_{B}\right]\right)\\ &=\ \frac{1}{\|\nabla f\|^{2}}\sum_{i=1}^{d}\lambda_{i}\,\big{(}\,\langle\nabla f,v_{i}\rangle^{2}\,+\,\mathbb{E}[\langle\zeta_{B},v_{i}\rangle^{2}]\,\big{)}.\end{split}$ (13)
However, after $T$ steps we can lower bound the loss with an increasing sequence which eventually diverges. Moreover, the loss asymptotically will approach the sequence with a linear convergence rate. Precisely, define the noise $\zeta_{B}$ such that $\nabla f_{B}=\nabla f+\zeta_{B}$. Then, for instance, if $\nabla f$ of the noise $\zeta_{B}$ have scalar product different than 0 with $v_{d}$, enforcing Equation (12) of (Lee & Jang, 2023) we may have decrease in loss in one step, indeed
$\begin{split}f(x_{T})-f(x_{0})\ &=\ \sum_{t=0}^{T-1}\left(\frac{\eta^{2}}{2}\sum_{i=1}^{d}\lambda_{i}\,\big{(}\,\langle\nabla f(x_{t}),v_{i}\rangle^{2}\,+\,\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{i}\rangle^{2}]\,\big{)}-\|\nabla f(x_{t})\|^{2}\right)\\ &\geq\ \sum_{t=0}^{T-1}\underbrace{\left(\frac{\eta^{2}}{2}\lambda_{1}-\eta\right)\,\langle\nabla f(x_{t}),v_{1}\rangle^{2}}_{positive}\,+\,\underbrace{\frac{\eta^{2}}{2}\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{1}\rangle^{2}]}_{positive}.\end{split}$ (14)
Both the quantities under-braced above–the parallel component to $v_{1}$–explode with the time $t$. This implies, in particular, that imposing $IAS\leq 2/\eta$ may implies one-step stability, but not trajectory stability.
Our quantity, in turn, factors in the variability that the eigenspaces have on different batches. Our quantity is also strictly bigger, thus imposing it to be equal to $2/\eta$ induces a lower sharpness, see, e.g, Appendices G and I.
Appendix D Stability for Noise-Injected Gradients Descents
We will see here that what we showed above is the case of mini-batch algorithms but not of many other versions of stochastic algorithms considered in theoretical work or practice.
We deal here with the case of noisy gradients, which is the case usually analyzed in the practice of theory works, see Section 6. While in minibatch stochastic gradient descent (SGD), the step function is
$x_{t+1}=x_{t}-\eta\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x_{t}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
and the noise comes from mini batching as shown above, in noisy gradient descents we do not care of a possible way to rewrite $A,b,c$ as sum of other terms. Just, there exists a random variable $\zeta_{t}$, independent for different times $t_{1}\neq t_{2}$, such that the $t-th$ step is
$x_{t+1}=x_{t}-\eta\left(\mathbf{A}\mathbf{(x_{t}+\zeta_{t})}+b\right).$
This implies that
$x_{t+1}-x^{*}=\left(I+\eta\mathbf{A}\right)\big{(}x_{t}-x^{*}\big{)}\ +\ A\zeta_{t}.$
This imply that the Hessian does not change from step to step. The expected stability formulation is actually the one of (Lee & Jang, 2023) in the sense of (Cohen et al., 2021) is obtained for $\eta_{\max}=2/\lambda_{\max}(\mathbf{A})$, as discussed above in Appendix C.
This, in particular, establishes mathematically and agrees with what we see empirically in Section 6.
Appendix E Sampling the Mini Batches
For simplicity let us work with the one dimensional case. The point in this section is that In the one dimensional case where $\mathbf{A}_{i}=\mathbf{a}_{i},\mathbf{b}_{i},\mathbf{c}_{i}\in\mathbb{R}$ for all $i\in\{1,2,\ldots,n\}$ we have that
$\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t+1}-x^{*}]=\underbrace{\mathbb{E}\left[\left(1-\eta\frac{1}{b}\sum_{i\in B}\mathbf{a}_{i}\right)\big{(}x_{t}-x^{*}\big{)}\right]}_{(1-\eta\mathbf{a})\cdot\mathbb{E}[x_{t+1}-x^{*}]\text{ when independent steps}}\ -\ \underbrace{\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}\left[\frac{1}{b}\sum_{i\in B}\mathbf{a}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right]}_{0\text{ when uniform sampling}}.$ (15)
When $a_{i},b_{i}$ are sampled uniformly within the dataset and independently of the step $t$, then this becomes
$\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t+1}-x^{*}]=(1-\eta\mathbf{a})\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t}-x^{*}]$ (16)
and the stability threshold is the same as for full-batch GD or any other momentum version, for instance.
In SGD without replacement the "independent steps" are now the epochs, we thus work wlog with an epoch at once exactly in the fashion of (Beneventano, 2023) as the tools previously developed by (Mishchenko et al., 2020; Smith et al., 2021) only work for infinitesimal step size and the tools in Appendix E of (Beneventano, 2023) are the only ones which allow us to consider more products and big learning rate. Let $k=n/b$ be the number of batches in an epoch and call $B_{j}$ the batch of $j-th$ step and $\mathbf{a}_{B_{j}},\mathbf{b}_{B_{j}}$ the averages of $\mathbf{a}$ and $\mathbf{b}$ over the batch $B_{j}$.
$x_{k}-x^{*}=\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\big{(}x_{t}-x^{*}\big{)}\ -\ \sum_{j=1}^{k}\left[\prod_{h=j+1}^{k}\left(1-\eta\mathbf{a}_{B_{h}}\right)\right]\big{(}\mathbf{a}_{B_{j}}x^{*}+\mathbf{b}_{B_{j}}\big{)}.$ (17)
Next note that by taking the expectation of the term $\prod_{j=1}^{k}\left(1+\eta\mathbf{a}_{B_{j}}\right)$ we obtain
$\mathbb{E}\left[\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\right]\ =\ (1-\eta\mathbf{a})^{k}\ +\ \frac{\eta^{2}}{n-1}\binom{k}{2}\left[\frac{1}{n}\sum_{i=1}^{n}a_{i}^{2}-a^{2}\right]\ +\ \ldots$
Here if $\eta=O\left(\frac{b^{2}}{n}[\mathrm{Var}(\mathbf{a}_{1})]^{-1}\right)$ which is generally very small, the second them may already be of size $O(1)$. We already see that in this case $\eta$ of the instability threshold has to be set also through the variance of the $\mathbf{a}_{i}$, the size of the batch or the number of batches. More generally, by applying Proposition ? of (Beneventano, 2023) we obtain that
$\mathbb{E}\left[\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\right]\ =\ (1-\eta\mathbf{a})^{k}\ -\ \kappa_{k}\big{(}1-\eta\mathbf{a}_{i}\big{)}$ (18)
and analogously
$\mathbb{E}\left[\sum_{j=1}^{k}\left[\prod_{h=j+1}^{k}\left(1+\eta\mathbf{a}_{B_{h}}\right)\right]\big{(}\mathbf{a}_{B_{j}}x^{*}+\mathbf{b}_{B_{j}}\big{)}\right]\ =\ -\ \sum_{j=1}^{k}\kappa_{j}\big{(}1+\eta\mathbf{a}_{k},1+\eta\mathbf{a}_{k-1},\ldots,1+\eta\mathbf{a}_{j+1},\mathbf{a}_{j}x^{*}+\mathbf{b}_{j}\big{)}$ (19)
where $\kappa_{k}\big{(}X\big{)}$ is the $k-th$ cumulant of the random variable $X$, which here has an empirical distribution from which we sample the $\mathbf{a}_{i}$s.
This implies that the fact of sampling without replacement already implies different $\eta$ for the edge of stability. We are now ready to deal with the main case.
Appendix F On Largest Eigenvalues of Sums of Matrices
In this section we establish mathematically how the gap between Batch Sharpness and $\lambda_{\max}$ scales with the batch size. Precisely, what size we can expect from the Batch Sharpness-$\lambda_{\max}$ gap for fixed network.
In particular, the following linear algebra results collectively enhance our understanding of the stability and scaling properties of the largest eigenvalues in the context of matrix sums.
F.1 Ordering the Largest Eigenvalues.
The largest singular value of the Hessian matrix derived from single data points is positive. This observation is crucial in establishing the following well-known property of matrix eigenvalues.
Lemma 4.
Let $m,b\in\mathbb{N}$ and consider $m$ matrices $M_{1},M_{2},\ldots,M_{b}\in\mathbb{R}^{m\times m}$ satisfying $\lambda_{\max}>|\lambda_{\min}|$. Then, the largest eigenvalue of their sum satisfies
$\lambda_{\max}\left(\sum_{i=1}^{b}M_{i}\right)\,\ \leq\ \ \sum_{i=1}^{b}\lambda_{\max}\left(M_{i}\right)$ (20)
with equality only if all $M_{i}$ are identical.
This lemma is a direct consequence of the convexity of the operator norm in matrices and the fact that the largest eigenvalue is positive in our setting. In our setting, it implies that with non-identical matrices, the maximum eigenvalue of the sum is strictly less than the sum of the maximum eigenvalues of the individual matrices. To illustrate, consider eigenvalue sequences for batch sizes that are powers of four, though the result generalizes to any $b_{1}<b_{2}$:
$\lambda_{\max}^{1}\ <\ \lambda_{\max}^{4}\ <\ \lambda_{\max}^{16}\ <\ \lambda_{\max}^{64}\ <\ \lambda_{\max}^{256}\ <\ \ldots$ (21)
F.2 Trends of $\lambda_{\max}$ given $b$
[…]
As we discuss in Section 5.3, while the behavior we observe for very small batches is not surprising, for bigger batch sizes it is. In computer vision tasks where there are way more parameters than datapoints, we observe that the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ decreases linearly with $1/b$ when evaluating the $\lambda_{\max}^{b}$ of any fixed model with different batch sizes, see Figure 8. This $1/b$ scaling is also what we expect from our mathematical analysis in Appendix F. We are indeed able to establish that for a fixed net, the gap scales as $1/b$ for small batch sizes, when the worst case $\lambda_{\max}^{b}$ is very different from $\lambda_{\max}$, and with $1/\sqrt{b}$ for big batch sizes [^2] Although big $b$ means such that the number of directions spanned in the parameter space by the vectors $\nabla_{\theta}f(\theta,x)$ are repeated multiple times, and that may be practically unrealistic with the current sizes of networks.. Indeed:
Proposition 1 (Expected Size of the Average of Matrices).
In the notations of Lemma 4. Under the same assumptions as Lemma 5, the spectral norm of the deviation of the average $\left\|\frac{1}{b}\sum_{i}M_{i}-M\right\|$ from its expectation $M$ satisfies:
$\left\|\frac{1}{b}\sum_{i}M_{i}-M\right\|\quad=\quad O_{b}\left(\sqrt{\frac{\sigma^{2}\log m}{b}}+\frac{B\log m}{b}\right)$
where $\sigma^{2}=\frac{1}{b}\max\{\|\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\|,\|\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\|\}$ is the expected second moment of the matrices and $B\geq\|M_{i}-M\|$, for all $i$, is a bound to the biggest random matrix $M_{i}$.
F.3 Random Matrix Theory for Scaling Eigenvalues.
While Section 5 establishes mathematically the order of Batch Sharpness, it lacks of mathematical quantification of their magnitudes. Random matrix theory helps bridging this gap at least for big batch sizes $b$ in the Online SGD case where instead of the full-batch Hessian we take as a reference its theoretical expectation.
Lemma 5 (Matrix Bernstein Inequality).
Let $n_{1},n_{2},b\in\mathbb{N}$, let $M_{1},M_{2},\ldots,M_{b}\in\mathbb{R}^{n_{1}\times n_{2}}$ be independent random matrices satisfying $\mathbb{E}[M_{i}]=M$ and $\|M_{i}-M\|\leq B$ for all $i$, let $v=\max\{\|\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\|,\|\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\|\}$ then for all $t>0$
$\mathbb{P}\left({\textstyle\left\|\sum_{i}M_{i}-M\right\|\geq t}\right)\quad\leq\quad(n_{1}+n_{2})\cdot\exp\bigg{(}-\frac{b^{2}t^{2}/2}{v+Bbt/3}\bigg{)}.$
This lemma provides a probabilistic upper bound on the deviation of the largest eigenvalue as the batch size increases. We now state a proposition that quantifies the expected spectral norm of the average of $b$ matrices $M_{i}$ based on this inequality.
Proof of 1.
The Matrix Bernstein inequality bounds the probability of deviation of $|\sum_{i=1}^{b}(M_{i}-M)|$ by $t$. Rescaling by $1/b$, we see that for the average $\overline{M}_{b}:=\frac{1}{b}\sum_{i=1}^{b}M_{i}$ we have
$\mathbb{P}\left({\textstyle\left\|\overline{M}_{b}-M\right\|\geq t/b}\right)\quad\leq\quad(n_{1}+n_{2})\cdot\exp\bigg{(}-\frac{b^{2}t^{2}/2}{v+Bbt/3}\bigg{)}.$
To bound the expectation $\mathbb{E}[\|\overline{M}_{b}-M\|]$, we use the following general inequality for random variables $X$ with tail bounds:
[…]
For $X=\|\overline{M}_{b}-M\|$, substitute the tail bound:
[…]
We now, introduce a substitution to handle the exponential term. Let:
$z=\frac{b^{2}t^{2}}{v},\quad\text{so that}\quad t=\sqrt{\frac{vz}{b^{2}}}\quad\text{and}\quad dt=\frac{1}{2}\sqrt{\frac{v}{b^{2}z}}\,dz.$
Rewriting the integral in terms of $z$:
[…]
While this integral is complex in its full form, we focus on the dominant terms by examining the asymptotics Large $b$:
  • Variance Contribution: The $v$-term dominates when $z$ is small. This leads to a contribution proportional to:
    $O\left(\frac{\sqrt{v\log(n_{1}+n_{2})}}{b}\right).$
  • Max Norm Contribution: The $B$-term dominates when $z$ is large. This leads to a contribution proportional to:
    $O\left(\frac{B\log(n_{1}+n_{2})}{b}\right).$
Combining these contributions gives:
$\mathbb{E}[\|\overline{M}_{b}-M\|]=O\left(\frac{\sqrt{v\log(n_{1}+n_{2})}}{b}+\frac{B\log(n_{1}+n_{2})}{b}\right).$
Next note that $v=b\cdot\sigma^{2}$. This concludes the proof of Proposition 1.
The proposition indicates that as $b$ increases, the expected deviation of $\overline{M}_{b}$ from $M$ diminishes, with a leading-order term scaling as:
  1. 1.
    Variance Decay: The term $\sqrt{\sigma^{2}/b}$ reflects how the variance contribution decreases as $b$ increases (similar to $1/\sqrt{b}$ scaling for scalar averages).
  2. 2.
    Norm Bound Decay: The term $B/b$ reflects how the worst-case individual matrix norm affects the average.
  3. 3.
    Logarithmic Dimension Dependence: The $\log(n_{1}+n_{2})$ factor accounts for the high-dimensional nature of the problem.
Appendix G Dependence of $\lambda_{\max}^{b}$-$\lambda_{\max}$ Gap on the Batch Size
In this appendix, we further discuss the dependence of the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ from the batch size. As mentioned earlier, there are two distinct cases - the static case, and the "trained" case. In the former, we fix a model at some point of the training, and vary the batch size. In particular, $\lambda_{\max}$ stays constant, and the only variation comes from $\lambda_{\max}^{b}$. In the latter, we fix a batch size at the beginning of the training, and look at the $\lambda_{\max}^{b}$-$\lambda_{\max}$ gap at the end of the training.
Figure 9: Log-log plots of (LHS) the gap between the maximum reached by the $\lambda_{\max}$ and the $\lambda_{\max}$ vs batch size; (RHS) $\lambda_{\max}^{b}$-$\lambda_{\max}$ gap vs batch size.
G.1 The static case
[…]
One can notice that at high batch sizes the observed slope is somewhat bigger ($-1.1$ if fitted to the $\lambda_{\max}^{b}-\lambda_{\max}$ computed for batch sizes larger than 1000). Now, for a dataset of size 8k, a batch size of 1000 constitutes $1/8$ of dataset, and thus has the $\lambda_{\max}^{b}$ very close to the $\lambda_{\max}$. This might potentially reveal a different scaling regime for batch sizes that are closer to dataset size. On the other hand, since the difference between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ becomes increasingly small when batch size approaches dataset size (especially in comparison to the value of each: $\lambda_{\max}^{b}-\lambda_{\max}$ being of order of  1, and each being around 500), the change in scaling might just be an effect of noise in estimating the highest eigenvalue of the hessian. Lastly, there is the effect of finiteness of the dataset size - that is, that the $1/b$ dependence would turn to 0 only when $b$ is infinite, although in reality the gap would be 0 when $b$ is equal to the dataset size. This dependence might effectively break the scaling. Answering the above questions necessitates further investigation. Nonetheless, the $1/b$ dependence appears to persist within the ’realistic’ SGD regime, characterized by batch sizes that are substantially smaller than the dataset size.
The $1/b$ scaling appears to hold throughout the training. In particular, it also applies at initialization, as showcased in the log-log in Figure 11.
Figure 10: Log-log plot of the $\lambda_{\max}^{b}$-$\lambda_{\max}$ for fixed model at convergence.
Figure 11: Log-log plot of the $\lambda_{\max}^{b}$-$\lambda_{\max}$ for a model at initialization
Moreover, the $1/b$ dependence is also architecture-independent. As illustrated in the log-log plot in Figure 12, it is also the case for a CNN architecture at convergence.
Figure 12: Log-log plot of the $\lambda_{\max}^{b}$-$\lambda_{\max}$ for fixed CNN model at convergence.
G.2 The trained case
[…]
Similar to the static case, we again see that the anomalous region at batch sizes that are larger than $1/8$, requiring further investigation. A distinct scaling regime emerges for very small batch sizes ($<10$), differing from the patterns described above. In this regime, the gap appears largely independent of the batch size. This anomaly might arise because, at such small batch sizes, the $\lambda_{\max}^{b}$ starts at levels at or beyond the EoSS level, bypassing the standard progressive sharpening phase and instead entering a regime where the $\lambda_{\max}$ decreases. Further investigation is necessary to rigorously characterize the scaling behavior in this regime.
Figure 13: Log-log plot of $\lambda_{\max}^{b}$-$\lambda_{\max}$ gap vs batch size at the EoSS. Notice how the scaling breaks for very small and very large batch sizes
Figure 14: $\lambda_{\max}^{b}$-$\lambda_{\max}$ gap vs $1/\sqrt{\text{batch\_size}}$ at the EoSS, for batch sizes in $[10,100]$.
[…]
Appendix H The Hessian and the Fisher Information Matrix Overlap
We show here empirically that at EoSS generally $\lambda_{\max}^{b}$ generally overlaps with the largest eigenvalue of the averaged mini-batch NTK and $\frac{1}{b}J_{B}^{\top}J_{B}$, which corresponds with the FIM in vision classification tasks. See Figure 16.
Figure 16: Ranging different batch sizes, the $\lambda_{\max}^{b}$ corresponds to the largest eigenvalue of the averaged mini-batch NTK and $\frac{1}{b}J_{B}^{\top}J_{B}$, which corresponds with the FIM in vision classification tasks.
Appendix I Exemplification Through a Simplified Model
[…]
$f(x)=a^{\top}B\cdot x,\qquad a\in\mathbb{R}^{2},\quad B\in\mathbb{R}^{2\times 2}.$
Then with a diagonal initialization, gradient descent will converge almost surely to a neural network of the following kind
$f(x)=(a_{1},a_{2})\cdot\begin{pmatrix}b_{1}&0\\ 0&b_{2}\end{pmatrix}\cdot x,\qquad\text{where}\quad|a_{1}\cdot b_{1}|\ =\ |a_{2}\cdot b_{2}|\ =\ 1.$
At convergence, the spectrum of the Hessian on the data point $z_{1}$ is $\{\lambda_{1},0,0,0,0,0\}$, with $\lambda_{1}:=a_{1}^{2}+b_{1}^{2}$, the Hessian on the data point $z_{2}$ is instead $\{\lambda_{2},0,0,0,0,0\}$, where $\lambda_{2}:=a_{2}^{2}+b_{2}^{2}$, and the two eigenvectors for these two eigenvalues are orthogonal between each other. This implies that the Hessian of the full-batch loss has spectrum $\{\lambda_{1}/2,\lambda_{2}/2,0,0,0,0\}$, while the Hessian on the mini batches of size one has either one of the spectra above.
This implies that
$\lambda_{\max}\quad=\quad\lambda_{\max}\left(\frac{1}{2}\mathcal{H}(z_{1})+\frac{1}{2}\mathcal{H}(z_{1})\right)\quad=\quad\max\left\{\frac{\lambda_{1}}{2},\ \frac{\lambda_{2}}{2}\right\}$ (22)
This is smaller than the average largest eigenvalue of the mini-batch Hessian which is
$\lambda_{\max}^{1}\quad=\quad\frac{1}{2}\lambda_{\max}(\mathcal{H}(z_{1}))+\frac{1}{2}\lambda_{\max}(\mathcal{H}(z_{2}))\quad=\quad\frac{\lambda_{1}}{2}\ \textbf{+}\ \frac{\lambda_{2}}{2}.$ (23)
  • Smaller size: Thus setting $\lambda_{\max}$ equal to $\lambda$ means that the max between $\lambda_{1}$ and $\lambda_{2}$ is exactly $2\lambda$. Note that the fact that $a_{1}\cdot b_{1}=a_{2}\cdot b_{2}=1$ and Cauchy-Schwartz imply that $\lambda_{1},\lambda_{2}\geq 2$. Setting $\lambda_{\max}^{1}$ to $\lambda$ thus implies that the maximum between $\lambda_{1}$ and $\lambda_{2}$ is at most $22\lambda-$ 2, generally smaller.
  • Higher alignment: Moreover, we have that the gradient $\nabla f(z_{i})$ on the data point $z_{i}$ exactly aligns with the eigenvector $v_{i}$ of the highest eigenvalue $\lambda_{i}$ of the Hessian in $z_{i}$. On the full batch, we are averaging them differently, precisely we have that there exist two constants $c_{1},c_{2}$ such that the gradient is $\frac{c_{1}}{2}v_{1}+\frac{c_{2}}{2}v_{2}$. Thus, where WLOG $\lambda_{1}>\lambda_{2}$ we have the alignments
    $\mathcal{H}(z_{1})\cdot\nabla L(z_{1})\ \sim\ c_{1}\lambda_{1}^{2}v_{1}\qquad\text{but}\qquad\mathcal{H}\cdot\nabla f\ \sim\ \frac{c_{1}}{2}\lambda_{1}^{2}v_{1}$ (24)
    Thus one half of it (batch size divided by number of data points).
This shows that in the same point of the gradient, SGD perceives the largest eigenvalue of the Hessian bigger and more relevant to the gradient then GD.
Appendix J Illustration of EoSS in Variety of Settings: Batch Sharpness
In this appendix, we provide further empirical evidence that EoSS arises robustly across a variety of models, step sizes, and batch sizes. Consistent with our main observations, we find that Batch Sharpness invariably stabilizes around $2/\eta$.
MLP (2-Layer) Baseline.
Figure 20 illustrates EoSS for our baseline network, an MLP with two hidden layers of dimension 512, trained on an $8192$-sample subset of CIFAR-10 with step size $\eta=0.004$. As the training proceeds, Batch Sharpness stabilizes around $2/\eta$, whereas $\lambda_{\max}$ plateaus strictly below Batch Sharpness. Decreasing the step size to $\eta=0.002$ (see Figure 17) rescales the plateau of Batch Sharpness around the new threshold $2/\eta$, in line with the behavior discussed in the main text.
5-Layer CNN.
We further confirm the EoSS regime in a five-layer CNN. As depicted in Figures 24 and 18, Batch Sharpness continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level. Notably, as we vary the batch size, the gap between Batch Sharpness and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 5.
ResNet-14.
Finally, we demonstrate that the EoSS regime also emerges for a canonical architecture commonly used in computer vision: ResNet-14. Note that we are using a version without Batc hNormalization. Figure 19 highlights the same qualitative behavior, with Batch Sharpness stabilizing at $2/\eta$.
Overall, these experiments provide further confirmation that EoSS is a robust phenomenon across different architectures, step sizes, and batch sizes.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 17: MLP: 2 hidden layers, hidden dimension 512; learning rate 0.01, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 18: CNN: 5 layers (3 convolutional, 2 fully-connected), learning rate 0.03, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 19: ResNet-10, learning rate 0.005, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
Appendix K Illustration of EoSS in Variety of Settings: $\lambda_{\max}^{b}$
[…]
MLP (2-Layer) Baseline.
[…]
Deeper MLP (4-Layer).
To assess whether increased depth alters the phenomenon, we use a deeper MLP (MLP_L) with four hidden layers, training again on the same CIFAR-10 subset. Figures 22 and 23 show that $\lambda_{\max}^{b}$ exhibits the same EoSS behavior for two different step sizes, reinforcing that depth alone does not invalidate our findings.
5-Layer CNN.
We further confirm the EoSS regime in a five-layer CNN. As depicted in Figures 24 and 25, $\lambda_{\max}^{b}$ continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level. Notably, as we vary the batch size, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 5.
ResNet-10.
[…]
Overall, these experiments provide further confirmation that EoSS is a robust phenomenon across different architectures, step sizes, and batch sizes. Although the specific magnitude of $\lambda_{\max}$ and the exact “hovering” value of $\lambda_{\max}^{b}$ can vary, the overarching pattern of $\lambda_{\max}^{b}\approx 2/\eta$ and $\lambda_{\max}<\lambda_{\max}^{b}$ persists in all our tested settings.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 20: MLP, 2 hidden layers, hidden dimension 512, learning rate 0.004, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical $\lambda_{\max}^{b}$ (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 21: MLP: 2 hidden layers, hidden dimension 512; learning rate 0.01, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical $\lambda_{\max}^{b}$ (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 22: MLP_L: 4 hidden layers, hidden dimension 512, learning rate 0.002, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical $\lambda_{\max}^{b}$ (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 23: MLP_L, 4 hidden layers, hidden dimension 512, learning rate 0.004, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical $\lambda_{\max}^{b}$ (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 24: CNN, 5 layers (3 convolutional, 2 fully-connected), learning rate 0.02, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical $\lambda_{\max}^{b}$ (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 25: CNN, 5 layers (3 convolutional, 2 fully-connected), learning rate 0.005, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical $\lambda_{\max}^{b}$ (green line), the $\lambda_{\max}$ (blue dots).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 26: ResNet-10, learning rate 0.005, 8k subset of CIFAR-10. Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical $\lambda_{\max}^{b}$ (green line), the $\lambda_{\max}$ (blue dots).

### v3

@@
 Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951), a stochastic variant of Gradient Descent (GD), has been effective in finding parameters that yield good test performance despite the complicated nonlinear nature of neural networks.
-In machine learning, full-batch GD and its adaptive versions have been shown to optimize in regime of instability (Xing et al., 2018; Jastrzębski et al., 2020; Cohen et al., 2021, 2022).
-Precisely, if the step size is $\eta>0$, the highest eigenvalue of the full-batch Hessian–which we will denote $\lambda_{\max}$ for simplicity–grows until $2/\eta$ in a first phase of training called progressive sharpening and hovers right above that value, subject to small oscillations.
+In machine learning, full-batch GD and its adaptive versions have been shown to optimize in regime of instability (Xing et al., 2018; Jastrzębski et al., 2019, 2020; Cohen et al., 2021, 2022).
+Precisely, if the step size is $\eta>0$ for full-batch GD, the highest eigenvalue of the full-batch Hessian–which we will denote $\lambda_{\max}$ for simplicity–grows until $2/\eta$ in a first phase of training called progressive sharpening and hovers right above that value, subject to small oscillations.
 Cohen et al.
@@
 In the case of mini-batch algorithms, (i) Jastrzębski et al.
-(2020) noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when.
+(2019, 2020) noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when.
 Moreover, Cohen et al.
@@
 Progressive sharpening.
-Early work noticed that the local shape of the loss landscape changes rapidly at the beginning of the training (LeCun et al., 2012; Keskar et al., 2016; Achille et al., 2017; Jastrzębski et al., 2018; Fort & Ganguli, 2019).
+Early work noticed that the local shape of the loss landscape changes rapidly at the beginning of the training (LeCun et al., 2012; Keskar et al., 2016; Achille et al., 2017; Jastrzębski et al., 2018, 2019; Fort & Ganguli, 2019).
 Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training (Keskar et al., 2016; Sagun et al., 2016; Fort & Scherlis, 2019; Jastrzębski et al., 2019).
 Later, Jastrzębski et al.
-(2020) and Cohen et al.
+(2019, 2020) and Cohen et al.
 (2021) managed to make this precise, noticing that, along the trajectories of SGD and GD, $\lambda_{\max}$ increases steadily, often after a small number of steps of decrease.
@@
 Jastrzębski et al.
-(2020); Cohen et al.
+(2019, 2020); Cohen et al.
 (2021, 2022) indeed showed that the phase transition comes at different points for different algorithms on the same landscapes.
 Jastrzębski et al.
-(2020) showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it.
+(2019, 2020) showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it.
 Cohen et al.
@@
 Full-batch edge of stability.
-After the phase transition above, GD and full-batch train in the EoS oscillatory regime (Cohen et al., 2021, 2022), where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value.
+After the phase transition above, GD and full-batch Adam train in the EoS oscillatory regime (Cohen et al., 2021, 2022), where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value.
 The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics.
@@
 Training with smaller batches constraints the dynamics in areas with smaller eigenvalues for the full-batch Hessian.
-This quantifies and characterizes prior observations that SGD tends to locate flat minima (Hochreiter & Schmidhuber, 1994, 1997) and that smaller batch sizes result in reduced Hessian sharpness (Keskar et al., 2016).
+This quantifies and characterizes prior observations that SGD tends to locate flat minima (Hochreiter & Schmidhuber, 1997) and that smaller batch sizes result in reduced Hessian sharpness (Keskar et al., 2016).
 Analogously, recent work strongly aligns with our result: Jastrzębski et al.
@@
 SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance (LeCun et al., 2012; Keskar et al., 2016; Goyal et al., 2017; Jastrzębski et al., 2018; Masters & Luschi, 2018; Smith et al., 2021).
-This advantage has been widely attributed to the flatness of the minima (Neyshabur et al., 2017; Wu et al., 2017; Dinh et al., 2017; Kleinberg et al., 2018; Jiang et al., 2019; Xie et al., 2020).
+This advantage has been widely attributed to the flatness of the minima (Murata et al., 1994; Hochreiter & Schmidhuber, 1994; Neyshabur et al., 2017; Wu et al., 2017; Dinh et al., 2017; Kleinberg et al., 2018; Jiang et al., 2019; Xie et al., 2020).
 Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks (Izmailov et al., 2019; Foret et al., 2021).
@@
 Figure 3: Comparing different sharpness measures.
-Red: observed Batch Sharpness on the batch used for the step; Green: average Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$.
+Red: observed sharpness on the step’s mini batch on the batch used for the step; Green: average Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$.
 Top row: MLP (2 hidden layers of width 512); middle: 5-layer CNN; bottom: ResNet-14; all trained on an 8k subset of CIFAR-10.
@@
 This is typically followed by renewed progressive sharpening, returning to the EoSS regime.
-This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, see, e.g., (Lewkowycz et al., 2020; Zhu et al., 2024).
+This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, see, e.g., Lewkowycz et al.
+(2020); Zhu et al.
+(2024).
 Figure 4: Catapults during EoSS.
@@
 Notably, the critical batch size we mention is the point beyond which training enters the full-batch regime.
-Works as (Zhang et al., 2024) study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation".
+Works as Zhang et al.
+(2024) study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation".
 We conjecture there may be a relation between the two quantities and we leave it to future work.
@@
 Let $(\lambda_{i},v_{i})$ be the ordered eigenvalue-eigenvector pairs of $\mathbf{A}$.
-With a deterministic Hessian matrix $\mathbf{A}$ we have that $x_{t+1}-x^{*}$ does not diverge if and only if the highest eigenvalue $\alpha_{i}$ such that $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ is smaller than $2/\eta$, see (Cohen et al., 2021).
+With a deterministic Hessian matrix $\mathbf{A}$ we have that $x_{t+1}-x^{*}$ does not diverge if and only if the highest eigenvalue $\alpha_{i}$ such that $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ is smaller than $2/\eta$, see Cohen et al.
+(2021).
 The idea is that if a component $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ exists, after $k$ steps it gets multiplied by $(1-\eta\alpha_{i})^{k}$.
@@
 $2\|\nabla f(x_{t})\|\quad\geq\quad\eta\nabla f(x_{t})^{\top}\cdot\nabla^{2}f(x_{t})\cdot\nabla f(x_{t}).$ (9)
-Thus the stability of (Cohen et al., 2021), from this perspective, is equivalent to enforcing for all $t$ the inequality Equation (9), or analogously enforcing it in expectation.
+Thus the stability of Cohen et al. (2021), from this perspective, is equivalent to enforcing for all $t$ the inequality Equation (9), or analogously enforcing it in expectation.
 Here, for full-batch GD we could argue that if $\langle x_{t}-x^{*},v_{1}\rangle$ and $\eta\lambda_{1}>2$ but $\eta\lambda_{i}<2$ for all $i\in\{2,3,\ldots,\}$ then there exists a time $T$, logarithmic in the initial values of $x_{0}$, such that $f(x_{T+1})>f(x_{T})$.
@@
 Taking this expectation over the ways in which we can sample the batches we conclude the proof. ∎
-Appendix C Comparison with (Lee & Jang, 2023)
-After we wrote the first version of this article, we realized that a (Lee & Jang, 2023) already dealt with this problem.
+Appendix C Comparison with Lee & Jang (2023)
+After we wrote the first version of this article, we realized that Lee & Jang (2023) already dealt with this problem.
 We draw here a comparison and show that they come to a different conclusion because they deal with a different notion of stability.
-While our batch sharpness and their Interaction-Aware Sharpness (IAS) are similar in the experiments–and about the same in case of bigger batch sizes, we discuss how our two notions of stability relate to the one of (Cohen et al., 2021).
+While our batch sharpness and their Interaction-Aware Sharpness (IAS) are similar in the experiments–and about the same in case of bigger batch sizes, we discuss how our two notions of stability relate to the one of Cohen et al.
+(2021).
 They propose that the quantity which stabilizes around $2/\eta$ is, in our notation,
@@
 However, it does not imply that the loss will not go up eventually.
-The instability threshold discussed in (Cohen et al., 2021), in the following work, and in our is stability about the whole trajectory, not about one precise step of the optimizer in consideration.
+The instability threshold discussed in Cohen et al.
+(2021), in the following work, and in our is stability about the whole trajectory, not about one precise step of the optimizer in consideration.
 […]
@@
 Precisely, define the noise $\zeta_{B}$ such that $\nabla f_{B}=\nabla f+\zeta_{B}$.
-Then, for instance, if $\nabla f$ of the noise $\zeta_{B}$ have scalar product different than 0 with $v_{d}$, enforcing Equation (12) of (Lee & Jang, 2023) we may have decrease in loss in one step, indeed
+Then, for instance, if $\nabla f$ of the noise $\zeta_{B}$ have scalar product different than 0 with $v_{d}$, enforcing Equation (12) of Lee & Jang (2023) we may have decrease in loss in one step, indeed
 $\begin{split}f(x_{T})-f(x_{0})\ &=\ \sum_{t=0}^{T-1}\left(\frac{\eta^{2}}{2}\sum_{i=1}^{d}\lambda_{i}\,\big{(}\,\langle\nabla f(x_{t}),v_{i}\rangle^{2}\,+\,\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{i}\rangle^{2}]\,\big{)}-\|\nabla f(x_{t})\|^{2}\right)\\ &\geq\ \sum_{t=0}^{T-1}\underbrace{\left(\frac{\eta^{2}}{2}\lambda_{1}-\eta\right)\,\langle\nabla f(x_{t}),v_{1}\rangle^{2}}_{positive}\,+\,\underbrace{\frac{\eta^{2}}{2}\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{1}\rangle^{2}]}_{positive}.\end{split}$ (14)
@@
 This imply that the Hessian does not change from step to step.
-The expected stability formulation is actually the one of (Lee & Jang, 2023) in the sense of (Cohen et al., 2021) is obtained for $\eta_{\max}=2/\lambda_{\max}(\mathbf{A})$, as discussed above in Appendix C.
+The expected stability formulation is actually the one of Lee & Jang (2023) in the sense of Cohen et al.
+(2021) is obtained for $\eta_{\max}=2/\lambda_{\max}(\mathbf{A})$, as discussed above in Appendix C.
 This, in particular, establishes mathematically and agrees with what we see empirically in Section 6.
@@
 and the stability threshold is the same as for full-batch GD or any other momentum version, for instance.
-In SGD without replacement the "independent steps" are now the epochs, we thus work wlog with an epoch at once exactly in the fashion of (Beneventano, 2023) as the tools previously developed by (Mishchenko et al., 2020; Smith et al., 2021) only work for infinitesimal step size and the tools in Appendix E of (Beneventano, 2023) are the only ones which allow us to consider more products and big learning rate.
+In SGD without replacement the "independent steps" are now the epochs, we thus work wlog with an epoch at once exactly in the fashion of Beneventano (2023) as the tools previously developed by Mishchenko et al.
+(2020); Smith et al.
+(2021) only work for infinitesimal step size and the tools in Appendix E of Beneventano (2023) are the only ones which allow us to consider more products and big learning rate.
 Let $k=n/b$ be the number of batches in an epoch and call $B_{j}$ the batch of $j-th$ step and $\mathbf{a}_{B_{j}},\mathbf{b}_{B_{j}}$ the averages of $\mathbf{a}$ and $\mathbf{b}$ over the batch $B_{j}$.
@@
 We already see that in this case $\eta$ of the instability threshold has to be set also through the variance of the $\mathbf{a}_{i}$, the size of the batch or the number of batches.
-More generally, by applying Proposition ?
-of (Beneventano, 2023) we obtain that
+More generally, by applying the results in Appendix E of Beneventano (2023) we obtain that
 $\mathbb{E}\left[\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\right]\ =\ (1-\eta\mathbf{a})^{k}\ -\ \kappa_{k}\big{(}1-\eta\mathbf{a}_{i}\big{)}$ (18)

### v4

@@
 Recent findings by Cohen et al.
-(2021) demonstrate that when training neural networks with full-batch gradient descent with step size $\eta$, the largest eigenvalue $\lambda_{\max}$ of the full-batch Hessian consistently stabilizes at $\lambda_{\max}=2/\eta$.
+[1] demonstrate that when training neural networks with full-batch gradient descent with step size $\eta$, the largest eigenvalue $\lambda_{\max}$ of the full-batch Hessian consistently stabilizes at $\lambda_{\max}=2/\eta$.
 These results have significant implications for convergence and generalization.
@@
 In this regime, what stabilizes at $2/\eta$ is Batch Sharpness: the expected directional curvature of mini-batch Hessians along their corresponding stochastic gradients.
-As a consequence, $\lambda_{\max}$–which is generally smaller than Batch Sharpness–is suppressed, aligning with the long-standing empirical observation that smaller batches and larger step sizes favor flatter minima.
+As a consequence, $\lambda_{\max}$—which is generally smaller than Batch Sharpness—is suppressed, aligning with the long-standing empirical observation that smaller batches and larger step sizes favor flatter minima.
 We further discuss implications for mathematical modeling of SGD trajectories.
-Edge of Stability, Optimization for deep learning, SGD
 1 Introduction
-Training algorithms are a key ingredient to the success of deep learning.
-Stochastic Gradient Descent (SGD) (Robbins & Monro, 1951), a stochastic variant of Gradient Descent (GD), has been effective in finding parameters that yield good test performance despite the complicated nonlinear nature of neural networks.
-In machine learning, full-batch GD and its adaptive versions have been shown to optimize in regime of instability (Xing et al., 2018; Jastrzębski et al., 2019, 2020; Cohen et al., 2021, 2022).
-Precisely, if the step size is $\eta>0$ for full-batch GD, the highest eigenvalue of the full-batch Hessian–which we will denote $\lambda_{\max}$ for simplicity–grows until $2/\eta$ in a first phase of training called progressive sharpening and hovers right above that value, subject to small oscillations.
+The choice of training algorithm is a key ingredient in the deep learning recipe.
+Extensive evidence, e.g.
+[2], indeed shows that performance consistently depends on the optimizer and hyper-parameters.
+How and why this is the case is a key question in optimization theory and deep learning.
+Gradient Descent (GD) was shown to optimize neural networks in a regime of instability, termed Edge of Stability (EoS) by Cohen et al.
+[1].
+With a constant step size $\eta$, the highest eigenvalue of the Hessian of the full-batch loss—denoted here as $\lambda_{\max}$—grows until $2/\eta$ and hovers right above that value, subject to small oscillations [1, 3, 4, 5, 6].
+Although classical convex optimization theory would call this step size “too large”, the loss continues to decrease.
+These works established a number of surprising facts: (1) that we require an optimization theory which works in more general scenarios then the classical $\eta<2/L$; (2) that GD for neural networks may not stabilize at stationary points but when $\lambda_{\max}=2/\eta$; (3) what a source of instability of (pre-)training is: progressive sharpening; (4) how location of stabilization depends on the choice of hyperparameters of full-batch optimizer.
+While real-world training is almost always mini-batch—given the large amount of data—existing EoS analyses explicitly do not apply to this case: no curvature-type quantities, as $\lambda_{\max}$, are known to similarly affect SGD while training neural networks.
+We bridge this gap by establishing that mini-batch SGD trains in a similar regime of instability which we term Edge of Stochastic Stability (EoSS).
+Precisely, a quantity we call Batch Sharpness, see Definition 1 and Figure 1, hovers around $2/\eta$ and implicitly functions as sharpness in the case of mini-batch algorithms.
+Organization and Contributions.
+Section 2 reviews related work and outlines the key open questions we tackle, namely how does the Edge of Stability phenomenon generalize to mini-batch algorithms and how can we quantify the characters in play.
+As oscillatory behaviors are central to these phenomena, Section 3 disentangles the two types of SGD oscillations into noise-driven (as in Robbins-Monro–type of stochastic optimization when the step size is kept fixed) and curvature-driven—which are the ones we are interested in.
+We establish here that most training happens under oscillations, only the second part of those are curvature-driven.
+In Section 4, we introduce, properly characterize, and empirically validate the phenomenon of Edge of Stochastic Stability.
+In Section 5 we discuss the theoretical motivation behind the quantity we term as Batch Sharpness.
+Batch Sharpness is the implicit notion of curvature that governs the phenomenon of Edge of Stability in the case of mini-batch algorithms, see Definition 1.
+It is the expected curvature along the step on the mini-batch landscape.
+We also prove that when Batch Sharpness bigger than $2/\eta$, the oscillations are unstable.
+A good feature of full-batch–EoS phenomenon was that the quantity of interest (the largest eigenvalue of the Hessian) has a key geometrical role which has been linked both to generalization and to speed of optimization.
+For Batch Sharpness this link is less direct, we thus investigate in Section 6 the relationship between these two quantities.
+Our results are another proof of the fact that the dynamics of noise-injected GD or SDEs and the dynamics of mini-batch SGD are qualitatively different and studying the firsts could be misleading for inducing properties of the second.
+We discuss this implication in Section 7, before concluding in Section 8.
+Figure 1: SGD at EoSS under different learning rates and batch sizes.
+MLP on an 8k subset of CIFAR-10 with step size $\eta>0$.
+Batch Sharpness stabilizes at the $2/\eta$ threshold across varying batch sizes and step sizes.
+2 Related Work
+Progressive sharpening.
+Early studies observed that the local shape of the loss landscape changes rapidly at the beginning of the training [2, 4, 7, 8, 9, 10].
+Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training [2, 4, 11, 12].
+Subsequently, Jastrzębski et al.
+[4, 5] and Cohen et al.
+[1] precisely characterized this behavior, demonstrating a steady rise in $\lambda_{\max}$ along GD and SGD trajectories, typically following a brief initial decline.
+This phenomenon was termed progressive sharpening by Cohen et al.
+[1].
+A phase transition.
+Early works [13, 14, 15, 16] revealed that large initial learning rates often enhance generalization at the cost of initial loss reduction.
+Jastrzębski et al.
+[5] attributed this to a sudden phase transition, termed the break-even point, marking the end of progressive sharpening.
+This transition slows the convergence by confining the dynamics to a region of lower sharpness.
+Unlike progressive sharpening, this phenomenon is considered to be induced by the gradient-based algorithm becoming unstable, not by the landscape.
+Jastrzębski et al.
+[4, 5], Cohen et al.
+[1, 3] indeed showed that the phase transition comes at different points for different algorithms on the same landscapes.
+Jastrzębski et al.
+[4, 5] showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it.
 Cohen et al.
-(2021) named this regime Edge of Stability (EoS).
-Surprisingly, unlike for quadratics, where this would imply divergence, this instability does not damage convergence when training neural networks.
+[1, 3] later showed indeed that it comes at the instability thresholds for the full-batch optimization algorithms.
+We manage here to quantify the value of the instability threshold for mini-batch SGD, thus characterizing when the phase transition happen for SGD.
+Importantly, the fact that progressive sharpening does not end on the landscape, but is limited by algorithmic properties implies that the locations of stabilization of the optimization algorithms surprisingly are not stationary points.
+Full-batch edge of stability.
+After the phase transition above, GD and full-batch Adam train in the EoS oscillatory regime [1, 3, 6], where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value.
+The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics.
+Observations from Cohen et al.
+[1, 3] indicate that, under mean square error (MSE), most training dynamics occur within this regime, effectively determining $\lambda_{\max}$ of the final solution.
+Lee and Jang [17] explained why in this regime $\lambda_{\max}$ often slightly exceeds $2/\eta$: this deviation arises primarily from nonlinearity of the loss gradient, which shifts the required value depending on higher-order derivatives, and the EoS being governed by the Hessian along the gradient direction, rather than $\lambda_{\max}$ alone.
+EoS and convergence.
+A growing body of research analyzes the surprising mechanism underlying EoS dynamics observed during training with GD.
+Classically, when gradients depend linearly on parameters, divergence occurs locally if $\eta>\frac{2}{\lambda_{\max}}$, as illustrated by one-dimensional quadratic models [1].
+In contrast, neural networks often converge despite violating this classical stability condition, presumably due to the problem’s non-standard geometry.
+Damian et al.
+[18] propose an explanation under some, empirically tested, assumptions of alignment of third derivatives and gradients.
+For further insights into convergence and implicit regularization phenomena in the EoS regime, we direct the reader to [19, 20, 21, 22, 23, 24, 25, 26].
 The work on EoS is about full-batch methods.
-The picture for full-batch algorithms seems (empirically) clear, both for GD (Cohen et al., 2021; Damian et al., 2023) and its adaptive and accelerated version (Cohen et al., 2022, 2024).
-Unfortunately, neural networks are typically trained with the mini-batch versions of these gradient-based methods, and how Cohen et al.
-(2021) in Section 6 and Appendices G and H noticed and stated in the introduction and limitations of their work, what they observed is not the case of mini-batch training.
-Precisely, to quote Cohen et al.
-(2021):
-while the sharpness does not flatline at any value during SGD (as it does during gradient descent), the trajectory of the sharpness is heavily influenced by the step size and batch size (Jastrzębski et al., 2019, 2020), which cannot be explained by existing optimization theory.
-Indeed, there are indications that the “Edge of Stability” intuition might generalize somehow to SGD, just in a way that does not center around the sharpness.
-[]delimited-[][$\ldots$][ … ] In extending these findings to SGD, the question arises of how to model “stability” of SGD .
-We define here a notion of mini-batch sharpness which acclimates to the learning rate and batch size in a way that mirrors $\lambda_{\max}$ in the Edge of Stability.
-What was empirically known for mini-batch SGD .
-In the case of mini-batch algorithms, (i) Jastrzębski et al.
-(2019, 2020) noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when.
+While the empirical behavior of EoS for full-batch algorithms is relatively well-understood, neural networks are predominantly trained using mini-batch methods.
+As explicitly noted by Cohen et al.
+[1, Section 6, Appendices G and H], their observations and analysis do not directly apply to mini-batch training.
+In particular, they emphasize:
+[…] while the sharpness does not flatline at any value during SGD (as it does during gradient descent), the trajectory of the sharpness is heavily influenced by the step size and batch size Jastrzębski et al.
+[4, 5], which cannot be explained by existing optimization theory.
+Indeed, there are indications that the “Edge of Stability” intuition might generalize somehow to SGD, just in a way that does not center around the (full-batch) sharpness.
+[…] In extending these findings to SGD, the question arises of how to model “stability” of SGD.
+In this work, we show that the EoS phenomenon does indeed generalize to SGD, and we identify the central quantity governing this generalization (Batch Sharpness in Definition 1).
+We modeled stability of SGD on the neural networks landscapes: our answer is that empirically SGD is stable if on average the step is stable on the mini-batch landscape—not on the full-batch landscape.
+What was empirically known for SGD.
+Figure 2: SGD on CIFAR-10: $\eta=1/200$. The full-batch Hessian’s $\lambda_{\max}$ plateaus below $2/\eta$. Smaller batch sizes lead to lower plateau values.
+In the context of mini-batch algorithms, (i) Jastrzębski et al.
+[4, 5] noticed that for SGD the phase transition happens earlier for smaller $\eta$ or smaller batch size $b$, but they did not quantify when.
 Moreover, Cohen et al.
-(2021) noticed that SGD somehow acclimates to the hyperparameters.
-However, they did not characterize how, if not negatively, (ii) it seems not to "flatline" at any value of $\lambda_{\max}$ and (iii) if it stabilizes, that always happens at a level they could not quantify which is below the $2/\eta$ threshold, see Figure 1, often without a proper progressive sharpening phase.
+[1] noticed that SGD somehow acclimates to the hyperparameters.
+Gilmer et al.
+[27] established effects of initialization and architecture choices on stability of SGD, without providing a definitive stability condition.
+These works did not characterize how, if not negatively, (ii) it seems not to "flatline" at any value of $\lambda_{\max}$ and (iii) if it stabilizes, that always happens at a level they could not quantify which is below the $2/\eta$ threshold, see Figure 2, often without a proper progressive sharpening phase.
 This scenario leaves the most basic questions open: In what way the location of convergence of SGD acclimates to the choice of hyperparameters?
@@
 Does SGD train in an unstable regime?
-Edge of Stochastic Stability.
-We address the questions above by introducing a general framework for SGD dynamics through the concept of Batch Sharpness: the expectation of the Hessians of mini-batch losses along the direction of the mini-batch gradients, see Section 3.1.
-This serves as a generalized and direct substitute for $\lambda_{\max}$ in the mini-batch SGD setting.
-Specifically, we identify a new regime–which we term Edge of Stochastic Stability (EoSS)–where Batch Sharpness stabilizes near $2/\eta$, while maintaining a critical gap with $\lambda_{\max}$.
-This dynamic explains the plateaus of $\lambda_{\max}$ at lower levels determined by batch size and learning rate, unifying the observed phenomena in mini-batch training.
-Organization of the rest of the paper.
-Section 2 reviews related work and outlines key open questions.
-In Section 3, we introduce and empirically validate the phenomenon of Edge of Stochastic Stability.
-Section 4 discusses what the observed notion of stability entails mathematically.
-In Section 5, we examine stability mathematically and we study values at which $\lambda_{\max}$ plateaus.
-In Section 6, we discuss the implications for the structure and modeling of SGD dynamics.
-Finally, Section 7 summarizes our findings and addresses the limitations of this work.
-Figure 1: Mini-batch SGD on CIFAR-10.
-The full-batch Hessian’s largest eigenvalue $\lambda_{\max}$ typically plateaus below $2/\eta$, here $\eta=1/200$.
-Smaller batch sizes lead to lower plateau values.
-Figure 2: SGD at EoSS under different learning rates and batch sizes.
-We train a fully-connected network on an 8k subset of CIFAR-10.
-Batch Sharpness stabilizes at the $2/\eta$ threshold across varying batch sizes and step sizes.
-2 Related Work
-2.1 Progressive Sharpening and Edge of Stability
-Progressive sharpening.
-Early work noticed that the local shape of the loss landscape changes rapidly at the beginning of the training (LeCun et al., 2012; Keskar et al., 2016; Achille et al., 2017; Jastrzębski et al., 2018, 2019; Fort & Ganguli, 2019).
-Multiple papers noticed growth of different estimators of $\lambda_{\max}$ in the early training (Keskar et al., 2016; Sagun et al., 2016; Fort & Scherlis, 2019; Jastrzębski et al., 2019).
-Later, Jastrzębski et al.
-(2019, 2020) and Cohen et al.
-(2021) managed to make this precise, noticing that, along the trajectories of SGD and GD, $\lambda_{\max}$ increases steadily, often after a small number of steps of decrease.
-This phenomenon was called progressive sharpening by Cohen et al.
-(2021).
-A phase transition.
-Early studies (Goodfellow et al., 2016; Li et al., 2019; Jiang et al., 2019; Lewkowycz et al., 2020) revealed that large initial learning rates often enhance generalization at the cost of initial loss reduction.
-Jastrzębski et al.
-(2020) attributed this to a sudden phase transition, termed the break-even point, marking the end of progressive sharpening.
-This transition slows the convergence by confining the dynamics to a region of lower sharpness.
-Unlike progressive sharpening, this phenomenon is considered to be induced by the gradient-based algorithm becoming unstable, not by the landscape.
-Jastrzębski et al.
-(2019, 2020); Cohen et al.
-(2021, 2022) indeed showed that the phase transition comes at different points for different algorithms on the same landscapes.
-Jastrzębski et al.
-(2019, 2020) showed that in the case of mini-batch SGD this phase comes earlier for bigger step sizes and smaller batch sizes, without quantifying it.
-Cohen et al.
-(2021, 2022) later showed indeed that it comes at the instability thresholds for the full-batch optimization algorithms.
-Full-batch edge of stability.
-After the phase transition above, GD and full-batch Adam train in the EoS oscillatory regime (Cohen et al., 2021, 2022), where the $\lambda_{\max}$ stabilizes and oscillates around a characteristic value.
-The name is due to the fact that, in the case of full-batch GD, the $\lambda_{\max}$ hovers at $2/\eta$ which is the stability threshold for optimizing quadratics.
-Observations from Cohen et al.
-(2021, 2022) indicate that most training dynamics occur within this regime, effectively determining $\lambda_{\max}$ of the final solution.
-Lee & Jang (2023) explained why $\lambda_{\max}$ in this regime is generally above $2/\eta$, rarely aligning exactly.
-The reasons include: (i) nonlinearity of the loss gradient, which shifts the required value depending on higher-order derivatives, and (ii) the EoS being governed by the Hessian along the gradient direction, rather than $\lambda_{\max}$ alone.
-EoS and convergence.
-There is a growing body of work analyzing the surprising mechanism of EoS in training dynamics with GD.
-Classically, when gradients depend linearly on parameters, divergence occurs locally if $\eta>\frac{2}{\lambda}$, as demonstrated with one-dimensional parabolas (Cohen et al., 2021).
-Yet, neural networks often converge even when $\eta\geq\frac{2}{\lambda}$, likely due to the problem’s non-standard geometry.
-Damian et al.
-(2023) proposed an explanation under some, empirically tested, assumptions of alignment of third derivatives and gradients.
-Wang et al.
-(2022) showed that the product-like structure of shallow networks guarantees convergence with learning rates exceeding the EoS threshold.
-Beneventano & Woodworth (2025) showed that although convergence is guaranteed, it may be extremely slow, this may, however, benefit implicit regularization effects.
-For further exploration of convergence and implicit regularizarion in the EoS regime, we direct the reader to Arora et al.
-(2022); Ahn et al.
-(2022, 2023); Zhu et al.
-(2023); Lyu et al.
-(2023).
-2.2 SGD, Hessian, and Generalization
-SGD finding flatter minima.
-Our result is inherently a result about mini-batch training improving flatness. Specifically, we explain why: {mdframed}
-Training with smaller batches constraints the dynamics in areas with smaller eigenvalues for the full-batch Hessian.
-This quantifies and characterizes prior observations that SGD tends to locate flat minima (Hochreiter & Schmidhuber, 1997) and that smaller batch sizes result in reduced Hessian sharpness (Keskar et al., 2016).
-Analogously, recent work strongly aligns with our result: Jastrzębski et al.
-(2021) demonstrates that large learning rates in SGD act similarly to penalizing the trace of the Fisher matrix–which approximates the Hessian during training and shares its top eigenspaces (Jastrzębski et al., 2018; Martens, 2020; Thomas et al., 2020)–in image classification tasks.
-Sharpness and Generalization.
-SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance (LeCun et al., 2012; Keskar et al., 2016; Goyal et al., 2017; Jastrzębski et al., 2018; Masters & Luschi, 2018; Smith et al., 2021).
-This advantage has been widely attributed to the flatness of the minima (Murata et al., 1994; Hochreiter & Schmidhuber, 1994; Neyshabur et al., 2017; Wu et al., 2017; Dinh et al., 2017; Kleinberg et al., 2018; Jiang et al., 2019; Xie et al., 2020).
-Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks (Izmailov et al., 2019; Foret et al., 2021).
-Moreover, Jastrzębski et al.
-(2021) shows that, in practice, penalizing the trace of the Fisher information matrix empirically improves generalization and reduces memorization—especially when training with noisy labels.
-We believe that our quantification and the introduction of a theoretical framework for these improvements in sharpness may be a stepping stone towards understanding these generalization benefits theoretically.
+Previous Works on SGD Stability.
+A series of works, Wu et al.
+[28], Ma and Ying [29], Granziol et al.
+[30], Wu et al.
+[31], Mulayoff and Michaeli [32], give sharp step-size stability thresholds for constant-step-size SGD on quadratic losses.
+However, the problem they tackle (is motivated by) but is not if and how SGD shows any EoS-type phenomenon when training neural networks [^1] For instance, in principle the level of stabilization of Frobenius norm for them could be independent of batch size, see Figure 6b in [31], differing from behavior of $\lambda_{\max}$ as described in Figure 2 and by [1, 5, 33])..
+Specifically, they leave open (1) what statistics of the Hessians are impacted by progressive sharpening [^2] I.e., in which directions progressive sharpening happens.
+We show that Batch Sharpness increases—not $\lambda_{\max}$—this implies that the covariance of the Hessians increases at least as fast as the mean; (2) if SGD stabilizes around stationary points or not; (3) in what precise way SGD adapts location of stabilization to hyper parameters, and why.
+These studies model stability of SGD on quadratics in general, we specify how to characterize the stability given the precise progressive sharpening phenomena of the neural networks landscape.
+Moreover, from their analysis is non trivial to understand if Batch Sharpness at $2/\eta$ is a possible edge of stability, thus we proved Theorem 1 with a slightly different approach.
+Empirically, multiple works—e.g., Lee and Jang [17]—establish the presence of oscillations in the SGD trajectory.
+However, did not distinguish between noise and curvature-driven oscillations—the ones directly relevant to the question of how the curvature adapts to the hyper parameters [^3] As we discuss in Section 3, 5 and Appendix G.
+.
+Flatness and Generalization.
+SGD-trained networks consistently generalize better than GD-trained ones, with smaller batch sizes further enhancing generalization performance [2, 7, 9, 34, 35, 36, 37, 38].
+This advantage has been widely attributed to some notion of flatness of the minima [15, 33, 39, 40, 41, 42, 43].
+Training algorithms explicitly designed to find flat minima have indeed demonstrated strong performance across various tasks [44, 45].
+Our result is inherently a result about mini-batch training improving flatness.
+Specifically, we explain why: Training with smaller batches constraints the dynamics to areas with smaller eigenvalues of the full-batch Hessian.
+This quantifies and characterizes prior observations that SGD tends to locate flat minima and that smaller batch sizes result in reduced Hessian sharpness [2, 33].
+(a) Trajectory
+(b) Step size schedule
+(c) Loss
+(d) Zoomed-in
+(e) GNI of Eq. (1)
+Figure 3: Quadratics: Dynamics of SGD on a 1-D quadratic with $N$ datapoints, $L(x)=\frac{1}{2N}\sum_{i}(x-a_{i})^{2}$, where $a_{i}\sim\mathcal{N}(0,1)$.
+At step size, oscillations are present.
+Yet, only when the step size becomes larger than 2/sharpness = 2 (last step size change, red line), the oscillations become unstable (d), the loss diverges (c).
+Meanwhile, the quantity of (1) consistently stays at $2/\eta$, indicating the presence of Type-1 oscillations.
+3 Two Types of Oscillations in SGD
+The key defining aspect of EoS is about the solutions found by the algorithm adapting to the optimizer’s hyperparameters.
+This in the case of full-batch algorithms happens through the appearance of an oscillatory regime.
+Mini-batch SGD, however, always oscillate because its gradient is noisy and the step size does not vanish.
+The central question is therefore which oscillations signal curvature-limited dynamics (EoS-like).
+Both types of oscillation involve quantities stabilizing near the critical threshold of $2/\eta$, yet they differ fundamentally:
+  1. 1.
+    Type-1 Noise-Driven Oscillation.
+SGD can wobble around a stationary point simply because gradients vary across batches and the step size is not annealed.
+This occurs even if the Hessian is small—can arise in flat regions of the landscape—as, fixed–step-size, mini-batch noise alone can induce perpetual “wiggling”.
+Such noise-driven behavior is well-documented in empirical work [17] and well-studied in classical stochastic approximation [46, 47, 48, 49].
+  2. 2.
+    Type-2 Curvature-Driven Oscillation.
+Here, the Hessian’s magnitude and the step size interact so that the algorithm repeatedly “overshoots” the current valley, creating catapult-like surges in loss.
+In other words, once the local, or perceived, curvature is too large relative to $2/\eta$, the updates become unstable in a manner analogous to the classic EoS [1].
+Type-1 oscillations occur even in simple quadratic settings no matter the step size—Figure 3.
+Proposition 1.
+Denote $L_{B}(\theta)$ the minibatch loss, $L(\theta)$ the full-batch loss, and $\mathcal{H}$ its Hessian, $\mathcal{P}_{b}$ the distribution of the batches of size $b$.
+Assume the trajectory oscillates (no matter the Type) then
+[…]
+Crucially, the appearance of some quantity—Gradient-Noise Interaction (GNI)—being $2/\eta$ implies the system is oscillating, not why.
+It does not mean, in principle, that the landscape or the curvature adapted to the hyper parameters.
+In the case of Type-1, $2/\eta$ is about the ratio between the covariance of the gradients and the size of the full-batch gradient, independently of the curvature [^4] See Appendix A for the proof of Proposition 1 and a classical treatment of Type-1 oscillations, which characterize any stochastic approximation/otpimization algorithm whenever the step size $\eta_{k}$ is not annealed such that $\sum_{k=1}^{\infty}\eta_{k}<\infty$..
+Importantly, a way to differentiate between the two types of oscillation is by changing the hyper parameters and observing the outcome.
+When the size of oscillations increases (bigger step or smaller batch) the dynamics diverges—catapults away—for Type-2 oscillations; it just increases the size of the oscillations—quickly restabilizing—for Type-1, see Figure 3 and Definition 2 in Section 5.
+(a) LR increased early
+(b) LR increased late
+(c) Batch decreased early
+(d) Batch decreased late
+[…]
 4 SGD Typically Occurs at the EoSS
-Extending Cohen et al.
-(2021), we identify a stable quantity in mini-batch SGD: Batch Sharpness.
-We empirically demonstrate that SGD operates in a regime we term Edge of Stochastic Stability, observed across multiple datasets and architectures.
-Batch Sharpness emerges as the key quantity governing this regime, serving as the analogous for mini-batch training of the sharpness.
+We introduce in this section Batch Sharpness, our notion of implicit curvature.
+We proceed characterizing the EoSS phenomenon, and establishing it empirically.
+We empirically find that when Type-2 oscillations occur, Batch Sharpness aligns closely with $2/\eta$.
+Interestingly, throughout most of the training—even during progressive sharpening—also Type-1 oscillations are present.
+4.1 Mini-Batch Sharpness
+Throughout this paper $B\subset\mathcal{D}$ denotes a random mini-batch of size $b$ drawn from a fixed sampling distribution $\mathcal{P}_{b}$.
+For model parameters $\theta\in\mathbb{R}^{d}$ let $L_{B}(\theta)=\frac{1}{b}\sum_{(x_{i},y_{i})\in B}\!\tilde{L}\!\bigl{(}f_{\theta}(x_{i}),\,y_{i}\bigr{)}$, $L(\theta)=\mathbb{E}_{B\sim\mathcal{P}_{b}}\!\bigl{[}L_{B}(\theta)\bigr{]}$ be the mini-batch and full-batch losses, respectively.
+Write $\mathcal{H}(L_{B})=\nabla_{\theta}^{2}L_{B}(\theta)$.
+Definition 1 (Batch Sharpness).
+We define Batch Sharpness as the ratio
+Batch $\textit{Batch Sharpness}\,(\theta)\quad:=\quad\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\nabla L_{B}(\theta)^{\top}\,\mathcal{H}(L_{\mathbf{B}})\,\nabla L_{B}(\theta)\bigr{]}}{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\|\nabla L_{\mathbf{B}}(\theta)\|^{2}\bigr{]}}\,.$ (2)
+Interpretation.
+For each mini-batch, $\nabla L_{B}$ is the stochastic search direction produced by SGD.
+Thus the normalized version of $\nabla L_{B}^{\top}\,\mathcal{H}(L_{B})\,\nabla L_{B}$, Batch Sharpness, measures the expected directional curvature of the mini-batch loss surface along the corresponding mini-batch gradients.
+For locally quadratic losses a constant step $\eta$ is, on average, stable on the mini-batch landscape iff Batch Sharpness< $\textit{Batch Sharpness}\,<\,2/\eta$, see Lemma 1; crossing the threshold induces the catapult–type divergence that marks the EoSS regime, see Section 5.
+Relation to earlier notions.
+Lee and Jang [17] showed that the ratio $\nabla L^{\top}\mathcal{H}\nabla L/\|\nabla L\|^{2}$ settles at $2/\eta$ during EoS and coincides with $\lambda_{\max}(\mathcal{H})$ only when the gradient aligns with the top eigenvector.
+Batch Sharpness extends that directional viewpoint to the stochastic setting by replacing $(\nabla L,\mathcal{H})$ with their mini-batch counterparts $\bigl{(}\nabla L_{B},\mathcal{H}(L_{B})\bigr{)}$ before taking expectations [^6] Importantly, see Appendix L, the right notion of curvature/stability for mini-batch algorithms has to depend on different statistics or moments of the mini-batch Hessians, not simply on the average of the Hessians as in full-batch.
+Batch Sharpness, interestingly, translates to a ratio between third and second moment in the quadratic setting..
+4.2 The Edge of Stochastic Stability
+We characterize here the phenomenon of the Edge of Stochastic Stability.
+We verify the emergence of EoSS across of a range of step sizes, batch sizes and architectures (Figure 6 and Appendix M); datasets (CIFAR-10 and SVHN, Appendix N); and dataset sizes (8k and 32k subsets, Figure 7).
+1. Stabilization of Batch Sharpness.
+SGD typically operates in a regime analogous to EoS: {mdframed}
+SGD tends to train in a regime we call Edge of Stochastic Stability. Precisely, after a phase of progressive sharpening, Batch Sharpness reaches a stability level of $2/\eta$, and hovers there.
+In particular, the level of plateau of Batch Sharpness is $2/\eta$ independent of the batch size (Figure 1).
+Importantly, Type-1 oscillations happen throughout most of the training as highlighted by the quantity of Proposition 1, see Figure 4, but they do not impact progressive sharpening which leads to the second phase of EoSS stabilization and Type-2 oscillations.
+Importantly, analogously to EoS, training continuous and the loss continues to decrease while Batch Sharpness is constrained by the learning rate size.
+2. Stabilization of $\lambda_{\max}$ and GNI.
+Crucially, stabilization of Batch Sharpness around $2/\eta$ happens while GNI is stabilized at $2/\eta$ and induces a corresponding stabilization of $\lambda_{\max}$.
+However, $\lambda_{\max}$ consistently settles at a lower level, due to a batch-size–dependent gap between the two (Figures 6).
+This is also influenced by the specific optimization trajectory (Figure 7).
+We make observations on this relationship and the factors determining the stabilization gap in Section 6.
+3. Catapults.
+Figure 5: Catapults at EoSS. During EoSS, Batch Sharpness goes through cycles of progressive sharpening and stabilizations. Notations follow Figure 6.
+Unlike in EoS, in the EoSS regime what is stabilized is the expectation of a quantity which the algorithm sees one observation at time.
+Occasionally, a sequence of sampled batches exhibits anomalously high sharpness–that is too high for the stable regime–and steps overshoot, triggering a catapult effect, where Batch Sharpness spikes before rapidly decreasing (Figure 5).
+This is typically followed by renewed progressive sharpening, eventually returning to the EoSS regime.
+This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, e.g., [16, 50].
 Refer to caption Refer to caption Refer to caption
@@
 Figure 6: Comparing different sharpness measures.
-Red: observed sharpness on the step’s mini batch on the batch used for the step; Green: average Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$.
+Red: observed sharpness on the step’s mini batch on the batch used for the step; Green: Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$.
 Top row: MLP (2 hidden layers of width 512); middle: 5-layer CNN; bottom: ResNet-14; all trained on an 8k subset of CIFAR-10.
-3.1 Mini-Batch Sharpness
-Let $L_{B}(\theta)\!=\!\frac{1}{b}\!\sum_{(x_{i},y_{i})\in B}\!\tilde{L}\bigl{(}f_{\theta}(x_{i}),y_{i}\bigr{)}$ be the empirical loss on a randomly sampled batch $B$ of size $b$.
-Denote by $\mathcal{H}(L_{B})$ the Hessian w.r.t.
-$\theta$ of $L_{B}$, and let $\nabla L_{B}(\theta)$ be its gradient.
-We define:
-Definition 1 (Batch Sharpness).
-Let $\mathcal{P}_{b}$ be the distribution of random batches of size $b$. The Batch Sharpness is
-Batch $\text{Batch Sharpness}\;=\;\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\nabla L_{B}^{\top}\,\mathcal{H}(L_{B})\,\nabla L_{B}\bigr{]}}{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\|\nabla L_{B}\|^{2}\bigr{]}}\,.$
-Denote by $\lambda_{\max}:=\lambda_{\max}\left[\mathcal{H}(L)\right]$ the largest eigenvalue of the Hessian of the full-batch loss.
-While Batch Sharpness behaves as $\lambda_{\max}$ in EoS, the two quantities differ fundamentally.
-We explore this distinction further in Section 4.
-3.2 The Edge of Stochastic Stability
-1. Stabilization of Batch Sharpness:
-We establish that SGD operates in a stability regime analogous to EoS, which we term Edge of Stochastic Stability. In this regime Batch Sharpness stabilizes at $2/\eta$: {mdframed}
-SGD tends to train in an instability regime we call Edge of Stochastic Stability.
-Precisely, after a phase of progressive sharpening, Batch Sharpness reaches a stability level of $2/\eta$, and hovers there.
-2. Stabilization of $\lambda_{\max}$:
-Also $\lambda_{\max}$ typically stabilizes.
-However, it does so at lower values that depends on the trajectory taken (Figures 3 and 5).
-We discuss the stabilization level of $\lambda_{\max}$ in Section 3.3 and 5.
-Consequently, standard EoS analysis that relies solely on $\lambda_{\max}$ cannot explain SGD’s actual training behavior in the small-batch setting.
-3. Catapults:
-Unlike EoS, EoSS stabilizes the expectation of a quantity which the algorithm sees one observation at time.
-Occasionally, a series of batches sampled have outlying sharpness–that is too high for the stable regime–and steps overshoot, triggering a catapult effect, where Batch Sharpness spikes before rapidly decreasing (Figure 4).
-This is typically followed by renewed progressive sharpening, returning to the EoSS regime.
-This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, see, e.g., Lewkowycz et al.
-(2020); Zhu et al.
-(2024).
-Figure 4: Catapults during EoSS.
-During EoSS, Batch Sharpness may through cycles of progressive sharpening and stabilizations.
-Red dots are the observed sharpnesses of the batches used for the step, the green line is Batch Sharpness.
+4.3 Batch Sharpness Governs EoSS
+Following Cohen et al.
+[1] and the discussion in Section 3, we track how the training dynamics change when perturbing the hyper-parameters mid-training.
+Overall, we find that Batch Sharpness governs EoSS behavior—mirroring how $\lambda_{\max}$ operates in the full-batch EoS—while the full-batch $\lambda_{\max}$ lags behind or settles inconsistently, underlining the mini-batch nature of SGD stability.
+Inducing catapults.
+Increasing the step size $\eta$ or decreasing the batch size $b$ triggers a catapult spike in all the quantities in considerations and the training loss, before Batch Sharpness re-stabilizes near the updated threshold $2/\eta$, see Figures 7(a) and 7(c).
+This therefore pushes $\lambda_{\max}$ lower.
+Crucially, if stability was governed solely by $\lambda_{\max}$, this step-size or batch-size adjustment would have left the dynamics unchanged, as $\lambda_{\max}$ remains below both the original and revised $2/\eta$ thresholds.
+Re-entering progressive sharpening.
+Conversely, reducing $\eta$ raises the $2/\eta$ threshold.
+Conversely, increasing the batch size reduces the gap between Batch Sharpness and $\lambda_{\max}$.
+These changes prompt a new phase of progressive sharpening of Batch Sharpness, see Figures 7(b) and 7(d).
+Here, $\lambda_{\max}$ also rises, but ultimately stabilizes at a lower value than if the entire training had run with the smaller step size.
+Again, if stability was governed by $\lambda_{\max}$, this step-size adjustment would have had the same effect as starting from scratch with the new step size.
+Notably, Batch Sharpness initially decreases when reducing the learning rate—a counterintuitive result potentially linked to shifting alignments between gradients and Hessian eigenvectors, warranting further investigation.
 (a) Increasing step size
@@
 Figure 7: Effects of changing step size or batch size in EoSS.
-(a) Increasing the learning rate causes a catapult spike before Batch Sharpness settles at the new $2/\eta$.
+(a) Increasing the learning rate causes a catapult spike before Batch Sharpness re-settles at the new $2/\eta$.
 (b) Decreasing the learning rate prompts renewed progressive sharpening.
-(c) Increasing the batch size narrows the gap between Batch Sharpness and $\lambda_{\max}$.
-(d) Decreasing the batch size.
-Triangles mark the steps at which the hyperparameters were changed.
-3.3 Batch Sharpness Governs EoSS
-Following Cohen et al.
-(2021) we show here how the dynamics of Batch Sharpness changes when hyper-parameters shift along the dynamics.
-Interestingly, while $\lambda_{\max}$ and other full-batch Hessian metrics may stabilize, we show here their specific plateaus or trajectories depend on initial conditions, not only on the hyperparameter.
-By contrast, $\lambda_{\max}(\mathcal{H}(L_{B}))$—the top eigenvalue of the mini-batch Hessian—does consistently stabilizes and shift with hyperparameters.
-Changing the step size.
-Increasing the step size $\eta$ triggers a catapult spike in all the quantities in considerations and the training loss, before Batch Sharpness re-stabilizes near the updated threshold $2/\eta$ (Fig.
-5(a)).
-This therefore pushes $\lambda_{\max}$ lower.
-Crucially, if stability was governed solely by $\lambda_{\max}$, this step-size adjustment would have left the dynamics unchanged, as $\lambda_{\max}$ remains below both the original and revised $2/\eta$ thresholds.
-Conversely, reducing $\eta$ (Fig.
-5(b) raises the $2/\eta$ threshold and prompts a new phase of progressive sharpening of Batch Sharpness.
-Here, $\lambda_{\max}$ also rises, but ultimately stabilizes at a lower value than if the entire training had run with the smaller step size.
-Again, if stability was governed solely by $\lambda_{\max}$, this step-size adjustment would have had the same effect as starting from scratch with the new step size.
-Notably, Batch Sharpness initially decreases—a counterintuitive result potentially linked to shifting alignments between gradients and Hessian eigenvectors.
-This behavior warrants further investigation.
-Changing the batch size.
-Increasing the batch size reduces the gap between Batch Sharpness and $\lambda_{\max}$, therefore making Batch Sharpness lower,(Fig.
-5(c)), and thus below the $2/\eta$ threshold.
-This causes a new progressive sharpening phase until Batch Sharpness again reaches the $2/\eta$ level, making $\lambda_{\max}$ stabilize at a new, higher level, again lower than the one of the training started from scratch with the new batch size.
-This shows that it is indeed the Batch Sharpness that is the "limiting" quantity in EoSS.
-Conversely, decreasing the batch size induces a brief catapult phase, after which Batch Sharpness returns to its previous plateau, as expected.
-This forces full-batch $\lambda_{\max}$ to adjust to a lower level.
-Overall, we find that Batch Sharpness governs EoSS behavior—mirroring how $\lambda_{\max}$ operates in the full-batch EoS—while the actual full-batch $\lambda_{\max}$ lags behind or settles inconsistently, underlining the mini-batch nature of SGD stability.
-4 Mathematics of Stability
-It may seem surprising at first that Batch Sharpness is not a quantity of the same nature as $\lambda_{\max}$ or is not a function of the full-batch Hessian.
-The reason is deep into the nature of mini-batch SGD.
-After recalling the computations of Cohen et al.
-(2021) in Section 4.1, we show what is the setting of mini-batch SGD and what is the notion of stability which depends on Batch Sharpness in Section 4.2.
-Later in Section 4.3 we motivate why the stability of mini-batch SGD, surprisingly, can not depend on full-batch quantities.
-4.1 A Quadratic Objective: Full-Batch GD
-We recall here the arguments of Cohen et al. (2021) and Lee & Jang (2023) for quadratic objective functions to later show how stability differs between the two settings.
-Assume we fix a learning rate $\eta$ for the whole training. Consider the example of the quadratic Taylor approximation of a loss function:
-$f(x)=\frac{1}{2}x^{\top}\mathbf{A}x+\mathbf{b}^{\top}x+\mathbf{c},$
-where $\mathbf{A}$ is the Hessian matrix (symmetric and positive semidefinite), $\mathbf{b}$ is the gradient vector, and $\mathbf{c}$ is a constant.
-Here, as Cohen et al.
-(2021) observed, when $\eta\lambda_{\max}(\mathbf{A})>2$ the component of $x_{t}-x^{*}$ parallel to the eigenvector relative to $\lambda_{\max}(\mathbf{A})$ explodes with a power law.
-This is equivalent to imposing the inequality (1) at every step.
-[…]
-Indeed, eventually, we have $f(x_{t+1})>f(x_{t})$, i.e., if and only if $\lambda_{\max}\leq 2/\eta$.
-4.2 Quadratic Objectives: Mini-Batch GD
-Empirically, progressive sharpening halts when the mini-batch analog of (1) reaches zero in expectation. Concretely, EoSS arises if, at each step,
-$\eta\,\|\nabla f_{B}(x_{t})\|^{2}\;\sim\;\frac{\eta^{2}}{2}\,\nabla f_{B}(x_{t})^{\top}\mathbf{A}_{B}\nabla f_{B}(x_{t}),$ (2)
-preventing the single-batch loss from diverging.
-This makes sense, as the gradient step is on the mini-batch loss, thus the training is stuck in an oscillatory regime when the single batch loss oscillates (on average) [^1] After completion of this manuscript we found Lee & Jang (2023).
-We compare this with their work in Appendix C..
-One might expect this to imply $\lambda_{\max}(\mathcal{H}(L_{B}))\leq 2/\eta$, as in full-batch (1), but that is typically not the case.
-What is inducing (1) to rely on $\lambda_{\max}$ is that the Hessians at every step share the eigenvectors relative to the highest eigenvalue.
-In mini-batch SGD, the sampled Hessians $\mathbf{A}_{B}$ generally have completely different eigenspaces than $\mathbf{A}$ and between each others.
-Indeed, the case of mini-batch SGD, we are working with the following quadratic Taylor approximation:
-$f_{B}(x)=\frac{1}{b}\sum_{i\in B}\left(\frac{1}{2}x^{\top}\mathbf{A}_{i}x+\mathbf{b}_{i}^{\top}x+\mathbf{c}_{i}\right),$ (3)
-where $B\subset\{1,2,\ldots,n\}$ is a randomly sampled mini batch of size $b$, and $\mathbf{A}_{i}$, $\mathbf{b}_{i}$, and $\mathbf{c}_{i}$ represent coefficients on the quadratic losses of individual data points.
-Here the full-batch Hessian $\mathbf{A}$ is given by $\mathbf{A}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i}$.
-Even if $\lambda_{\max}(\mathbf{A})>2/\eta$ with eigenvector $v_{\max}$, certain $\mathbf{A}_{i}$ may map $v_{\max}$ to 0 or a completely different vector, avoiding the blow-up seen in the full-batch regime.
-Enforcing (2) in this setting (with $\mathbf{b}_{i}=0$) thus results in
-$\frac{2}{\eta}\ \geq\ \sup_{v\in\mathbb{S}^{d}}\ \frac{v^{\top}\mathbb{E}_{i}\left[(\mathbf{A}_{i})^{3}\right]v}{v^{\top}\mathbb{E}_{i}\left[(\mathbf{A}_{i})^{2}\right]v}\ \begin{subarray}{c}\text{ which does not}\\ \text{simplify to}\end{subarray}\ \lambda_{\max}(\mathbf{A}).$ (4)
-4.3 Why Stability cannot Depend on Full-Batch
-We show here with an example that in the mini-batch setting the stability thresholds can not depend only on full-batch Hessian or gradients, but it has to depend on the higher moments of them over batch sampling, as we can always construct a counterexample which diverges otherwise.
-Imagine on two data points we have $\mathbf{A}_{1}=\alpha\cdot I+M$ and $\mathbf{A}_{2}=\alpha\cdot I-M$, with $\alpha,\gamma>0$ and $M=\begin{pmatrix}0&0\\ 0&\gamma\end{pmatrix}$, and $\mathbf{b}_{1}=\mathbf{b}_{2}=0$.
-Here $\mathbf{A}$ is $\alpha I$.
-$\mathbf{A}_{1}$ has eigenvalues $\alpha,\alpha+\gamma$ and $\mathbf{A}_{2}$ has eigenvalues $\alpha,\alpha-\gamma$.
-Let us now look at possible full-batch stability results, as developed by Cohen et al.
-(2021).
-If the right stability notion for mini-batch SGD depended on the full-batch Hessian or gradients, then it would be independent on $\gamma$, and this can not be the case.
-For instance, if a notion as the one in Equation (1) would explain stability, then $\alpha$ could grow until $\alpha=2/\eta$.
-However, no matter the size of $\alpha$, if $\gamma>\max\big{\{}\alpha,\tfrac{2}{\eta}-\alpha\big{\}}$ a mini-batch iteration $x_{t+1}=(1-\eta\mathbf{A}_{i})x_{t}=$ increases the size of the second component of $x$ by a multiplicative factor bigger than 1.
-Thus a study of the stability of the system can not depend on the full-batch Hessian but has to depend on how big the oscillations due to the size of $\gamma$ are, i.e., on the higher moments of the distribution of the mini-batch Hessian.
-Note that this situation would be even more extreme if we had (as in the practice of deep learning) the top eigenvectors of the mini-batch Hessians to point in completely different directions, not just to have high variance.
-As a sanity check, any stability threshold dependent on the higher moments of the mini-batch Hessian or gradients–as Batch Sharpness or $\mathbb{E}_{B}[\lambda_{\max}(\mathcal{H}(L_{B})]$–would induce $\eta_{\max}$ to depend on $\gamma$ too.
-5 The Geometry of the Landscape at EoSS
-We devote this section to the discussion of the sizes during training of different quantities which may be or interest of practitioners of theoreticians, as the $\lambda_{\max}$ of full-batch and mini-batch Hessians in terms of learning rate, batch size, and critical batch size.
-For a batch size $b$, denote by
-$\lambda_{\max}^{b}:=\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\lambda_{\max}(\mathcal{H}(L_{B}))\big{]}.$
-{mdframed}
-Then, we establish here that:
-  • Also $\lambda_{\max}^{b}$ stabilizes.
-  • $\lambda_{\max}^{b}$ stabilizes at a level that ranges between $4/\eta$ and $2/\eta$. The level is lower for very small and very large batch sizes.
-  • $\lambda_{\max}^{b}$ is always greater than $\lambda_{\max}$.
-  • We find that the level at which $\lambda_{\max}$ stabilizes is characterized by two different regimes.
-The threshold is what we call critical batch size.
-This critical batch size depends on the complexity of the data-model.
-  • However, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ typically goes as $1/b^{\alpha}$ generally with $\alpha=0.70$.
-This is surprising, as when fixing the network, the gap goes as $1/b$ or $1/b^{1/2}$ both empirically and mathematically.
-5.1 Average $\lambda_{\max}^{b}$ for the mini-batch Hessians
-We establish that $\lambda_{\max}^{b}$–the average over the possible batches sampled of the highest eigenvalue of the mini-batch Hessian–generally stabilizes at a value just bigger than $2/\eta$.
-This happens on a wide range or models and datasets.
-We observe that its stabilization levels is generally very close to $2/\eta$ for very small batch sizes, it increases until the critical batch size, then it decreases again.
-Interestingly, we observe that $\lambda_{\max}^{b}$ responds to the change of parameters exactly as Cohen et al. (2021) observed for $\lambda_{\max}$ in the full-batch setting.
-5.2 Size of $\lambda_{\max}$
-Although we show that the quantity of interest is the mini-batch Hessian, in case of mini-batch SGD, there are both historical and generalization-related reasons to investigate what the spectrum of the full-batch Hessian is.
-We establish that also $\lambda_{\max}$ generally stabilizes in the EoSS at a level which is strictly smaller than $2/\eta$.
-We show that the exact level at which $\lambda_{\max}$ does not only depend on learning rate and batch size but on the full trajectory taken, unlike for EoS (Cohen et al., 2021).
-Indeed, in Section 3.3 we showed that it does not fully adapt to changes in hyper parameters.
-However, we observed that when we do not induce any shock in the system by changing hyper parameters during training, we can characterize two regimes for the stabilization levels, see Figure 6.
-Precisely
-  • (i) $b\leq$ critical batch size: $\lambda_{\max}$ grows with $b$ with a power law.
-  • (ii) $b\geq$ critical batch size: $\lambda_{\max}$ grows with $b$ much slower while approaching $2/\eta$ from below.
-Unfortunately, we could not find a way to relate the size of $\lambda_{\max}$ to learning rate and batch size–see Figure 9–if not through the gap it has with the average $\lambda_{\max}$ of the mini-batch Hessian $\lambda_{\max}^{b}$.
-Notably, the critical batch size we mention is the point beyond which training enters the full-batch regime.
-Works as Zhang et al.
-(2024) study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation".
-We conjecture there may be a relation between the two quantities and we leave it to future work.
-5.3 The Gap Between $\lambda_{\max}$ and $\lambda_{\max}^{b}$
-Note that conceptually $\lambda_{\max}^{b}$ is inherently of the same nature as $\lambda_{\max}$, with the only difference being the "part" of the loss it is computed on-the mini-batch loss for the former, and full-batch loss for the latter.
-Yet, this difference proves to be crucial in being descriptive of the SGD training dynamics.
-Indeed, e.g.: {mdframed}
-While the mini-batch gradients are an unbiased estimator of the full-batch gradient, $\lambda_{\max}\big{(}\mathcal{H}(L_{B})\big{)}$ is a positively biased estimator of $\lambda_{\max}$.
-We indeed show in Lemma 4, that $\lambda_{\max}^{b}$ is strictly bigger than $\lambda_{\max}$ and, as a consequence, decreasing the batch size the gap increases. In particular,
-$\lambda_{\max}^{1}\ <\ \lambda_{\max}^{4}\ <\ \lambda_{\max}^{16}\ <\ \lambda_{\max}^{64}\ <\ \lambda_{\max}^{256}\ <\ \ldots$ (5)
-Figure 6: Batch Sharpness vs.
-$\lambda_{\max}^{b}$ vs.
-$\lambda_{\max}$ at EoSS.
-The values of Batch Sharpness (green), mini-batch $\lambda_{\max}^{b}$ (orange), and full-batch $\lambda_{\max}$ (blue) against varying batch sizes.
-We now empirically show what the gaps $\lambda_{\max}^{b}-\lambda_{\max}$ are, and we discuss why that is surprising.
-As shown in Appendix G and Figure 9, the log-log plot of the gap against batch size reveals three distinct regimes:
-  • Very small $b$: for extremely small batch sizes, the gap plateaus. This is unsurprising, as we are in the (i) regime for $\lambda_{\max}$.
-  • Typical/moderate $b$: the gap consistently scales as $1/b^{0.7}$ in the experiment.
-This trend dominates for a wide range of practical batch sizes, highlighting the systematic scaling behavior of the $\lambda_{\max}^{b}-\lambda_{\max}$ gap.
-It is important to note that we establish in Appendix G the exponent to be $0.7$ independently from the depth, the seeds, the subsets of Cifar10 considered, and the architecture (MLP, CNN, ResNets).
-Also note that the critical batch size falls within this range.
-The limitation of this observation in describing $\lambda_{\max}$ is that this consistent trend is consequence both of $\lambda_{\max}$ growing and $\lambda_{\max}^{b}$ decreasing, not just one of the two.
-  • Very big $b$: here generally $\lambda_{\max}^{b}$ is essentially $\lambda_{\max}$ and the gap is converging to zero, we conjecture that some form of scaling limit from random matrix theory applies here.
-The scaling law is generally of the form: $1/b^{\alpha}$ with $\alpha\geq 1$.
-The exact value of $\alpha$ may vary with the setting, but the overall trend is robust.
-While the behavior for very small batches is not surprising, for bigger batch sizes it is.
-In computer vision tasks where there are way more parameters than datapoints, we observe that the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ decreases linearly with $1/b$ when evaluating the $\lambda_{\max}^{b}$ of any fixed model with different batch sizes, see Figure 8.
-This $1/b$ scaling is also what one expects from the analysis in Appendix F.
-6 Implications: Why Noise-Injected GD Differs from Mini-Batch SGD
-6.1 SGD vs. Noisy Gradient Descent
+(c) Increasing the batch size lowers Batch Sharpness and re-starts progressive sharpening.
+(d) Decreasing the batch size increases Batch Sharpness and causes a catapult.
+The experiments are conducted on a 32k subset of CIFAR-10 to ensure sufficient complexity remains in the dataset, which is necessary for observing renewed progressive sharpening, consistent with observations by Cohen et al.
+[1].
+5 On Stability
+The previous section empirically demonstrated that mini-batch SGD generally settles into the EoSS regime, where Batch Sharpness hovers around $2/\eta$.
+But why does this threshold matter, and what does it mean for the algorithm’s stability?
+In classical (full-batch) gradient descent, the condition $\eta<2/\lambda_{\max}$ guarantees local stability by preventing divergence along the direction of the largest eigenvalue of a fixed Hessian.
+Conversely, the stability of mini-batch SGD is complicated by the variability of Hessians across iterations, making traditional stability criteria insufficient, see Appendix L.
+In this section, we investigate what the notion of stability for SGD on the landscapes of neural networks is and clarify how mini-batch SGD’s “edge” behavior emerges from the evolving distribution of mini-batch Hessians aligned with their own gradients—rather than from any single, fixed Hessian.
+5.1 Mathematizing Instability
+For our analysis, the relevant definition of Type-2 stability must capture the scenario in which perturbations of the step size trigger significant shifts to different regions of the loss landscape—rather than merely resulting in marginal re-stabilization at slightly higher loss values.
+This perspective naturally leads to a curvature-based criterion that distinguishes genuinely disruptive oscillations from benign noise-driven fluctuations.
+Specifically, when mini-batch loss surfaces are locally well-approximated by quadratics along the directions of SGD steps, the nature of the oscillations becomes governed by these local quadratic approximations.
+Under such conditions, oscillations qualify as disruptive precisely if they would result in divergence when evaluated on the locally approximating quadratic landscape.
+Definition 2 (Quadratic instability and Catapults).
+Consider the quadratic approximation of all the data point landscapes $(\theta-x_{i})^{\top}\mathcal{H}_{i}(\theta-x_{i})$.
+We say that a step size is unstable if the trajectory exits all the compacts in which the quadratic approximation holds up to $\mathcal{O}(\eta)$ [^7] This means that either SGD seen as a linear dynamical systemis diverging or that the re-stabilization would happen be at a level which exits the largest region in which the quadratic approximation holds and so the dyanmics changes region..
+We say the algorithm experienced a catapult when this event happened.
+Under this definition, Type-1 oscillations are correctly classified as stable.
+Even if increasing the step size causes these oscillations to locally increase in size but re-stabilize at the new level, the divergence remains bounded, remaining within a compact region, see for instance Figures 3 and 4.
+5.2 Stability and Batch Sharpness
+The next theorem implies that SGD is quadratically unstable in two cases, when $\lambda_{\max}>2/\eta$, but also more surprisingly when $\lambda_{\max}<2/\eta$ but Batch Sharpness is bigger than $2/\eta$.
+While in the first case one can show very easily that the parameters explode, in the second setting we can not and we instead show that the second moment of the mini-batch gradients $\mathbb{E}[\lVert\nabla L_{B}\rVert^{2}]$ increases exponentially.
+Theorem 1.
+Let $\eta\leq 2/\lambda_{\max}$.
+There exists a absolute constant $c>0$ such that if Batch Sharpness is strictly bigger than $2/\eta+c\eta$ then $\mathbb{E}[\lVert\nabla L_{B}\rVert^{2}]$ locally increases exponentially with the SGD step and the trajectory is quadratically unstable (Definition 2).
+The proof is a little involved but essentially relies on Jensen and Cauchy-Schwarz inequalities, see Appendix C and Proposition 5.
+Path dependency and implications on progressive sharpening.
+understanding the inner working and the phenomenology of progressive sharpening is a key open ingredient to understand stability and location of convergence of optimization algorithms.
+In principle, instability could emerge while moving toward infinity in any possible direction on the space.
+Previous work dealt with sufficient and necessary conditions for the stability of SGD in general, [29, 32], i.e., when does SGD become unstable given any possible notion of progressive sharpening.
+We show that the trajectory goes to infinity in precise directions, the fact that Batch Sharpness is the quantity to look at implies that progressive sharpening is acting in a certain way on the datapoint Hessians.
+Precisely, that progressive sharpening takes the trajectory in that place of the boundary of the high dimensional open set of stability such that $GNI=BatchSharpness=2/\eta$.
+Empirically, indeed, the iterates become unstable precisely because of Batch Sharpness reaching $2/\eta$, not because of, e.g., $\lambda_{\max}$ increasing but its variance remaining stable (which would keep Batch Sharpness $<GNI-constant$ until the instability threshold.
+This, we think, shows one more time why it is key for a deep learning optimization theory to advance to focus on path-dependent properties of the correct paths.
+Importantly, in Appendix B we also establish exactly when $GNI\geq\textit{Batch Sharpness}$.
+This fact depends only on the kurtosis of the gradients (which rarely changes much), $\mathbb{E}[\lVert\mathcal{H}(L_{B})-\mathcal{H}\rVert^{2}_{2}]$, and the alignment of the steps with the full-batch Hessian.
+We conjecture that Batch Sharpness reaches $GNI$ at $2/\eta$ by means of increasing $\mathbb{E}[\lVert\mathcal{H}(L_{B})-\mathcal{H}\rVert^{2}_{2}]$ more than the alignment of gradients with the top eigenvectors the full-batch Hessian.
+This would imply that progressive sharpening increases the variance of the Hessian over the batches at least as fast as $\lambda_{\max}$.
+This conjecture would agree with the red cloud amplifying around the green line in Figure 6.
+Stability on the mini-batch step.
+It is also worth commenting on maybe the most unsurprising property of EoSS: that, up to higher orders in the step size, the average single mini batch step is stable on the mini-batch landscape if and only if Batch Sharpness is below $2/\eta$.
+Curiously, however, we were not able to deduce properties of the stability of the whole system solely from this observation, Theorem 1 was proved unrelatedly.
+Lemma 1.
+$\mathbb{E}_{\begin{subarray}{c}\mathcal{P}_{b}\end{subarray}}[L_{B}(x_{t+1})-L_{B}(x_{t})]\leq O(\eta^{2})$ if and only if $\eta\leq 2/\textit{Batch Sharpness}$.
+If the mini-batch landscapes satisfy the PL condition, also the mini-batch gradients do not blow up $\mathbb{E}_{\begin{subarray}{c}\mathcal{P}_{b}\end{subarray}}[\lVert\nabla L_{B}(x_{t+1})\rVert^{2}-\lVert\nabla L_{B}(x_{t})\rVert^{2})]\leq O(\eta^{2})$.
+The importance of stating this result is that Lemma 1 together with Theorem 1 imply that: the curvature-driven oscillations on the average mini-batch landscape imply full-batch loss diverges; the oscillations on the full-batch landscape which are stable on the mini-batch one do not.
+When the dynamics on the mini-batches is stable (Type-1), the dynamics is stable and progressive sharpening still can happen (see Figure 4) until Batch Sharpness reaches $2/\eta$.
+6 On the fate of ${\lambda_{\max}}$
+In this section we examine how $\lambda_{\max}$ behaves once EoSS is reached and clarify its relationship to Batch Sharpness.
+A key aspect of the original EoS analysis is, indeed, that the controlling quantity—the largest eigenvalue of the full-batch Hessian $\lambda_{\max}$—has an immediate geometric interpretation.
+There exists an extensive literature about $\lambda_{\max}$ size and role in neural networks, and it is a main ingredient of any proof of convergence.
+The EoSS picture replaces $\lambda_{\max}$ with Batch Sharpness, a statistic whose connection to generalization and role in optimization theory is largely unexplored.
+(a) MLP, $\eta=0.01$
+(b) MLP, $\eta=0.004$
+(c) CNN, $\eta=0.03$
+Figure 8: Stabilisation level of $\boldsymbol{\lambda_{\max}}$ across learning rates and architectures.
+Top: final-epoch $\lambda_{\max}$ vs.
+batch size.
+Bottom: log–log plots of the gap  $\,2/\eta-\lambda_{\max}\,$ for the same runs.
+All experiments use CIFAR-10 8k; further details in Appendix H.
+6.1 Empirical facts
+Below the phenomena we extensively observe in vision classification tasks trained with MSE, ablating on batch sizes, learning rates, architectures, datasets.
+See Figure 6 for a good reference of what generally goes on.
+  • Fact 1:  Progressive Sharpening. $\lambda_{\max}$ increases at most as long as Batch Sharpness increases.
+  • Fact 2:  Phase Transition. Once Batch Sharpness plateaus at $2/\eta$, $\lambda_{\max}$ stops increasing. If it moves, it only decreases from this time on.
+  • Fact 3:  Path-dependence.
+If changes to hyper parameters are made, Batch Sharpness changes abruptly or restart growing and $\lambda_{\max}$ also changes.
+Stabilization of both happen as Batch Sharpness reaches $2/\eta$.
+The trajectory of $\lambda_{\max}$ is inconsistent with the size of hyper parameters (see Figure 7).
+The level of $\lambda_{\max}$ is path-dependent: it inherits the history of progressive sharpening up to the moment EoSS is reached.
+  • Fact 4:  Smaller batches $\Rightarrow$ flatter minima.
+Across every setting we tested, reducing the batch size monotonically decreases the plateau level of $\lambda_{\max}$.
+This aligns with the long-standing empirical observation that smaller batches locate flatter minima see, e.g., Keskar et al.
+[2], Jastrzębski et al.
+[33]).
+  • Fact 5:  A critical batch size marks the SGD $\to$ GD crossover.
+Each curve in Figure 8 exhibits a bend at $b\approx b_{c}(\eta)$: for $b<b_{c}$ the plateau falls rapidly with $b$, while for $b\!>\!b_{c}$ it flattens and approaches the full-batch value.
+This $b_{c}$ corresponds to the regime in which the mini-batch landscapes approximate well enough the full-batch landscape, restoring GD-like dynamics (Appendix H).
+  • Fact 6:  No universal power law.
+From static analysis, one would expect a scaling $2/\eta-\lambda_{\max}=O(b^{-\alpha})$ for some $\alpha$, see Appendix I.
+The log–log plots (bottom row of Figure 8) show no robust straight-line behaviour, ruling out such law for any possible exponent $-\alpha$.
+6.2 Implications and Open Questions
+The findings above lead to the following main conclusions.
+  1. (C1)
+    $\lambda_{\max}$ is not the stability limiter for mini-batch training.
+Batch Sharpness governs EoSS; $\lambda_{\max}$ follows.
+$\lambda_{\max}$ is capped from above by the value it reaches at the phase transition characterized by Batch Sharpness reaching $2/\eta$.
+This and Facts 1—3 above imply that: {mdframed} The stabilization of $\lambda_{\max}$ is a by-product of EoSS, not the quantity that governs it.
+  2. (C2)
+    A theory of $\lambda_{\max}$ has to account for the correct progressive sharpening.
+By fixing the model and changing batch size $b$, the gap between the maximal eigenvalue of $\mathbb{E}\big{[}\lambda_{\max}(\mathcal{H}(L_{B}))\big{]}$ and $\lambda_{\max}=\lambda_{\max}(\mathcal{H})=\lambda_{\max}\mathbb{E}[\mathcal{H}(L_{B})]$ scales as $1/b$, see a proof in Appendix I.
+Any theory that keeps the parameter vector fixed and only varies $b$, or anyways leads to a power law, misses the path-dependent descent that determines where training arrives and where $\lambda_{\max}$ stabilizes.
+Facts 3 and 6 thus imply that analysis of $\lambda_{\max}$ is insufficient if it does not account for (1) the precise and correct effect of progressive sharpening on the higher moments of the Hessian and (2) the correct alignment between mini-batch steps and Hessians.
+Quantifying the plateau of $\lambda_{\max}$ is thus still an (important) open problem.
+A complete account will require a dynamical theory through the progressive-sharpening phase and beyond.
+Not just properties at its endpoint as for full-batch methods.
+Remark: Why not $\lambda_{\max}\mathbb{E}[\mathcal{H}(L_{B})]$.
+Figure 9: Ranging different batch sizes.
+One might hope that the largest eigenvalue of each mini-batch Hessian could act as a stochastic proxy for curvature.
+The reason is that the step is generally not aligned with the eigenvector of the largest eigenvalue and thus $\lambda_{\max}\mathbb{E}[\mathcal{H}(L_{B})]$ stabilizes right above $2/\eta$, however, the level of stabilization slightly changes with $b$—see Appendix J—unlike for Batch Sharpness.
+7 Implications: How Noise-Injected GD Differs from SGD
+SGD vs. Noisy Gradient Descent.
+Figure 10: SGD vs.
+Noisy GD.
+Only noise preserving the mini-batch structure of SGD leads to $\lambda_{\max}$ plateauing below $2/\eta$ (akin to EoSS and as observed by [2]).
+Noise injection fails to reproduce this behavior even with the same covariance SGD’s.
 A common belief is that SGD’s regularization stems from its “noisy” gradients, which find flatter minima.
@@
 To test this, we compare mini-batch SGD (batch size 16) against three noisy GD variants:
-  • Directional Noise: Gaussian noise matching the directions and signs of mini-batch SGD noise (Simsekli et al., 2019).
-  • Diagonal Noise: Gaussian noise restricted to the diagonal part of the SGD noise covariance.
-  • Isotropic Noise: Gaussian noise with isotropic covariance.
-As shown in Fig.
-7, only directionally accurate noise leads to an EoSS-like regime with $\lambda_{\max}$ stabilizing well below $2/\eta$.
-More generic (diagonal or isotropic) noise fails to reproduce this behavior, indicating that standard SDE approximations—which typically ignore directional correlations—cannot fully capture the observed dynamics.
-Figure 7: SGD vs.
-Noisy GD.
-Only noise preserving the covariance structure and direction of SGD leads to $\lambda_{\max}$ plateauing below $2/\eta$ (akin to EoSS).
-Diagonal or isotropic noise fails to reproduce this behavior.
-Note that batch sharpness is still plotted for comparison, though for noisy GD it loses its meaning.
-From a theoretical perspective (Appendix D), these experiments suggest that stability thresholds differ fundamentally between mini-batch SGD (governed by Batch Sharpness) and noise-injected GD (governed by $\lambda_{\max}$).
-6.2 The Mini-Batch Landscape Matters
-Classical analyses of neural network optimization often assume a single, static landscape: (i) Online perspective, modeling each sample’s gradient as a noisy unbiased estimator of the expected risk, or (ii) Offline perspective, treating the dataset as fixed and SGD as noisy GD on the empirical loss.
+  • Anisotropic Sampling Noise: Gaussian reweighting on the samples [51], which is different from SGD but maintains the mini-batch structure.
+  • Diagonal Noise: Gaussian noise restricted to the diagonal part of the SGD noise covariance [52].
+  • Isotropic Noise: Gaussian noise with isotropic covariance [52].
+As shown in Figure 10 and Appendix D, only noise which maintains the higher moments of the Hessian(s) leads to an EoSS-like regime with $\lambda_{\max}$ stabilizing well below $2/\eta$.
+More generic (e.g., diagonal or isotropic) noise fails to reproduce this behavior.
+These experiments suggest that stability thresholds differ fundamentally between mini-batch SGD (governed by Batch Sharpness) and noise-injected GD (governed by $\lambda_{\max}$).
+Unsurprisingly, in the case in which the noise affects only the gradients—not the Hessians—indeed, EoSS comes for $\lambda_{\max}=2/\eta$ as for GD [29, 32].
+Even in the quadratic setting, the appearance of Type-1 oscillations and GNI are not affected by the structure and distribution of the Hessian on the mini-batches, see Appendix A.
+The stability threshold, however, is.
+It depends on the Hessian’s higher moments, see Theorem 1 or [29, 32].
+Challenges for SDE Modeling.
+Classical analyses of neural network optimization often assume a single, static landscape: (i) Online perspective, modeling each step’s gradient as a noisy unbiased estimator of the expected gradient, or (ii) Offline perspective, treating the dataset as fixed and SGD as noisy GD on the empirical loss.
 In both views, it is the full-batch Hessian that supposedly drives curvature.
-Our results instead highlight the mini-batch landscape: each update sees a Hessian $\mathcal{H}(L_{B})$ that can differ significantly from $\mathcal{H}(L)$, leading to Batch Sharpness stabilizing at $2/\eta$ even when full-batch $\lambda_{\max}$ is much smaller.
-6.3 Challenges for SDE Modeling
-Prior works already note limitations of SDE-based approaches for SGD implicit regularization: they may be mathematically ill-posed (Yaida, 2018), fail except under restrictive conditions (Li et al., 2021), converge to qualitatively different minima (HaoChen et al., 2020), or miss higher-order effects (Damian et al., 2021; Li et al., 2022).
-Recent discrete analyses (Roberts, 2021; Smith et al., 2021; Beneventano, 2023) attempt to address some of these issues.
+Our results instead highlight that each update sees a Hessian $\mathcal{H}(L_{\mathbf{B}})$ that generally differs significantly from $\mathcal{H}$, leading to Batch Sharpness stabilizing at $2/\eta$ when $\lambda_{\max}$ is smaller.
+{mdframed}
+Standard SDE—or analogous—approximations of SGD cannot thus describe the location of convergence of SGD or its behavior for neural networks under the assumption of progressive sharpening.
+Indeed, they typically ignore any statistics of the Hessians except for the mean.
+Prior works already note limitations of SDE-based approaches for SGD implicit regularization: they may be mathematically ill-posed [53], fail except under restrictive conditions [54], converge to qualitatively different minima [55], or miss higher-order effects [56, 57].
+Recent discrete analyses [36, 37, 58] attempt to address some of these issues.
 Nonetheless, our findings expose a deeper gap: when batch sizes are small, the geometry of the mini-batch Hessian differs markedly from that of the full-batch, altering both eigenvalues and eigenvector alignments.
 Conventional SDE models, which assume a static or average Hessian, cannot easily capture these rapid fluctuations.
-Fully accounting for the distribution of $\{\mathbf{A}_{i}\}$ remains an open challenge, emphasizing the need for further research into truly batch-specific dynamics.
 8 Conclusions, Limitations, and Future Work
 Conclusions.
-We introduced Batch Sharpness as a key descriptor of mini-batch SGD dynamics, showing that it stabilizes near $2/\eta$.
-This defines an Edge of Stochastic Stability, analogous to the classic EoS (Xing et al., 2018; Cohen et al., 2021, 2022), but now in the small-batch regime most commonly used in practice.
-Empirically, Batch Sharpness remains at $2/\eta$ while the full-batch $\lambda_{\max}$ hovers below, at a level that systematically depends on batch size; indeed, the gap $\lambda_{\max}^{b}-\lambda_{\max}$ follows a power law $(\text{batch size})^{-0.7}$ across various models and datasets.
+We have addressed the longstanding question of if and how mini-batch SGD enters a regime reminiscent of the “Edge of Stability” previously observed in full-batch methods.
+Contrary to the usual focus on the global Hessian’s top eigenvalue, we uncovered that Batch Sharpness—the expected directional curvature of the mini-batch Hessian in the direction of its own gradient—consistently rises (progressive sharpening) and then hovers around $2/\eta$, independent of batch size.
+This behavior characterizes a new “Edge of Stochastic Stability”, which explains how mini-batch training can exhibit catapult-like surges and settle into flatter minima even when the full-batch Hessian remains below $2/\eta$.
+Our analysis clarifies why smaller batch sizes and larger learning rates both constrain the final curvature to a lower level, thereby linking these hyperparameters to flatter solutions and often improved generalization.
+Furthermore, we show that this phenomenon depends on the directional structure of mini-batch noise and does not straightforwardly emerge from generic isotropic or diagonal noise injections, highlighting important limitations of SDE-based approximations.
+Overall, the EoSS framework unifies several empirically observed effects—catapult phases, dependence on batch size, and progressive sharpening—under a single perspective focused on the mini-batch landscape and its directional curvature.
 Limitations.
 (i) We have tested only image-classification tasks, leaving open whether similar phenomena arise in NLP, RL, or other domains.
-(ii) Although we observe empirical batch-size scaling laws, we lack a full theoretical account, especially at moderate batch sizes.
+(ii) Although we observe empirical EoSS, we lack full theoretical account.
 (iii) Our experiments mainly use fixed learning rates and standard architectures, so very large-scale or large-batch settings remain less explored.
 (iv) We have not analyzed momentum-based or adaptive methods (e.g.
-Adam), even though full-batch EoS has been seen there (Cohen et al., 2022).
+Adam), even though full-batch EoS has been seen there [3].
 Future Work.
-Beyond addressing these limitations, several directions remain: (i) Extending higher-order analyses (Damian et al., 2023; Cohen et al., 2021; Ahn et al., 2022; Lyu et al., 2023) to mini-batch SGD may clarify why $\textit{Batch Sharpness}\sim 2/\eta$, (ii) Refining our batch-size scaling argument via random matrix theory might explain the “typical” exponent (around 0.7) and its possible crossovers for very large or very small batches.
-In summary, Batch Sharpness provides a more faithful perspective on SGD than the full-batch $\lambda_{\max}$.
-Our findings highlight how mini-batch Hessians drive training instabilities and ultimately shape learned solutions, hopefully inspiring further research on how step size and batch size jointly govern deep network training.
+Beyond addressing these limitations, several directions remain: Understanding (i) where $\lambda_{\max}$ stabilizes; (ii) how EoSS and EoS affect performances and the features learned by the neural network [1, 18, 21, 24]; (iii) consequently if it is benign effect or not; (iv) what the other sources of instability are there in the training; (v) better describing the phenomenon of progressive sharpening; and (vi) understanding its causes.
+Moreover, we conjecture, and experiments seem to suggest, that the phenomenon for SGD is very similar to the one described by [18, 26], but a mathematical verification is a necessary future work.
 Acknowledgement
-A special thanks to Stanisław Jastrzębski, Alex Damian, Jeremy Cohen, Afonso Bandeira, Boris Hanin, Mark Lowell, and Renée Carmona for invaluable discussions, which were crucial to the development of this project.
+A special thanks to Stanisław Jastrzębski, Alex Damian, Jeremy Cohen, Afonso Bandeira, Boris Hanin, and Renée Carmona for invaluable discussions, which were crucial to the development of this project.
+We also want to thank Mark Lowell, Yee Whye Teh, and the referees of ICML 2025, for comments and feedbacks on the first version of the preprint.
 We acknowledge the use of DeepSeek, Claude, and ChatGPT for providing code assistance, debugging support, and editing suggestions.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
-Appendix A On the Stability
-A.1 On Equivalent Formulations of Stability
-Consider as working example the quadratic Taylor approximation of a loss function:
-$f(x)=\frac{1}{2}x^{\top}\mathbf{A}x+\mathbf{b}^{\top}x+\mathbf{c},$
-where $\mathbf{A}$ is the Hessian matrix (symmetric and positive semidefinite), $\mathbf{b}$ is the gradient vector, and $\mathbf{c}$ is a constant.
-The question of stability is very nuanced, and we decided to define stability and $\eta_{\max}$ as follows
-Definition 2 (Stability).
-We say that the optimization algorithm $\mathcal{A}$ on the objective $f$ is stable if it does not diverge.
-By Folklore, in the quadratic objective $f$ above, enforcing stability is equivalent to choosing a step size smaller than $\eta_{\max}$ as follows.
-Definition 3 (Step size at the edge for an algorithm).
-Specify a discrete-time optimization algorithm of choice $\mathcal{A}$.
-We define "step size at the edge for $\mathcal{A}$" the maximum step size $\eta_{\max}(\mathcal{A})>0$ under which the dynamics do not diverge along the whole trajectory in the quadratic objective case, meaning there does not exist a $t$ such that $|x_{t+1}-x^{*}|>|x_{t}-x^{*}|$.
-In case of a stochastic update, we define $\eta_{\max}(\mathcal{A})$ analogously, but in expectation.
-This is a good definition in the quadratic setting above, no matter how we define divergence, indeed
-Lemma 1.
-Let $f(x)$ be quadratic, specify a discrete-time optimization algorithm of choice $\mathcal{A}$.
-Then $\eta_{\max}(\mathcal{A})>0$ is the maximum step size under which the dynamics does not diverge to infinity.
-Both in terms of $f(x_{t})$ and in terms of $\|x_{t}-x^{*}\|$.
-And we have, analogously, stability in terms of non-increasing loss.
-Lemma 2.
-Let $f(x)$ be quadratic, specify a discrete-time optimization algorithm of choice $\mathcal{A}$.
-Then $\eta_{\max}(\mathcal{A})>0$ is the maximum step size under which there does not exist a $t$ such that $f(x_{t+1})>f(x_{t})$.
-A.2 Specialization of the Setting to Mini-Batch SGD
-In the case of mini-batch SGD, we are working, instead, with the following 2-order Taylor approximation:
-$f_{B}(x)=\frac{1}{b}\sum_{i\in B}\left(\frac{1}{2}x^{\top}\mathbf{A}_{i}x+\mathbf{b}_{i}^{\top}x+\mathbf{c}_{i}\right),$ (6)
-where $B\subset\{1,2,\ldots,n\}$ is a randomly sampled mini batch of size $b$, and $\mathbf{A}_{i}$, $\mathbf{b}_{i}$, and $\mathbf{c}_{i}$ represent contributions from individual data points.
-The full objective matrices are instead given by:
-$\mathbf{A}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i},\quad\mathbf{b}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{b}_{i}.$
-Update rule for mini-batch SGD.
-Here, the mini-batch SGD update rule is:
-$x_{t+1}=x_{t}-\eta\nabla f_{B}(x_{t}),$
-where the gradient for mini batch $B$ is:
-$\nabla f_{B}(x_{t})\ =\ \frac{1}{b}\sum_{i\in B}\left(\mathbf{A}_{i}x_{t}+\mathbf{b}_{i}\right)\ =\ \left[\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}\right](x_{t}-x^{*})\,+\,\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x^{*}+\mathbf{b}_{i}\right).$ (7)
-Substituting this gradient, the update becomes:
-$x_{t+1}=x_{t}-\eta\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x_{t}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
-To analyze stability, one generally focuses on the evolution of the error $x_{t}-x^{*}$, where $x^{*}=-\mathbf{A}^{-1}\mathbf{b}$ is the optimum.
-By adding and subtracting $x^{*}$, we obtain that error evolves as:
-$x_{t+1}-x^{*}=\left(I-\eta\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}\right)\big{(}x_{t}-x^{*}\big{)}\ -\ \eta\left(\sum_{i\in B}\mathbf{A}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
-We can notice here the first difference with the full-batch case:
+Appendix A On the Two Types of Oscillations in SGD Dynamics
+A fundamental challenge in analyzing SGD compared to GD stems from the inherent oscillations induced by mini-batch gradient noise.
+This appendix, together with Appendix C, extends the discussion in Section 3 by formally distinguishing between two distinct types of oscillations: noise-driven (Type-1) and curvature-driven (Type-2).
+This distinction is crucial because Type-1 oscillations occur independently of the loss landscape’s curvature and thus do not exert a regularizing effect on the sharpness of the final solution.
+In contrast, Type-2 oscillations are directly caused by landscape curvature and induce an implicit regularization effect by discouraging convergence towards sharp minima.
+We begin with a minimalistic example to illustrate the nature of Type-1
+A.1 A Minimalistic Quadratic Example.
+Consider a regression problem with two datapoints, $1$ and $-1$, and a linear model $f(x)=x$ under the quadratic loss. The (scaled) full-batch loss is given by:
+$L(x)\;=\;\frac{1}{4}(x-1)^{2}\;+\;\frac{1}{4}(x+1)^{2}.$
+Batch-1 SGD updates with step-size $20<\eta<$ 2 result in oscillatory behavior around the optimum $x=0$ due entirely to gradient noise, with amplitude approximately $\sqrt{\frac{\eta}{2-\eta}}$.
+Crucially, the Hessian in this example is small ($\frac{d^{2}L}{dx^{2}}=1$), demonstrating that these persistent oscillations are entirely noise-driven (Type-1).
+Formally, the SGD update is:
+$x_{t+1}=x_{t}-\eta\nabla\ell_{i_{t}}=(1-\eta)x_{t}+\eta\xi_{t}$
+where $l_{i_{t}}$s are the individual datapoint losses, and $\xi_{t}$s are i.i.d Rademacher random variables. Thus, we obtain the first two moments explicitly:
+$\mathbb{E}[x_{t}]=(1-\eta)\mathbb{E}[x_{t-1}]=(1-\eta)^{t}x_{0}$
+$\mathbb{E}[x_{t}^{2}]=(1-\eta)^{2}\mathbb{E}[x_{t-1}^{2}]+\eta^{2}=(1-\eta)^{2t}x_{0}+\frac{\eta^{2}}{1-(1-\eta)^{2}}\Bigl{(}1-(1-\eta)^{2t}\Bigr{)}$
+This implies convergence in expectation for $20<\eta<$ 2, with a limiting variance given by:
+$\lim_{t\to\infty}\mathbb{E}[x_{t}^{2}]=\frac{\eta}{2-\eta}$
+and divergence for $\eta>2$.
+A key observation is that increasing $\eta$ to any value $\eta_{1}<2$ merely changes the amplitude of oscillations to $\sqrt{\frac{\eta_{1}}{2-\eta_{1}}}$ without triggering any catapult-like behavior.
+The only step size for which we observe Type-2 (curvature-driven) oscillations and an EoS-like [^8] The key difference between these oscillations and genuine EoS behavior in neural networks is that, in the quadratic case, the full-batch loss does not decrease, making this scenario inherently less informative.
+In contrast, neural networks exhibit a surprising, albeit non-monotonic, decrease in loss within this instability regime, an effect arising from the multidimensional nature of their optimization landscape [18] instability is precisely $\eta=2$, where the dynamics effectively become a random walk, and any larger step size leads to divergence.
+See Figure 3 for a plot of this phenomena for a one-dimensional example with many datapoints, with the same stability thresholds (also see Example 3.1 in Bottou et al. [48]).
+{mdframed}
+Crucially, when $\eta<2$ oscillations occur persistently on the full-batch loss, despite the individual steps on the mini-batch loss remaining stable.
+The oscillation is due to the fact that the mini-batch loss landscape shifts from step to step, not to the fact that the steps are unstable.
+A.2 General Case: From One-Dimensional Toy to Multidimensional Lyapunov Analysis
+The simple one-dimensional regression in §A.1 already demonstrates how noise-driven (Type-1) oscillations can persist indefinitely and yield a stable “two-cycle” around the optimum, independent of the actual Hessian magnitude.
+In higher dimensions, the story is similar: when the step size $\eta$ is fixed, the randomness in mini-batch gradients still injects a continual “kick” at each iteration, causing the iterates to hover in a noisy neighborhood of the minimum.
+The main difference is that now there can be many directions—some with higher curvature than others, or even flat ($\lambda=0$) directions.
+Nonetheless, the essential mechanism remains:
+$\Delta_{t+1}\;=\;\bigl{(}I-\eta\,\mathcal{H}\bigr{)}\;\Delta_{t}\;-\;\eta\,\xi_{t},$
+where $\Delta_{t}=x_{t}-x^{\star}$ is the displacement from the optimum, $\mathcal{H}$ is the Hessian at $x^{\star}$, and $\xi_{t}$ encodes the random fluctuation (gradient noise).
+Once $\Delta_{t}$ settles into a stationary distribution, the covariance $\Sigma_{x}$ can be found by solving a discrete Lyapunov equation similar to the one-dimensional case.
+Key References and the Road Ahead.
+A number of works formalize this “SGD noise equilibrium” by viewing the updates as a linear Markov chain in a neighborhood of $x^{\star}$.
+Classical references include Mandt et al.
+[47] for the stochastic differential analogy (Ornstein–Uhlenbeck process), and Bottou et al.
+[48] for a thorough discussion of how the constant stepsize prevents exact convergence.
+Intuitively, the argument for Proposition 1 goes thus as follows:
   1. 1.
-    The rest $\sum_{i\in B}\mathbf{A}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}$ it is different than 0 and it is different than 0 even in expectation when sampling is not uniform and independent.
+    The iterates oscillate with a stationary covariance $\Sigma_{x}$ around $x^{*}$.
   2. 2.
-    The multiplier $\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}$ is not $\mathbf{A}$.
-Next note that if we work with an full-batch GD with noise injected with unbiased and independent steps in the fashion of Robbins & Monro (1951), their expectation evens these terms out.
-We refer the reader to Appendix D for that.
-If we work with a biased algorithm, as Random Reshuffling (or SGD without replacement) or Shuffle Once, this is not the case (Mishchenko et al., 2020; Smith et al., 2021; Beneventano, 2023).
-We refer the reader to Appendix E for what changes in this setting.
-Furthermore, while in the deterministic case working with $x_{t+1}-x^{*}$ or with $\|x_{t+1}-x^{*}\|^{2}$ is the same, this is not the case for mini-batch algorithms.
-It is thus non-obvious how to generalize stability to start with.
-We will see in the next section how to do it.
-A.3 Formulating Stability for Mini-Batch SGD
-Unlike in the deterministic case, we have to be a little more careful with the formulation here.
-Let $(\lambda_{i},v_{i})$ be the ordered eigenvalue-eigenvector pairs of $\mathbf{A}$.
-With a deterministic Hessian matrix $\mathbf{A}$ we have that $x_{t+1}-x^{*}$ does not diverge if and only if the highest eigenvalue $\alpha_{i}$ such that $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ is smaller than $2/\eta$, see Cohen et al.
-(2021).
-The idea is that if a component $\langle x_{t+1}-x^{*},v_{i}\rangle\neq 0$ exists, after $k$ steps it gets multiplied by $(1-\eta\alpha_{i})^{k}$.
-This implies that the system will diverge if and only if $|1-\eta\alpha_{i}|>1$.
-When our step is a noise-injected GD similar stability phenomena happens, $\mathbf{A}$ is not impacted by the noise, see Appendix D.
-However, in the case of mini-batch SGD we have to be much more precise, and note that at every step, stability does not depend on $\mathbf{A}$, but on the "$\mathbf{A}$" of the mini batch.
-We have an important changes here:
-We do not have persistence of a Hessian matrix "$\mathbf{A}$".
-In particular, there is no persistence in the eigenspaces of the Hessian from step to step.
-This implies that we can not make an argument as the one above tracking the components of $x_{t}-x^{*}$.
-This will have an important implication, that at the step $t$ we care about the direction we are moving towards, not about the eigenvector of the highest eigenvalue.
-We are now dealing with objects of the form $\nabla f^{\top}\cdot\nabla^{2}f\cdot\nabla f$, not of the form $\lambda_{\max}\nabla^{2}f$.
-This imply that we may not really work on the eigenspaces of $\mathbf{A}$ as before, we decide to work with the analogous formulation of stability given by Lemma 2: For quadratic objectives, $|x_{t+1}-x^{*}|\leq|x_{t}-x^{*}|$ reduces in norm if and only if $f(x_{t+1})\leq f(x_{t})$.
-We will work from now on with the formulation of stability of $f(x_{t+1})\leq f(x_{t})$.
-At the step $t$ we can write the following thing update for a quadratic objective $f$.
+    Full-batch (expected) gradient is zero at $x^{*}$ and grows roughly linearly with distance for small deviations (by Taylor expansion $\nabla L(x)\approx\mathcal{H}(x-x^{*})$).
+So on average over the iterations we have
+    $\mathbb{E}_{k}[\lVert\nabla L(x_{k})\rVert^{2}]=\mathrm{Tr}\big{(}\mathcal{H}\Sigma_{x}\mathcal{H}).$
+  3. 3.
+    The stationary covariance of the gradients is $\Sigma_{g}$ satisfies:
+    $\Sigma_{x}\approx\frac{\eta}{2}(\mathcal{H}^{-1}\Sigma_{g}).$
+Putting all together, this implies that $\Sigma_{g}$ of the gradients is about $\frac{2}{\eta}\mathcal{H}^{-1}$ bigger than the full-batch $\nabla L\nabla L^{\top}$.
+Precisely the following quantity (where $L_{i}$ is the loss on the $i-th$ data point):
+$\frac{\mathbb{E}_{i}\big{[}\nabla L_{i}(\mathbf{x})^{\top}\mathcal{H}\nabla L_{i}(\mathbf{x})\big{]}}{\lVert\nabla L\rVert^{2}}\,=\,\frac{\mathrm{Tr}\big{(}\mathcal{H}\Sigma_{g}\big{)}}{\mathrm{Tr}\big{(}\mathcal{H}\Sigma_{x}\mathcal{H})},$ (3)
+and this can be rewritten as
 […]
-which is equivalent to
-$2\|\nabla f(x_{t})\|\quad\geq\quad\eta\nabla f(x_{t})^{\top}\cdot\nabla^{2}f(x_{t})\cdot\nabla f(x_{t}).$ (9)
-Thus the stability of Cohen et al. (2021), from this perspective, is equivalent to enforcing for all $t$ the inequality Equation (9), or analogously enforcing it in expectation.
-Here, for full-batch GD we could argue that if $\langle x_{t}-x^{*},v_{1}\rangle$ and $\eta\lambda_{1}>2$ but $\eta\lambda_{i}<2$ for all $i\in\{2,3,\ldots,\}$ then there exists a time $T$, logarithmic in the initial values of $x_{0}$, such that $f(x_{T+1})>f(x_{T})$.
-This time $T$ is defined by the components $2+$ being small enough and the first one big enough that the $|1-\eta\lambda_{1}|>1$ update on the first components outweigh the decrease in the others.
-Appendix B Stability for Mini-Batch SGD
-In the case of mini-batch SGD we have an update on the mini-batch loss.
-To get to the loss from the mini-batch loss, we can just take an expectation over the batch-sampling process.
-Thus, by taking the expectation at every algorithm step we can write the stability requirement as
-$\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[f_{B}(x_{t+1})-f_{B}(x_{t})\right]\ =\ -\eta\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\nabla f_{B}(x_{t})\right]\,+\,\frac{\eta^{2}}{2}\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\nabla^{2}f_{B}(x_{t})\nabla f_{B}(x_{t})\right]\ \leq\ 0.$ (10)
-This leads to our lemma
-Lemma 3.
-$\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]$ if and only if the step size $\eta$ is smaller than $2$ divided by
-Batch $\textit{Batch Sharpness}\ =\ \frac{\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\cdot\nabla^{2}f_{B}(x_{t})\cdot\nabla f_{B}(x_{t})\right]}{\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\|\nabla f_{B}(x_{t})^{2}\|\right]}.$
+This $2/\eta$ thus comes out of the only fact of oscillating and it is unrelated to the Hessian value.
+Moreover, EoS happens only in the eigenspace of the highest eigenvalue, Type-1 noise on the whole subspace spanned by the eigenspaces of the non-negative eigenvalues.
+On Stability
+For this reason, linear stability analyses of stochastic gradient descent and noise-injected gradient descent on quadratic objectives–originally explored by Wu et al.
+[28] and further developed by Ma and Ying [29], Wu et al.
+[31], Mulayoff and Michaeli [32]—explicitly exclude Type-1 oscillations by categorizing them as stable.
+Specifically, Ma and Ying [29] restrict their analysis to interpolating minima, where all individual gradients vanish, thus effectively eliminating noise-driven oscillations and isolating the curvature-driven (Type-2) scenario.
+Mulayoff and Michaeli [32] extends this to a more general class of minima by restricting the analysis to the orthogonal complement of the null space of the Hessian, and demonstrating that the noise-driven oscillations do not affect stability.
+Lee and Jang [17] empirically established that Gradient-Noise Interaction (GNI) consistently remains around $2/\eta$ throughout training.
+From the above, this implies that most training occurs in an oscillatory regime (at least Type-1)—see Figure 12 and Appendix G.
+In contrast, our study specifically investigates the emergence and implications of Type-2 oscillations, given their significant role in implicitly regularizing the loss landscape.
+From Informal to Formal.
+In the next subsection (§A.3), we present a general discrete Lyapunov proof of Lemma 1, allowing $\mathcal{H}\succeq 0$ to be possibly degenerate and not necessarily commuting with the noise covariance.
+The result is summarized in Proposition 2, showing rigorously that $\mathrm{GNI}\approx 2/\eta$ arises under any stable constant-stepsize mini-batch SGD orbit.
+This “$\tfrac{2}{\eta}$-law” is precisely the high-dimensional extension of the toy one-dimensional phenomenon above.
+{mdframed} In particular, we show here that the appearance of some quantity being $2/\eta$ means that the system is oscillating but does not mean in principle that the landscape or the curvature adapted to the hyper parameters.
+A.3 Proof of Proposition 1
+Setup and Notation.
+Let
+$L(x)\;=\;\tfrac{1}{n}\sum_{i=1}^{n}\ell_{i}(x)$
+be twice–continuously differentiable with a (possibly non–isolated) minimiser $x^{\star}$.
+Denote by $\,\mathcal{H}:=\nabla^{2}L(x^{\star})\succeq 0$ the positive–semidefinite Hessian at $x^{\star}$.
+Decompose the ambient space $\mathbb{R}^{d}$ into
+$\mathbb{R}^{d}\;=\;E_{+}\oplus E_{0},\quad E_{+}:=\operatorname{Im}(\mathcal{H}),\quad E_{0}:=\ker(\mathcal{H}),$
+with corresponding orthogonal projectors $P_{+},P_{0}$. Let $\mathcal{H}^{\dagger}$ denote the Moore–Penrose pseudoinverse of $\mathcal{H}$. It will be convenient to define the Kronecker–sum operator
+$\mathcal{K}:\mathbb{R}^{d\times d}\to\mathbb{R}^{d\times d},\quad\mathcal{K}(X)\;=\;\mathcal{H}\,X\;+\;X\,\mathcal{H}.$
+Assumptions Near $x^{\star}$.
+We work under the following assumptions in a neighborhood of $x^{\star}$.
+  1. (A1)
+    Local linearity.  Each $\ell_{i}$ is (locally) twice differentiable, and in a sufficiently small neighborhood of $x^{\star}$ we have
+    $\nabla\ell_{i}(x)\;=\;\nabla\ell_{i}\bigl{(}x^{\star}\bigr{)}\;+\;\mathcal{H}\,\bigl{(}x-x^{\star}\bigr{)}\;+\;\mathcal{O}\!\bigl{(}\|x-x^{\star}\|^{2}\bigr{)},\quad\mathcal{H}\;=\;\nabla^{2}L\bigl{(}x^{\star}\bigr{)}\succeq 0.$
+  2. (A2)
+    Finite, compatible noise.  Define
+[…]
+    We assume that in the flat subspace $\ker(\mathcal{H})$, there is no large random forcing. Formally, either
+[…]
+    or else $\|P_{0}\,\Sigma_{g}\,P_{0}\|\lesssim\eta$ so that the random walk does not blow up in flat directions.
+  3. (A3)
+    Stepsize stability.  Let
+    $\lambda_{\max}^{+}\;:=\;\max\,\bigl{\{}\lambda>0:\lambda\in\sigma(\mathcal{H})\bigr{\}}.$
+    We take a constant stepsize $\eta$ such that
+    $0\;<\;\eta\;<\;\frac{2}{\lambda_{\max}^{+}}.$
+    This guarantees mean–square stability on the subspace $\operatorname{Im}(\mathcal{H})$, since $\rho\bigl{(}I-\eta\,\mathcal{H}\bigr{)}\leq 1-\eta\,\lambda_{\min}^{+}$.
+Remark 1 (Remarks on the assumptions).
+Exact vs. Lipschitz Hessian (on (A1))
+When each $\ell_{i}$ is strictly quadratic, the local linearity $\nabla\ell_{i}(x)=\nabla\ell_{i}(x^{\star})+\mathcal{H}\,(x-x^{\star})$ holds exactly.
+In general, if $\nabla^{2}\ell_{i}$ is $L_{2}$–Lipschitz, the second–order expansion yields a small $\mathcal{O}(\|x-x^{*}\|^{2})$ remainder.
+For small $\eta$, typical SGD iterates remain within $\mathcal{O}(\sqrt{\eta}\,)$ of $x^{\star}$, causing only $\mathcal{O}(\eta^{2})$ corrections in the Lyapunov equation—less than the main $\mathcal{O}(\eta)$ term.
+Small drift in flat directions (on (A2))
+The condition $P_{0}\,\Sigma_{g}\,P_{0}=0$ can be relaxed to $\|P_{0}\,\Sigma_{g}\,P_{0}\|\leq\delta$.
+A discrete–Lyapunov analysis shows that $\Sigma_{x}$ remains finite provided $\delta=\mathcal{O}(\eta)$.
+Concretely, if $\|P_{0}\,\Sigma_{g}\,P_{0}\|\leq\tfrac{1}{2}\,b\,\lambda_{\min}^{+}\,\eta$, then the null–space covariance $\Sigma_{x}^{00}$ does not diverge, matching the same $\mathcal{O}(\eta)$ scale in $\mathrm{Im}(\mathcal{H})$.
+Near–critical slow mixing (on (A3))
+As $\eta\to 2/\lambda_{\max}^{+}$ from below, the mixing time $\tau_{\mathrm{mix}}\sim 1/(\eta\,\lambda_{\min}^{+})$ can become quite large.
+Reviewers may appreciate a remark that we assume $\eta\lambda_{\min}^{+}\ll 1$, so that the chain quickly enters stationarity and the covariance arguments hold.
+Mini–Batch SGD and Gradient Noise.
+For each iteration $t$, sample a mini–batch $B_{t}$ of size $b$ (either i.i.d. or well–shuffled from a finite dataset) and define
+$g_{t}\;:=\;\frac{1}{b}\sum_{\,i\in B_{t}}\nabla\ell_{i}\bigl{(}x_{t}\bigr{)},\qquad x_{t+1}\;:=\;x_{t}\;-\;\eta\,g_{t}.$
+Let
+$\Delta_{t}\;:=\;x_{t}-x^{\star},\quad\xi_{t}\;:=\;g_{t}-\nabla L(x^{\star}).$
+Then $\mathbb{E}[\xi_{t}]=0$ and $\mathbb{E}[\xi_{t}\,\xi_{t}^{\top}]=\tfrac{1}{b}\,\Sigma_{g}$.
+In particular, for with–replacement (i.i.d.) sampling we have $\mathbb{E}[\xi_{t}\mid x_{t}]=0$, so the cross–term $\mathbb{E}[\xi_{t}\,\Delta_{t}^{\top}]=0$.
+For without–replacement sampling, one can show $\mathbb{E}[\xi_{t}\,\Delta_{t}^{\top}]$ remains $\mathcal{O}(\eta)$ [36, 37, 49], hence appearing only at order $\mathcal{O}(\eta^{2})$ in the final covariance.
+Proposition 2 (General Gradient–Noise Interaction).
+Under the above Setup/Notation and assumptions (A1)(A3), consider the mini–batch SGD updates
+$x_{t+1}\;=\;x_{t}-\eta\,g_{t},\quad g_{t}\;=\;\tfrac{1}{b}\sum_{\,i\in B_{t}}\nabla\ell_{i}(x_{t}),\quad B_{t}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}_{b}.$
+Then the linearized error process $\Delta_{t}:=x_{t}-x^{\star}$ admits a unique stationary covariance $\Sigma_{x}$ on $E_{+}=\mathrm{Im}(\mathcal{H})$. In particular,
+$\Sigma_{x}\;=\;\frac{\eta}{b}\,\mathcal{K}^{\dagger}\bigl{(}\Sigma_{g}\bigr{)}\;+\;\mathcal{O}(\eta^{2}),\qquad\mathcal{K}(X)\;=\;\mathcal{H}\,X+X\,\mathcal{H},$ (5)
+and, under stationarity,
+[…]
+Furthermore, all hidden constants scale at most linearly in $\|\Sigma_{g}\|$ and inversely in $\lambda_{\min}^{+}$ and one can make the constant in $\mathcal{O}(\eta)$ explicit by Lipschitz bounds on $\mathcal{H}$.
+In particular, this implies Proposition 1
+$\mathrm{GNI}(\mathbf{x})\;:=\;\frac{\mathbb{E}_{B}\![\,\nabla L_{B}(\mathbf{x})^{\top}\,\mathcal{H}\,\nabla L_{B}(\mathbf{x})\bigr{]}}{\|\nabla L(\mathbf{x})\|^{2}}\;\approx\;\frac{2}{\eta}.$
+Proof of Proposition 2.
+Linearized Dynamics and Lyapunov Equation.
+By (A1) (or its Lipschitz extension) we have $\nabla L(x_{t})=\mathcal{H}\,\Delta_{t}+\mathcal{O}(\|\Delta_{t}\|^{2})$.
+Restricting to the leading linear term, the recursion is
+$\Delta_{t+1}\;=\;\bigl{(}I-\eta\,\mathcal{H}\bigr{)}\,\Delta_{t}\;-\;\eta\,\xi_{t}.$
+At stationarity, let $\Sigma_{x}:=\mathbb{E}[\Delta_{t}\,\Delta_{t}^{\top}]$. Taking outer products in the linear approximation gives the discrete Lyapunov equation
+$\Sigma_{x}\;=\;\bigl{(}I-\eta\,\mathcal{H}\bigr{)}\,\Sigma_{x}\,\bigl{(}I-\eta\,\mathcal{H}\bigr{)}^{\top}\;+\;\frac{\eta^{2}}{b}\,\Sigma_{g}\;+\;\mathcal{O}(\eta^{3}),$ (7)
+where higher–order terms come from the $\mathcal{O}(\|\Delta_{t}\|^{2})$ nonlinearity and possible small correlation $\mathbb{E}[\xi_{t}\,\Delta_{t}^{\top}]$.
+These can be shown to contribute only at order $\mathcal{O}(\eta^{2})$ or higher in $\Sigma_{x}$.
+Vectorization and Solving for $\Sigma_{x}$.  Define $\mathrm{vec}(\cdot)$ so that $\mathrm{vec}(A\,X\,B)=(B^{\top}\!\otimes A)\,\mathrm{vec}(X)$. Then (7) rewrites as
+$\Bigl{[}I\;-\;\bigl{(}I-\eta\,\mathcal{H}\bigr{)}\otimes\bigl{(}I-\eta\,\mathcal{H}\bigr{)}\Bigr{]}\,\mathrm{vec}(\Sigma_{x})\;=\;\frac{\eta^{2}}{b}\,\mathrm{vec}\bigl{(}\Sigma_{g}\bigr{)}\;+\;\mathcal{O}(\eta^{3}).$
+Since $(I-\eta\,\mathcal{H})^{\otimes 2}=I-\eta\,\mathcal{K}+\mathcal{O}(\eta^{2})$, the bracket equals $\eta\,\mathcal{K}+\mathcal{O}(\eta^{2})$.
+Restricting to $E_{+}\otimes E_{+}$, the operator $\mathcal{K}$ is positive–definite, so its Moore–Penrose inverse $\mathcal{K}^{\dagger}$ exists there (while $\mathcal{K}=0$ on null directions).
+Hence
+$\mathrm{vec}(\Sigma_{x})\;=\;\eta\,\mathcal{K}^{\dagger}\Bigl{(}\tfrac{1}{b}\,\mathrm{vec}(\Sigma_{g})\Bigr{)}\;+\;\mathcal{O}(\eta^{2}).$
+Reverting to matrix form proves the key statement $\Sigma_{x}=\tfrac{\eta}{b}\,\mathcal{K}^{\dagger}(\Sigma_{g})+\mathcal{O}(\eta^{2})$.
+Full–Batch vs. Mini–Batch Gradients.  Next, we compare the second–moment of $\nabla L(x_{t})$ versus $\nabla L_{B}(x_{t})$. By definition,
+$\nabla L_{B}(x_{t})\;=\;\xi_{t}\;+\;\nabla L(x_{t}).$
+Under i.i.d. sampling, $\xi_{t}$ is conditionally uncorrelated with $\Delta_{t}$, hence also with $\nabla L(x_{t})=\mathcal{H}\,\Delta_{t}$. We then get:
+$\mathbb{E}_{B}\bigl{[}\nabla L_{B}(x_{t})^{\top}\,\mathcal{H}\,\nabla L_{B}(x_{t})\bigr{]}\;=\;\tfrac{1}{b}\,\mathrm{tr}\bigl{(}\mathcal{H}\,\Sigma_{g}\bigr{)}\;+\;\mathrm{tr}\bigl{(}\mathcal{H}\,\Sigma_{x}\,\mathcal{H}\bigr{)}.$
+Meanwhile,
+$\mathbb{E}\bigl{[}\|\nabla L(x_{t})\|^{2}\bigr{]}\;=\;\mathrm{tr}\bigl{(}\mathcal{H}\,\Sigma_{x}\,\mathcal{H}\bigr{)}.$
+Substituting $\Sigma_{x}\approx\tfrac{\eta}{b}\,\mathcal{K}^{\dagger}(\Sigma_{g})$ makes both terms proportional to $\mathrm{tr}(\mathcal{H}\,\Sigma_{g})$, and one finds
+$\frac{\mathbb{E}_{B}\bigl{[}\nabla L_{B}^{\top}\,\mathcal{H}\,\nabla L_{B}\bigr{]}}{\mathbb{E}\bigl{[}\|\nabla L(x_{t})\|^{2}\bigr{]}}\;=\;\frac{2}{\eta}\,\Bigl{(}1\;+\;\mathcal{O}(\eta)\Bigr{)}.$
+This completes the proof of (6) and Proposition 2. ∎
+Remark 2 (Discussion and Constants).
+  • Explicit constants in $\mathcal{O}(\eta)$.
+If $\|\mathcal{H}\|\leq L_{2}$ and $\lambda_{\min}^{+}>0$, then $\|\mathcal{K}^{\dagger}\|\lesssim(\lambda_{\min}^{+})^{-1}$.
+Consequently, the $\mathcal{O}(\eta^{2})$ terms in (5) scale linearly with $\|\Sigma_{g}\|$ and at most quadratically with $(\lambda_{\min}^{+})^{-1}$.
+Concretely,
+    $\bigl{\|}\Sigma_{x}-\tfrac{\eta}{b}\,\mathcal{K}^{\dagger}(\Sigma_{g})\bigr{\|}_{2}\;\leq\;C_{\mathrm{nonlin}}\,\eta^{2}\,\|\Sigma_{g}\|_{2},\quad\text{where}\quad C_{\mathrm{nonlin}}\;=\;\frac{L_{2}}{\bigl{(}2\,\lambda_{\min}^{+}\bigr{)}^{2}},$
+    and thus $\|\Sigma_{x}\|_{2}\;\leq\;\tfrac{\eta}{2\,b\,\lambda_{\min}^{+}}\|\Sigma_{g}\|_{2}+\mathcal{O}(\eta^{2})$.
+  • Without–replacement sampling.
+For large $n$ or for well–shuffled datasets, the correlation introduced by sampling without replacement typically appears only in the cross–term $\mathbb{E}[\xi_{t}\,\Delta_{t}^{\top}]$ at order $\mathcal{O}(\tfrac{b}{n}\,\eta)$, which is again absorbed into $\mathcal{O}(\eta^{2})$ at the level of $\Sigma_{x}$.
+Hence all conclusions remain valid to leading order $\mathcal{O}(\eta)$.
+  • Flat directions.
+If $\|P_{0}\,\Sigma_{g}\,P_{0}\|$ is nonzero but $\lesssim\eta$, the same discrete Lyapunov analysis shows there is still a finite stationary $\Sigma_{x}$ with $\mathcal{O}(\eta)$ scale in $E_{+}$ and at most $\mathcal{O}(\eta)$ in $E_{0}$.
+If $\|P_{0}\,\Sigma_{g}\,P_{0}\|\gg\eta$, the iterates execute an unbounded random walk in the null space and the stationary covariance diverges in those directions.
+Remark 3.
+  1. 1.
+    Universality.
+The factor $2/\eta$ emerges without assuming $[\mathcal{H},\Sigma_{g}]=0$ or $\mathcal{H}\succ 0$; commutativity affects only the size of the $\mathcal{O}(\eta)$ remainder.
+Thus every stable constant–stepsize SGD orbit satisfies the Gradient–Noise Interaction (GNI) rule.
+  2. 2.
+    Flat directions.
+Condition $P_{0}\Sigma_{g}P_{0}=0$ in (A2) is essential: if violated, the iterates execute an uncontrolled random walk in $E_{0}$, for which no finite stationary covariance exists and the ratio is undefined.
+  3. 3.
+    Edge of Stability.
+Observing $\text{GNI}\approx 2/\eta$ in practice is therefore necessary but not sufficient for curvature–driven (“Edge-of-Stability”) dynamics.
+One must additionally check that the top Hessian eigenvalue $\lambda_{\max}^{+}$ itself approaches $2/\eta$.
+  4. 4.
+    Higher–order corrections.
+Retaining the $\eta^{2}(H\otimes H)$ term in the Neumann expansion of $(I-\eta\mathcal{H})^{\otimes 2}$ refines $\Sigma_{x}$ and hence the GNI ratio; see [47, 48] for explicit bounds.
+Discussion of Key Assumptions & Broader Context.
+(1) Local linearity and neighborhood size.
+We rely on the locally linear approximation $\nabla\ell_{i}(x)\approx\nabla\ell_{i}(x^{\star})+\mathcal{H}\,\bigl{(}x-x^{\star}\bigr{)}$ from Assumption (A1), which is exact in the strictly quadratic case and otherwise follows from a second-order Taylor expansion with Lipschitz Hessians.
+In practice, as long as the step size $\eta$ is small enough, the SGD iterates remain close to $x^{\star}$.
+Consequently, the higher-order ($\mathcal{O}(\|x-x^{\star}\|^{2})$) terms contribute only $\mathcal{O}(\eta^{2})$ to the discrete Lyapunov equation, negligible compared to our main $\mathcal{O}(\eta)$ term.
+A related subtlety is that our stationarity result is exact for the linearized process; transferring it to the fully nonlinear dynamics requires “small-noise ergodicity” arguments [e.g.
+59, 60], ensuring that for sufficiently small $\eta$, the nonlinear SGD inherits nearly the same stationary covariance.
+(2) Flat directions and step-size stability.
+Assumption (A2) states $P_{0}\,\Sigma_{g}\,P_{0}=0$ (or at least $\|P_{0}\,\Sigma_{g}\,P_{0}\|\lesssim\eta$) to prevent unbounded drift along $\ker(\mathcal{H})$.
+If there is too much noise in flat directions, the covariance $\Sigma_{x}$ diverges, and the notion of a stable “oscillation” no longer holds.
+Assumption (A3) with $\eta<2/\lambda_{\max}^{+}$ ensures $\Delta_{t}$ is mean-square stable in the curved subspace $E_{+}=\mathrm{Im}(\mathcal{H})$.
+In changing or “progressively sharpening” landscapes, one must still check that $\eta$ remains safely below $2/\lambda_{\max}^{+}$.
+Another practical concern is whether $\xi_{t}$ is truly independent of $x_{t}$.
+For i.i.d.
+mini-batches, the tower property $\mathbb{E}[\xi_{t}\mid x_{t}]=0$ justifies the cross-term in the Lyapunov approach; without-replacement sampling, one obtains a small correlation bounded by $\mathcal{O}(\tfrac{b}{n}\,\eta)$, still absorbed by our $\mathcal{O}(\eta^{2})$ remainder.
+(3) Universal ratio vs. Edge-of-Stability. A key outcome is that the ratio
+$\frac{\mathbb{E}_{B}\bigl{[}\nabla L_{B}(x)^{\top}\,\mathcal{H}\,\nabla L_{B}(x)\bigr{]}}{\|\nabla L(x)\|^{2}}\;\approx\;\frac{2}{\eta}$
+holds broadly once the system is in a stable constant-stepsize regime.
+However, by itself it does not guarantee that the top Hessian eigenvalue is exactly $2/\eta$.
+Both curvature-driven and noise-driven oscillations can produce the same measured ratio.
+Thus, to detect a genuine “Edge-of-Stability” (EoS), one must also verify that $\lambda_{\max}^{+}(\mathcal{H})\approx 2/\eta$.
+Otherwise, the $2/\eta$ ratio simply reflects noise-dominated plateaus.
+Historical note on the $2/\eta$ law and Lyapunov analysis.
+The observation that fixed–stepsize SGD does not converge to a point but rather equilibrates at a noise–controlled “temperature” originates in classical stochastic approximation.
+Already in the seminal paper of Robbins and Monro [46] a diminishing step was prescribed precisely to avoid this residual variance.
+In the neural–network community, Bottou [61] drew the first explicit analogy between a constant learning rate and the temperature of a thermodynamic system, noting that the algorithm must reach a stationary distribution whose spread scales with $\eta$.
+This qualitative picture foreshadows the
+$\frac{\mathbb{E}[\nabla L_{B}^{\top}\mathcal{H}\,\nabla L_{B}]}{\|\nabla L\|^{2}}\;\approx\;\frac{2}{\eta}$
+identity proved in Proposition 1.
+Control–theoretic formalization.
+A rigorous derivation of the stationary covariance appeared with the discrete–time Lyapunov treatment of Mandt et al.
+[47], who viewed constant–stepsize SGD on a local quadratic as a Markov chain converging to an Ornstein–Uhlenbeck law.
+Solving the Lyapunov equation yields $\Sigma_{x}=\frac{\eta}{b}\mathcal{K}^{\dagger}(\Sigma_{g})+\mathcal{O}(\eta^{2})$, and, after a short calculation, the gradient–noise interaction plateaus at $\approx 2/\eta$.
+For strongly convex objectives this covariance was independently derived in the comprehensive optimisation survey of Bottou et al.
+[48, Thm.
+4.6], who emphasised that the full–batch gradient cannot decay below a floor proportional to $\eta\,\mathrm{Tr}(H\Sigma_{g})$.
+The commutativity and invertibility constraints of these early analyses were removed by Yaida [53], who solved the exact discrete Lyapunov equation in the general non–normal, possibly degenerate case.
+His fluctuation–dissipation relation shows that even without $[\mathcal{H},\Sigma_{g}]=0$ (or $\mathcal{H}\succ 0$) the first–order term in $\eta$ still enforces the same $2/\eta$ ratio proved in our Lemma.
+A.4 Full- vs. mini-batch gradient norms
+Lemma 2 (Bias–variance identity.).
+For every $\theta\in\mathbb{R}^{d}$
+$\mathbb{E}_{B}\!\bigl{[}\|\nabla L_{B}\|^{2}\bigr{]}\;=\;\|\nabla L\|^{2}\;+\;\mathbb{E}_{B}\!\bigl{[}\|\nabla L_{B}-\nabla L\|^{2}\bigr{]}.$ (8)
 Proof.
-Note that the Taylor expansion as in Equation 10 ensures that
-$\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]\ =\ \mathbb{E}\left[-\eta\nabla f_{B}(x_{t})^{\top}\nabla f_{B}(x_{t})\,+\,\frac{\eta^{2}}{2}\nabla f_{B}(x_{t})^{\top}\cdot\nabla^{2}f_{B}(x_{t})\nabla f_{B}(x_{t})\right]$
-Thus, by enforcing $\mathbb{E}[f_{B}(x_{t+1})]\leq\mathbb{E}[f_{B}(x_{t})]$ is equivalent to enforcing
-$2\ \geq\ \eta\cdot\textit{Batch Sharpness}.$
-Taking this expectation over the ways in which we can sample the batches we conclude the proof. ∎
-Appendix C Comparison with Lee & Jang (2023)
-After we wrote the first version of this article, we realized that Lee & Jang (2023) already dealt with this problem.
-We draw here a comparison and show that they come to a different conclusion because they deal with a different notion of stability.
-While our batch sharpness and their Interaction-Aware Sharpness (IAS) are similar in the experiments–and about the same in case of bigger batch sizes, we discuss how our two notions of stability relate to the one of Cohen et al.
-(2021).
-They propose that the quantity which stabilizes around $2/\eta$ is, in our notation,
-$\mathrm{IAS}\quad:=\quad\frac{\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}^{\top}\cdot\nabla^{2}f\cdot\nabla f_{B}\right]}{\|\nabla f\|^{2}}\quad=\quad\frac{\mathrm{tr}\left(\nabla^{2}f\cdot\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}\nabla f_{B}^{\top}\right]\right)}{\mathrm{tr}\left(\nabla f\nabla f^{\top}\right)}.$ (11)
-The very natural way they come up with this quantity is by imposing the following inequality, similar to our Equation (10).
-$\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\underbrace{f(x_{t+1})-f(x_{t})}\right]\ =\ -\eta\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla\underbrace{f(x_{t})^{\top}}\nabla f_{B}(x_{t})\right]\,+\,\frac{\eta^{2}}{2}\mathbb{E}_{\begin{subarray}{c}\text{sampling}\\ \text{batches}\end{subarray}}\left[\nabla f_{B}(x_{t})^{\top}\cdot\underbrace{\nabla^{2}f(x_{t})}\nabla f_{B}(x_{t})\right]\ \leq\ 0.$ (12)
-The difference between our Equation (10) and this equation above if that the terms in which we put a under brace are here evaluated on the whole dataset and not on the mini batch considered.
-While the inequalities seem analogous and both equally motivated at first glance, we claim they carry an important difference.
-Enforcing Equation (12) does imply that in 1 step the loss does not go up.
-However, it does not imply that the loss will not go up eventually.
-The instability threshold discussed in Cohen et al.
-(2021), in the following work, and in our is stability about the whole trajectory, not about one precise step of the optimizer in consideration.
+Expand $\|\nabla L_{B}\|^{2}=\|\nabla L+(\nabla L_{B}-\nabla L)\|^{2}$ and take $\mathbb{E}_{B}[\cdot]$. The cross term vanishes because $\mathbb{E}_{B}[\nabla L_{B}]=\nabla L$. ∎
+Explicit variance.
+Let $\Sigma(\theta):=\mathrm{Cov}_{(x,y)\sim\mathcal{D}}[\,\nabla L_{B}(\theta)\,]$, $B\sim\mathcal{P}_{b}$, $b=1$, be the covariance of a single sample gradient. Then, for i.i.d. batches of size $b$,
+$\mathbb{E}_{B}\!\bigl{[}\|\nabla L_{B}-\nabla L\|^{2}\bigr{]}=\frac{1}{b}\,\operatorname{tr}\bigl{(}\Sigma(\theta)\bigr{)}.$
+For a finite dataset of size $n$ sampled without replacement the factor $1/b$ is replaced by ${(n-b)}/\bigl{(}{b\,(n-1)}\bigr{)}$.
+Stationary, small-step regime.
+In Appendix A.3 we solved the discrete Lyapunov equation and obtained (first order in the step size $\eta$):
+$\|\nabla L(\theta)\|^{2}=\frac{\eta}{b}\,\operatorname{tr}\!\bigl{(}\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H}\bigr{)}+\mathcal{O}(\eta^{2})$
+where $\mathcal{K}(X)=\mathcal{H}X+X\mathcal{H}$ and $\mathcal{K}^{\dagger}$ is its Moore–Penrose inverse on $\operatorname{Im}\mathcal{H}\otimes\operatorname{Im}\mathcal{H}$. This implies that
+$\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]}\quad=\quad\eta\,\frac{\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}{\operatorname{tr}(\Sigma)+\eta\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}+\mathcal{O}(\eta^{2}).$ (9)
+Spectral bounds on the bias–variance ratio.
+Let the eigen–decomposition of the full-batch Hessian be $\mathcal{H}=\sum_{i=1}^{r}\lambda_{i}\,v_{i}v_{i}^{\!\top}$ with strictly positive eigen-values $0<\mu:=\lambda_{1}\leq\dots\leq\lambda_{r}=\lambda_{\max}$ and let $P_{+}:=\sum_{i=1}^{r}v_{i}v_{i}^{\!\top}$ be the projector onto $\operatorname{Im}\mathcal{H}$.
+Define the curved-subspace covariance $\Sigma_{+}:=P_{+}\Sigma P_{+}$.
+Then, up to the $\mathcal{O}(\eta^{2})$ term already displayed,
+$\frac{\mu}{2}\,\operatorname{tr}(\Sigma_{+})\;\leq\;\operatorname{tr}\bigl{(}\mathcal{H}\,\mathcal{K}^{\dagger}(\Sigma)\,\mathcal{H}\bigr{)}\;\leq\;\frac{\lambda_{\max}}{2}\,\operatorname{tr}(\Sigma_{+})\;\leq\;\frac{\lambda_{\max}}{2}\,\operatorname{tr}(\Sigma).$ (10)
+Consequently
+Lemma 3 (Bound on the ratio).
+In the oscillatory regime, we have
+$\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]}\;=\;\eta\,\frac{\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}{\operatorname{tr}(\Sigma)+\eta\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})}+\mathcal{O}(\eta^{2})$
+and
+$\frac{\eta\,\mu}{2}\;\frac{\operatorname{tr}(\Sigma_{+})}{\operatorname{tr}(\Sigma)+\tfrac{\eta\lambda_{\max}}{2}\operatorname{tr}(\Sigma_{+})}\;\leq\;\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]}\;\leq\;\frac{\eta\,\lambda_{\max}}{2}\;\frac{\operatorname{tr}(\Sigma_{+})}{\operatorname{tr}(\Sigma)+\tfrac{\eta\mu}{2}\operatorname{tr}(\Sigma_{+})}.$
+If the gradient noise has no component in the null-space of $\mathcal{H}$ $(\Sigma_{+}=\Sigma)$,
+$\frac{\eta\mu}{2+\eta\mu}\;\leq\;\varepsilon^{2}(\theta)\;\leq\;\frac{\eta\lambda_{\max}}{2+\eta\lambda_{\max}},\qquad\frac{\|\nabla L\|^{2}}{\mathbb{E}_{B}[\|\nabla L_{B}\|^{2}]})\in\bigl{[}\tfrac{\eta\mu}{2},\tfrac{\eta\lambda_{\max}}{2}\bigr{]}+\mathcal{O}(\eta^{2}).$
+Proof.
+Write $\Sigma=\sum_{i,j}s_{ij}\,v_{i}v_{j}^{\!\top}$. Because $\mathcal{K}^{\dagger}(v_{i}v_{j}^{\!\top})=(\lambda_{i}+\lambda_{j})^{-1}v_{i}v_{j}^{\!\top}$ for $\lambda_{i}+\lambda_{j}>0$,
+$\mathcal{H}\,\mathcal{K}^{\dagger}(\Sigma)\,\mathcal{H}=\sum_{i,j}\frac{\lambda_{i}\lambda_{j}}{\lambda_{i}+\lambda_{j}}\,s_{ij}\;v_{i}v_{j}^{\!\top}.$
+Taking the trace removes all off-diagonal terms and gives
+$\operatorname{tr}\bigl{(}\mathcal{H}\,\mathcal{K}^{\dagger}(\Sigma)\,\mathcal{H}\bigr{)}=\sum_{i=1}^{r}\frac{\lambda_{i}^{2}}{2\lambda_{i}}\,s_{ii}=\frac{1}{2}\sum_{i=1}^{r}\lambda_{i}\,s_{ii}.$
+Each coefficient $\lambda_{i}$ lies between $\mu$ and $\lambda_{\max}$, while $\sum_{i}s_{ii}=\operatorname{tr}(\Sigma_{+})$.
+This yields the two inequalities in (10).
+Insert them into the fraction $\eta\,\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})/\bigl{[}\operatorname{tr}(\Sigma)+\eta\operatorname{tr}(\mathcal{H}\mathcal{K}^{\dagger}(\Sigma)\mathcal{H})\bigr{]}$ and simplify.
+∎
+A.5 Extending to Neural Networks
+Differentiating Oscillations in Neural Network Optimization
+Our analytical treatment of SGD on one-dimensional quadratic objectives leverages the simplicity of having a single curvature measure–the second derivative–which facilitates a precise landscape characterization and explicit stability conditions.
+However, extending this analysis to multidimensional quadratics already introduces significantly more intricate dynamics, necessitating advanced analytical frameworks as developed by [28, 29, 32].
+Transitioning further to neural network optimization increases this complexity dramatically, since training predominantly occurs away from the manifold of minima, including the EoS-like instabilities themselves (as evidenced by the continuous reduction in loss)—and therefore requires to go beyond linear stability of quadratics near the manifold of minima.
+Given the current absence of robust theoretical tools to comprehensively analyze such dynamics, distinguishing between curvature-driven and noise-driven oscillations necessitates empirical experimentation.
+Specifically, we probe the dynamics by systematically varying hyperparameters (e.g., step size or batch size), as illustrated in Figure 3, allowing us to differentiate curvature-induced (Type-2) oscillations from purely noise-induced (Type-1) oscillations (Figure 4).
+Type-2 Oscillations Are Unique to NN Optimization
+This complexity inherent in neural network optimization is not merely an analytical inconvenience; rather, it is intrinsically tied to the emergence and significance of Type-2 oscillations and EoS-style phenomena.
+Notably, Type-2 oscillations emerge naturally [^10] We define an as emerging naturally if it arises inherently from the training dynamics, and not a result of precisely-selected hyperparameters or initializations, reflecting a fundamental characteristic of the optimization process itself.
+Formally, it needs to happen over a range of hyperparameter choices and initializations.
+only in the case of neural network optimization, but not in the case of quadratic objectives.
+In the one-dimensional quadratic scenario analyzed previously, curvature-driven oscillations require the step size to precisely match the stability threshold $2/\lambda_{\max}$, or exceed it, in which case we have divergence—in either case, it means that optimization of quadratics does naturally enter a regime of instability.
+In contrast, neural network optimization uniquely exhibits progressive sharpening, a third-order derivative phenomenon [18], where curvature naturally increases during training.
+This progressive increase in curvature means that training with a fixed step size can transition into an EoS-like regime of instability without any explicit adjustment of the hyperparameters, and stay there due to self-stabilization effects [18].
+Hence, Type-2 oscillations emerge naturally and robustly within neural network training dynamics due to this intrinsic change of the loss landscape.
+Consequently, Type-2 oscillations and EoS-like regimes are fundamentally driven by progressive sharpening, which doesn’t happen in quadratics, making it a purely neural network optimization phenomena.
+Appendix B Quadratic Setting: Batch Sharpness and GNI
+B.1 Setting: Data and risk.
+Intuitively, we minimise a random quadratic loss whose curvature changes with each sample. Let $(H_{i},x_{i})_{i\geq 1}$ be i.i.d. with
+$H_{i}\in\mathbb{R}^{d\times d},\;H_{i}\succeq 0,\quad x_{i}\in\mathbb{R}^{d},\quad\mathbb{E}\!\bigl{[}\lVert H_{i}\rVert_{\mathrm{F}}^{4}+\left\|x_{i}\right\|^{4}\bigr{]}<\infty.$
+Define [^11] All eigenvalues of $H:=\mathbb{E}[H_{i}]$ are positive because $\lambda_{\min}(H)>0$ is assumed; the ‘$+$’ superscript in earlier drafts is therefore redundant.
+$H:=\mathbb{E}[H_{i}],\quad G:=\mathbb{E}[H_{i}x_{i}],\quad\theta^{\star}:=H^{-1}G.$
+The population risk is $L(\theta):=\mathbb{E}_{i}\bigl{[}(\theta-x_{i})^{\top}H_{i}(\theta-x_{i})\bigr{]}$.
+B.2 SGD dynamics.
+With constant stepsize $\eta$ satisfying
+$0<\eta<\frac{2}{\lambda_{\max}(H)},$ (11)
+the update
+$\theta_{t+1}=\theta_{t}-2\eta\,H_{i_{t}}\bigl{(}\theta_{t}-x_{i_{t}}\bigr{)},$ (12)
+where $i_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}\{1,\dots,n\}$ and $i_{t}\!\perp\!\!\!\perp\!\theta_{t}$, generates a Markov chain that is geometrically ergodic [60].
+Let $\pi_{\eta}$ denote its unique stationary law and write $m_{\eta}:=\mathbb{E}_{\pi_{\eta}}[\theta]$, $\Sigma_{\eta}:=\operatorname{Cov}_{\pi_{\eta}}(\theta)$.
+Key intuition: the stochastic gradient $g(\theta,i):=2H_{i}(\theta-x_{i})$ is unbiased, $\mathbb{E}_{i}[g(\theta,i)]=\nabla L(\theta)=2H(\theta-\theta^{\star})$, so the mean iterate should coincide with $\theta^{\star}$.
+B.3 Stationary mean and covariance.
+The discussion in Appendix A implies that in this setting
+Lemma 4 (Lyapunov solution).
+For any stepsize satisfying (11),
+$m_{\eta}$ $=\theta^{\star},$ (13)
+$\Sigma_{\eta}$ $=\frac{\eta}{2}\,H^{-1}\Sigma_{g}H^{-1}+\mathcal{O}(\eta^{2}),\qquad\Sigma_{g}:=4\,\operatorname{Var}_{i}\!\bigl{[}H_{i}(\theta^{\star}-x_{i})\bigr{]},$ (14)
+where the $\mathcal{O}(\eta^{2})$ constant depends polynomially on $\lambda_{\min}(H)^{-1}$ and on the $4^{\text{th}}$-moment bound above.
+Sketch.
+Set $\Delta_{t}:=\theta_{t}-\theta^{\star}$ and $\xi_{t}:=2H_{i_{t}}(\theta^{\star}-x_{i_{t}})$ (mean 00, variance $\Sigma_{g}$). Subtracting (12) at $\theta^{\star}$ gives
+$\Delta_{t+1}=(I-2\eta H)\Delta_{t}-\eta\,\xi_{t}-2\eta\bigl{(}H_{i_{t}}-H\bigr{)}\Delta_{t}.$ (15)
+Mean.
+Taking expectations and using $i_{t}\!\perp\!\!\!\perp\!\theta_{t}$, $\mathbb{E}[\Delta_{t+1}]=(I-2\eta H)\,\mathbb{E}[\Delta_{t}],$ whose unique fixed point is 00, proving $m_{\eta}=\theta^{\star}$ (13).
+Covariance.
+Let $S_{t}:=\mathbb{E}[\Delta_{t}\Delta_{t}^{\top}]$.
+Multiplying (15) by its transpose and exploiting $\mathbb{E}[\Delta_{t}]=0$, one obtains the discrete Lyapunov recursion $S_{t+1}=(I-2\eta H)S_{t}(I-2\eta H)^{\top}+\eta^{2}\Sigma_{g}+R_{t}$ with $\|R_{t}\|=O(\eta^{3})$ thanks to $\mathbb{E}\|H_{i}\|_{\mathrm{F}}^{4}<\infty$.
+Passing to the limit and inverting $\bigl{[}I-(I-2\eta H)^{\otimes 2}\bigr{]}$ by a Neumann series yields (14).
+∎
+Take-away.
+Constant-stepsize SGD is unbiased in the mean but has an $\mathcal{O}(\eta)$ stationary variance that blows up if $\eta$ is too large; the $2/\lambda_{\max}(H)$ is thus necessary for avoiding the blowup, although not sufficient.
+B.4 Per-sample residuals.
+Define, for any $\theta\in\mathbb{R}^{d}$,
+$Y_{i}(\theta):=H_{i}(\theta-x_{i}),\quad\mu(\theta):=\mathbb{E}_{i}[Y_{i}(\theta)]=H\theta-G.$
+At stationarity write $Y:=Y_{i}(\theta)$, $\mu:=\mu(\theta)$, and set
+$\widetilde{Y}\;:=\;Y-\mathbb{E}_{i}[Y\mid\theta]\;=\;Y-\mu\quad\Longrightarrow\quad\mathbb{E}_{i}[\widetilde{Y}\mid\theta]=0.$
+Finally define the scalars [^12] Notation: $\mathbb{E}_{i,\pi}$ integrates $i$ and $\theta\sim\pi_{\eta}$.
+$\mathcal{A}$ $:=\mathbb{E}_{i,\pi}\bigl{[}Y^{\!\top}H_{i}Y_{i}\bigr{]},$$\quad\mathcal{B}$ $:=\mathbb{E}_{i,\pi}\bigl{[}Y^{\!\top}HY_{i}\bigr{]},$ $\quad\mathcal{C}$ $:=\mathbb{E}_{i,\pi}\bigl{[}\left\|Y_{i}\right\|^{2}\bigr{]},$ (16)
+$C_{0}$ $:=\mathbb{E}_{i,\pi}\bigl{[}\left\|\widetilde{Y}_{i}\right\|^{2}\bigr{]},$ $\quad\mathcal{D}$ $:=\mu^{\top}\mathbb{E}_{i,\pi}\bigl{[}H_{i}Y_{i}\bigr{]}.$
+Why these symbols? $\mathcal{A}$ is the average loss at stationarity; $\mathcal{B}$ and $\mathcal{C}$ act as mixed-moment controls that will upper-bound $\mathcal{A}$ via a trace inequality.
+B.5 Bounding curvature–fluctuation interaction.
+Lemma 5 (Cross term).
+Let $\widetilde{\Delta}:=\mathbb{E}_{i,\pi}\bigl{[}\widetilde{Y}^{\!\top}(H_{i}-H)\widetilde{Y}\bigr{]}$ and $\Delta:=\mathcal{A}-\mathcal{B}=\mathbb{E}_{i,\pi}\bigl{[}Y^{\!\top}(H_{i}-H)Y\bigr{]}$.
+Then $\Delta:=\mathcal{A}-\mathcal{B}=\widetilde{\Delta}+2\mathcal{D}$,
+$\widetilde{\Delta}\;\leq\;\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]},$ (17)
+and
+$\Delta\;\leq\;\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}.$ (18)
+Proof.
+In the following two steps we establish the first inequality.
+Step 1 (Cauchy - Schwarz). For any matrix $A$ and vector $v$, $|v^{\top}Av|\leq\|A\|_{2}\,\|v\|^{2}$. With $A:=H_{i}-H$ and $v:=Y$:
+$|\,\widetilde{Y}^{\!\top}(H_{i}-H)\widetilde{Y}|\leq\|H_{i}-H\|_{2}\,\|\widetilde{Y}\|^{2}.$
+Note that this is sharp if and only if $\widetilde{Y}$ aligns with the eigenvectors of the maximal eigenvalues of $H_{i}-H$.
+Step 2 (Average and separate: Jensen). Independence of $i$ and $\theta$ at stationarity implies, taking the expectation that
+$|\Delta|\leq\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]}.$
+Second inequality. Note that
 […]
-$\begin{split}2/\eta\ \geq\ \mathrm{IAS}\ &=\ \frac{1}{\|\nabla f\|^{2}}\left(\nabla f^{\top}\cdot\nabla^{2}f\cdot\nabla f\mathbb{E}\left[\zeta_{B}^{\top}\cdot\nabla^{2}f\cdot\zeta_{B}\right]\right)\\ &=\ \frac{1}{\|\nabla f\|^{2}}\sum_{i=1}^{d}\lambda_{i}\,\big{(}\,\langle\nabla f,v_{i}\rangle^{2}\,+\,\mathbb{E}[\langle\zeta_{B},v_{i}\rangle^{2}]\,\big{)}.\end{split}$ (13)
-However, after $T$ steps we can lower bound the loss with an increasing sequence which eventually diverges.
-Moreover, the loss asymptotically will approach the sequence with a linear convergence rate.
-Precisely, define the noise $\zeta_{B}$ such that $\nabla f_{B}=\nabla f+\zeta_{B}$.
-Then, for instance, if $\nabla f$ of the noise $\zeta_{B}$ have scalar product different than 0 with $v_{d}$, enforcing Equation (12) of Lee & Jang (2023) we may have decrease in loss in one step, indeed
-$\begin{split}f(x_{T})-f(x_{0})\ &=\ \sum_{t=0}^{T-1}\left(\frac{\eta^{2}}{2}\sum_{i=1}^{d}\lambda_{i}\,\big{(}\,\langle\nabla f(x_{t}),v_{i}\rangle^{2}\,+\,\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{i}\rangle^{2}]\,\big{)}-\|\nabla f(x_{t})\|^{2}\right)\\ &\geq\ \sum_{t=0}^{T-1}\underbrace{\left(\frac{\eta^{2}}{2}\lambda_{1}-\eta\right)\,\langle\nabla f(x_{t}),v_{1}\rangle^{2}}_{positive}\,+\,\underbrace{\frac{\eta^{2}}{2}\mathbb{E}[\langle\zeta_{B}(x_{t}),v_{1}\rangle^{2}]}_{positive}.\end{split}$ (14)
-Both the quantities under-braced above–the parallel component to $v_{1}$–explode with the time $t$.
-This implies, in particular, that imposing $IAS\leq 2/\eta$ may implies one-step stability, but not trajectory stability.
-Our quantity, in turn, factors in the variability that the eigenspaces have on different batches.
-Our quantity is also strictly bigger, thus imposing it to be equal to $2/\eta$ induces a lower sharpness, see, e.g, Appendices G and I.
-Appendix D Stability for Noise-Injected Gradients Descents
-We will see here that what we showed above is the case of mini-batch algorithms but not of many other versions of stochastic algorithms considered in theoretical work or practice.
-We deal here with the case of noisy gradients, which is the case usually analyzed in the practice of theory works, see Section 6.
-While in minibatch stochastic gradient descent (SGD), the step function is
-$x_{t+1}=x_{t}-\eta\left(\frac{1}{b}\sum_{i\in B}\mathbf{A}_{i}x_{t}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right).$
-and the noise comes from mini batching as shown above, in noisy gradient descents we do not care of a possible way to rewrite $A,b,c$ as sum of other terms.
-Just, there exists a random variable $\zeta_{t}$, independent for different times $t_{1}\neq t_{2}$, such that the $t-th$ step is
-$x_{t+1}=x_{t}-\eta\left(\mathbf{A}\mathbf{(x_{t}+\zeta_{t})}+b\right).$
-This implies that
-$x_{t+1}-x^{*}=\left(I+\eta\mathbf{A}\right)\big{(}x_{t}-x^{*}\big{)}\ +\ A\zeta_{t}.$
-This imply that the Hessian does not change from step to step.
-The expected stability formulation is actually the one of Lee & Jang (2023) in the sense of Cohen et al.
-(2021) is obtained for $\eta_{\max}=2/\lambda_{\max}(\mathbf{A})$, as discussed above in Appendix C.
-This, in particular, establishes mathematically and agrees with what we see empirically in Section 6.
-Appendix E Sampling the Mini Batches
-For simplicity let us work with the one dimensional case.
-The point in this section is that In the one dimensional case where $\mathbf{A}_{i}=\mathbf{a}_{i},\mathbf{b}_{i},\mathbf{c}_{i}\in\mathbb{R}$ for all $i\in\{1,2,\ldots,n\}$ we have that
-$\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t+1}-x^{*}]=\underbrace{\mathbb{E}\left[\left(1-\eta\frac{1}{b}\sum_{i\in B}\mathbf{a}_{i}\right)\big{(}x_{t}-x^{*}\big{)}\right]}_{(1-\eta\mathbf{a})\cdot\mathbb{E}[x_{t+1}-x^{*}]\text{ when independent steps}}\ -\ \underbrace{\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}\left[\frac{1}{b}\sum_{i\in B}\mathbf{a}_{i}x^{*}+\frac{1}{b}\sum_{i\in B}\mathbf{b}_{i}\right]}_{0\text{ when uniform sampling}}.$ (15)
-When $a_{i},b_{i}$ are sampled uniformly within the dataset and independently of the step $t$, then this becomes
-$\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t+1}-x^{*}]=(1-\eta\mathbf{a})\mathbb{E}_{\begin{subarray}{c}\text{sampling B}\end{subarray}}[x_{t}-x^{*}]$ (16)
-and the stability threshold is the same as for full-batch GD or any other momentum version, for instance.
-In SGD without replacement the "independent steps" are now the epochs, we thus work wlog with an epoch at once exactly in the fashion of Beneventano (2023) as the tools previously developed by Mishchenko et al.
-(2020); Smith et al.
-(2021) only work for infinitesimal step size and the tools in Appendix E of Beneventano (2023) are the only ones which allow us to consider more products and big learning rate.
-Let $k=n/b$ be the number of batches in an epoch and call $B_{j}$ the batch of $j-th$ step and $\mathbf{a}_{B_{j}},\mathbf{b}_{B_{j}}$ the averages of $\mathbf{a}$ and $\mathbf{b}$ over the batch $B_{j}$.
-$x_{k}-x^{*}=\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\big{(}x_{t}-x^{*}\big{)}\ -\ \sum_{j=1}^{k}\left[\prod_{h=j+1}^{k}\left(1-\eta\mathbf{a}_{B_{h}}\right)\right]\big{(}\mathbf{a}_{B_{j}}x^{*}+\mathbf{b}_{B_{j}}\big{)}.$ (17)
-Next note that by taking the expectation of the term $\prod_{j=1}^{k}\left(1+\eta\mathbf{a}_{B_{j}}\right)$ we obtain
-$\mathbb{E}\left[\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\right]\ =\ (1-\eta\mathbf{a})^{k}\ +\ \frac{\eta^{2}}{n-1}\binom{k}{2}\left[\frac{1}{n}\sum_{i=1}^{n}a_{i}^{2}-a^{2}\right]\ +\ \ldots$
-Here if $\eta=O\left(\frac{b^{2}}{n}[\mathrm{Var}(\mathbf{a}_{1})]^{-1}\right)$ which is generally very small, the second them may already be of size $O(1)$.
-We already see that in this case $\eta$ of the instability threshold has to be set also through the variance of the $\mathbf{a}_{i}$, the size of the batch or the number of batches.
-More generally, by applying the results in Appendix E of Beneventano (2023) we obtain that
-$\mathbb{E}\left[\prod_{j=1}^{k}\left(1-\eta\mathbf{a}_{B_{j}}\right)\right]\ =\ (1-\eta\mathbf{a})^{k}\ -\ \kappa_{k}\big{(}1-\eta\mathbf{a}_{i}\big{)}$ (18)
-and analogously
-$\mathbb{E}\left[\sum_{j=1}^{k}\left[\prod_{h=j+1}^{k}\left(1+\eta\mathbf{a}_{B_{h}}\right)\right]\big{(}\mathbf{a}_{B_{j}}x^{*}+\mathbf{b}_{B_{j}}\big{)}\right]\ =\ -\ \sum_{j=1}^{k}\kappa_{j}\big{(}1+\eta\mathbf{a}_{k},1+\eta\mathbf{a}_{k-1},\ldots,1+\eta\mathbf{a}_{j+1},\mathbf{a}_{j}x^{*}+\mathbf{b}_{j}\big{)}$ (19)
-where $\kappa_{k}\big{(}X\big{)}$ is the $k-th$ cumulant of the random variable $X$, which here has an empirical distribution from which we sample the $\mathbf{a}_{i}$s.
-This implies that the fact of sampling without replacement already implies different $\eta$ for the edge of stability. We are now ready to deal with the main case.
-Appendix F On Largest Eigenvalues of Sums of Matrices
+Note that the fact that $\mu^{\top}(H_{i}-H)\mu$ is centered in zero, implies that
+[…]
+This establish the first thesis of the lemma and implies that
+$|Y^{\!\top}(H_{i}-H)Y|\;\leq\;\|H_{i}-H\|_{2}\,\left(\|\widetilde{Y}\|^{2}+2|\mu^{\top}\widetilde{Y}|+\lVert\mu\rVert^{2}\right).$
+Taking the expectation we have
+$|\Delta|\leq\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2\lVert\mu\rVert_{2}^{2}\cdot\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{2}]+\lVert\mu\rVert_{2}^{4}}.$
+This concludes the proof of the lemma. ∎
+$\Delta$ measures how curvature noise $(H_{i}-H)$ correlates with noise $Y$. Lemma 5 shows this interaction cannot exceed the root product of the moments of two natural quantities.
+B.6 Main trace inequality
+Proposition 3.
+We have that Batch Sharpness is smaller than GNI if and only if
+$\mathcal{B}C_{0}\;\geq\;\lVert\mu\rVert_{2}^{2}\cdot(\tilde{\Delta}+2\mathcal{D}).$ (21)
+In particular,
+$\mathcal{B}\,\mathcal{C}-\mathcal{A}\,\|\mu\|^{2}\;\geq\;\mathcal{B}C_{0}-\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}\lVert\mu\rVert^{2}.$ (22)
+This implies that Batch Sharpness is smaller than GNI up to $O(\eta)$ at the stationary distribution when
+$C_{0}\geq\frac{\eta}{2}\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}.$ (23)
+Proof.
+Note that $\mathcal{A}\lVert\mu\rVert^{2}\leq\mathcal{B}\lVert\mu\rVert^{2}+|\Delta|\lVert\mu\rVert^{2}$.
+Applying Lemma 5 and noticing that $\mathcal{C}=C_{0}+\lVert\mu\rVert^{2}$ concludes the proof of (22).
+Next note that up to $\mathcal{O}(\eta)$ we have that GNI $=\mathcal{B}/\lVert\mu\rVert^{2}=2/\eta$ at the stationary distribution.
+This proves the second part (23) and establishes Proposition 3.
+∎
+B.7 The moments of the oscillations
+1. Centering first simplifies the ratio
+Because $C_{0}$ is defined with the centred residual $\widetilde{Y}$, we compare
+$\frac{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}{C_{0}^{2}}\;=\;\frac{\mathbb{E}[\|\widetilde{Y}\|^{4}]}{C_{0}^{2}}+2\,\frac{\|\mu\|^{2}}{C_{0}}+\frac{\|\mu\|^{4}}{C_{0}^{2}}.$
+The first term is the kurtosis-type ratio $\kappa_{Y}\;:=\;\mathbb{E}[\|\widetilde{Y}\|^{4}]/C_{0}^{2}.$
+2. Exact formula in the Gaussian quadratic setting
+Assume:
+  • $x_{i}\!\mid\!H_{i}\sim\mathcal{N}(\theta^{\star},\Sigma_{x})$ – an isotropic sampling model for simplicity;
+  • $H_{i}$ is deterministic, i.e. $H_{i}\equiv H\succeq 0$.
+Then $\xi:=Y=H(\theta^{\star}-x_{i})\sim\mathcal{N}\!\bigl{(}0,\;S\bigr{)}$ with $S:=H\Sigma_{x}H$ and $\widetilde{Y}\equiv Y$ because $\mu=0$. By Isserlis’ theorem
+$\mathbb{E}[\|Y\|^{4}]\;=\;(\operatorname{tr}S)^{2}+2\operatorname{tr}(S^{2}),\qquad C_{0}\;=\;\operatorname{tr}S.$
+Therefore
+[…]
+  • Isotropic case $S=\sigma^{2}I_{d}$Longrightarrow $\kappa=1+\tfrac{2}{d}$, e.g. $=3$ in $d=1$, $=1.4$ in $d=5$, $\to 1$ for large $d$.
+  • Rank-one noise $S=\sigma^{2}uu^{\!\top}$ ($\|u\|=1$$\Longrightarrow$ $\kappa=3$ – the maximal value compatible with Gaussianity.
+  • The interval bounds follow from $\operatorname{tr}(S^{2})\in[\tfrac{1}{d}(\operatorname{tr}S)^{2},\,(\operatorname{tr}S)^{2}]$.
+3. Effect of random curvature $H_{i}$
+If $H_{i}$ varies but is independent of $x_{i}$ and of $\theta$ a priori, the second moment becomes
+$C_{0}\;=\;\mathbb{E}_{H}\!\bigl{[}\operatorname{tr}(H\Sigma_{x}H)\bigr{]},$
+while
+$\mathbb{E}[\|Y\|^{4}]\;=\;\mathbb{E}_{H}\Bigl{[}(\operatorname{tr}(H\Sigma_{x}H))^{2}+2\,\operatorname{tr}\!\bigl{(}(H\Sigma_{x}H)^{2}\bigr{)}\Bigr{]}.$
+Define $a:=\mathbb{E}_{H}[\operatorname{tr}(H\Sigma_{x}H)],\;b:=\mathbb{E}_{H}[\operatorname{tr}(H\Sigma_{x}H)^{2}],\;c:=\mathbb{E}_{H}[\operatorname{tr}((H\Sigma_{x}H)^{2})].$ Then
+[…]
+[…]
+B.8 Closing up
+In the setting above, thus, we can transform (23) in the following inequality
+$1\geq\frac{\eta}{2}\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\sqrt{\frac{\mathbb{E}_{\pi}[\|\widetilde{Y}\|^{4}]+2C_{0}\lVert\mu\rVert^{2}+\lVert\mu\rVert^{4}}{C_{0}^{2}}}$ (26)
+and the previous subsection tells us that there exists a universal constant $\kappa_{Y}$ which is a form of kurtosis of our steps (which depends on our distribution) such that Batch Sharpness is smaller than GNI whenever
+$\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\leq\;\frac{2}{\eta(\kappa_{Y}+\mathcal{O}(\eta))}.$ (27)
+Proposition 4 (Stationary trace inequality).
+For every $\eta$ such that $0<\eta<2/\lambda_{\max}$ we have
+$\mathcal{A}\,\|\mu\|^{2}\;\leq\;\mathcal{B}\,\mathcal{C}$ (28)
+up to $\mathcal{O}(\eta)$ when
+$\sqrt{\mathbb{E}_{i}[\|H_{i}-H\|_{2}^{2}]}\;\leq\;\frac{2}{\eta\kappa_{Y}}+\mathcal{O}(\eta).$ (29)
+In particular, at the beginning of the training for randomly initialized weights this is the case as the expectation of the LHS is $\mathcal{O}(\log(d))$ where $d$ is the number of parameters.
+Proof.
+The theorem follows from (27) and (25). ∎
+Interpretation.
+Inequality (28) says: “the loss under random curvature cannot align too strongly with the squared prediction bias.” When either (i) the stationary iterate is unbiased ($\mu=0$) or (ii) the curvature is deterministic, the bound is tight.
+Appendix C Gradients explode above the EoSS: Proof of Theorem 1
+We now compute the update of the norm of the gradients $\mathbb{E}_{i}[\lVert Y_{i}\rVert_{2}^{2}]$ after one step in the setting in which we are at the EoSS.
+Precisely we are computing here the value of $\mathbb{E}_{t}\mathbb{E}_{i}[\lVert Y_{i}^{t+1}\rVert_{2}^{2}]$ so the average over the iterations of the update to the quantity $\mathcal{C}$ above.
+Precisely we here prove the following Proposition.
+Proposition 5.
+In the setting and notations of Appendix B.1 and B.4.
+Assume $\eta\leq 2/\lambda_{\max}(\mathcal{H})$.
+Then there exists an absolute constant $c>0$ such that when Batch Sharpness   $>2/\eta+c\eta$, then $\mathcal{C}$ increases in size exponentially and the trajectory diverges (is quadratically unstable, see Definition 2).
+Note that assumptions on $GNI$ are not necessary.
+Proof.
+In the proof we use the notations of Appendix B.
+Step 1: One step on the gradient’s second moment
+. Remind that the SGD iterate satisfy
+$\theta_{t+1}\;=\;\theta_{t}-\eta\,Y_{j_{t}}(\theta_{t}),\quad i_{t}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{D},$
+and define a fresh, independent index $j$ used only for the outer expectation in $\mathcal{C}^{t+1}$. Because $j\perp i_{t}$ we may write
+$Y_{i}(\theta_{t+1})\;=\;H_{i}\!\bigl{(}\theta_{t+1}-x_{i}\bigr{)}=Y_{i}(\theta_{t})\;-\;\eta\,H_{i}Y_{j_{t}}(\theta_{t}).$
+Squaring, expanding, and averaging over $j$ gives
+$\begin{split}\mathcal{C}^{t+1}&=\mathbb{E}_{i}\bigl{\|}Y_{i}(\theta_{t})-\eta H_{i}Y_{j_{t}}(\theta_{t})\bigr{\|}^{2}\\ &=\mathcal{C}^{t}-2\eta\,\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{i}(\theta_{t})^{\!\top}H_{i}Y_{j_{t}}(\theta_{t})\bigr{]}}_{\text{cross term}}+\eta^{2}\,\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{j_{t}}(\theta_{t})^{\!\top}H_{i}^{2}Y_{j_{t}}(\theta_{t})\bigr{]}}_{\text{variance term}}.\end{split}$ (30)
+Step 2: Decoupling the indices.
+Note that (20) in the proof of Lemma 5 establishes that
+$2\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{i}(\theta_{t})^{\!\top}H_{i}Y_{j_{t}}(\theta_{t})\bigr{]}\;=\;\mathcal{A}-\mathcal{B}-\widetilde{\Delta}.$ (31)
+This implies that we can rewrite
+$\begin{split}\mathcal{C}^{t+1}&=\mathcal{C}^{t}-\eta\,(\mathcal{A}-\mathcal{B}-\widetilde{\Delta})+\eta^{2}\,(\text{variance term}).\end{split}$ (32)
+Next note that if we are at the EoSS, then $\mathcal{A}\approx\tfrac{2}{\eta}(1+\delta)\mathcal{C}$ for some $\delta\in\mathbb{R}$. This implies that we can rewrite the term above as
+$\begin{split}\mathcal{C}^{t+1}\;&\approx\;-(1+2\delta)\,\mathcal{C}^{t}\;+\;\underbrace{\eta\mathcal{B}+\eta\widetilde{\Delta}+\eta^{2}\,(\text{variance term})}_{\text{rest}}.\end{split}$ (33)
+Let us know understand the size of the rest, the trajectory diverges if and only if:
+[…]
+Next note that by applying Jensen inequality to the term multiplied by $\eta^{2}$ we obtain that
+$\sqrt{\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{j_{t}}(\theta_{t})^{\!\top}H_{i}^{2}Y_{j_{t}}(\theta_{t})\bigr{]}}_{\text{variance term}}\cdot\underbrace{\mathbb{E}_{i}\!\bigl{[}Y_{i}(\theta_{t})^{\!\top}Y_{i}(\theta_{t})\bigr{]}}_{\mathcal{C}}}\;\geq\;\underbrace{\mathbb{E}_{i,j_{t}}\!\bigl{[}Y_{j_{t}}(\theta_{t})^{\!\top}H_{i}\cdot Y_{i}(\theta_{t})\bigr{]}}_{\mathcal{D}}.$ (35)
+Step 3: Final algebra.
+Plugging this above, we obtain that the trajectory diverges when
+[…]
+Again applying (31) we obtain that this is equivalent to
+[…]
+Since $\eta\mathcal{A}=2(1+\delta)\mathcal{C}$, then $\eta^{2}\mathcal{A}^{2}=4(1+\delta)^{2}\mathcal{C}^{2}$ to asking
+[…]
+Furthermore, equivalent to asking
+$\eta\mathcal{B}+\eta\widetilde{\Delta}-\frac{2(1+\delta)}{2}\eta\widetilde{\Delta}-\frac{2(1+\delta)}{2}\eta\mathcal{B}+\eta^{2}\frac{\mathcal{B}^{2}+\widetilde{\Delta}^{2}+2\mathcal{B}\widetilde{\Delta}}{4\mathcal{C}}\;>\;(1-\delta+\delta^{2})\,\mathcal{C}$ (39)
+or, even further simplified
+$\eta\delta(\mathcal{B}+\widetilde{\Delta})+\eta^{2}\frac{(\mathcal{B}+\widetilde{\Delta})^{2}}{4\mathcal{C}}\;>\;(1-\delta+\delta^{2})\,\mathcal{C}.$ (40)
+Here we plug in (20) again and we can rewrite this as
+$\eta\delta(\mathcal{A}-2\mathcal{D})+\eta^{2}\frac{(\mathcal{A}-2\mathcal{D})^{2}}{4\mathcal{C}}\;>\;(1-\delta+\delta^{2})\,\mathcal{C}.$ (41)
+By plugging, as before, $\eta\mathcal{A}=2(1+\delta)\mathcal{C}$ we obtain
+$2\delta(1+\delta)\mathcal{C}-2\eta\delta 2\mathcal{D}-2\eta(1+\delta)\mathcal{D}+\eta^{2}\frac{\mathcal{D}^{2}}{\mathcal{C}}\;>\;\big{(}1-\delta+\delta^{2}-(1+\delta)^{2}\big{)}\,\mathcal{C}$ (42)
+which simplifies as
+$\underbrace{2\eta(1+2\delta)\mathcal{D}}_{\mathcal{O}(\eta^{2})}-\underbrace{\eta^{2}\frac{\mathcal{D}^{2}}{\mathcal{C}}}_{\mathcal{O}(\eta^{4})}\;<\;\delta(5+2\delta)\,\underbrace{\mathcal{C}}_{\mathcal{O}_{\eta}(1)}.$ (43)
+Thus there exists a constant $c>0$, such that if $\delta>c\eta^{2}$ the trajectory diverges exponentially, if $\delta<c\eta^{2}$ the trajectory is stable.
+Appendix D When $H_{i}\equiv H$: Pure Gradient–Noise Oscillations
+Set–up.
+Assume in the Setting of Appendix B every sample shares the same curvature: $H_{i}\equiv H\succeq 0$ for all $i$.
+SGD still sees noisy gradients $g_{t}=H\Delta_{t}+\xi_{t}$ with $\mathbb{E}_{i_{t}}[\xi_{t}]=0$, $\operatorname{Var}_{i_{t}}(\xi_{t})=\Sigma_{g}$, but now there is no curvature noise.
+Batch Sharpness is capped by $\lambda_{\max}(H)$.
+For any vector $v$, $v^{\top}Hv\leq\lambda_{\max}(H)\,\|v\|^{2}$. Averaging over the mini–batch therefore yields
+$\textit{Batch Sharpness}(\theta)=\frac{\mathbb{E}_{i}[g^{\!\top}Hg]}{\mathbb{E}_{i}[\|g\|^{2}]}\;\leq\;\lambda_{\max}(H)\quad\forall\theta.$ (44)
+Hence, as long as the classical stability condition $\eta<2/\lambda_{\max}(H)$ holds,
+[…]
+Gradient–Noise Interaction still plateaus at $\tfrac{2}{\eta}$.
+The derivation in Appendix A did not use $H_{i}-H$ fluctuations, only the fixed stepsize.
+Consequently the stationary covariance satisfies $\Sigma_{x}=\eta\,\mathcal{K}^{\dagger}(\Sigma_{g})+\mathcal{O}(\eta^{2})$, and
+$\mathrm{GNI}=\frac{\mathbb{E}_{i}[g^{\!\top}Hg]}{\|\nabla L\|^{2}}=\frac{2}{\eta}\bigl{[}1+\mathcal{O}(\eta)\bigr{]}.$ (45)
+Thus SGD still wiggles with the universal Type–1 ratio $2/\eta$, even though the curvature never adapts.
+Instability threshold reverts to $\lambda_{\max}(H)$.
+Because (44) enforces Batch $\textit{Batch Sharpness}\leq\lambda_{\max}(H)$, the only way to reach the critical value $2/\eta$ is to push the stepsize past the quadratic limit:
+[…]
+Summary.
+With $H_{i}\equiv H$ the dynamics exhibits only noise–driven (Type–1) oscillations:
+  • $\mathrm{GNI}\!\to\!2/\eta$ at stationarity (same as the general case);
+  • Batch Sharpness remains below $2/\eta$, bounded by $\lambda_{\max}(H)$;
+  • the classical GD threshold $\eta=2/\lambda_{\max}(H)$ once again marks the onset of true instability.
+In short, removing curvature variability collapses the Edge-of-Stochastic Stability back to the familiar quadratic picture.
+Appendix E Mini-Batch Without Replacement
+The main text and Appendix AC treated SGD with i.i.d. sampling. Here we show how the two key quantities
+$\mathrm{GNI}\;=\;\frac{\mathbb{E}_{B}\!\bigl{[}g^{\!\top}\mathcal{H}g\bigr{]}}{\|\nabla L\|^{2}},\qquad\textit{Batch Sharpness}\;=\;\frac{\mathbb{E}_{B}[g^{\!\top}H_{B}g]}{\mathbb{E}_{B}[\|g\|^{2}]},$
+behave when each epoch is a random permutation of the $n$ samples (batch size $b$, $k:=n/b$ steps per epoch).
+We argue that without replacement sampling does not shift the EoSS in any practically relevant way.
+The only visible effect is a $1-\tfrac{b-1}{n-1}$ reduction in gradient-noise variance, which cancels in the GNI and Batch Sharpness and appears only as a second-order $\mathcal{O}(b/n)$ correction.
+The differences in effect between with and without replacements are thus about speed of convergence [49, 62] or in terms of traveling of the manifold of minima [36, 37], not in terms of EoSS.
+Noise statistics under permutation sampling.
+Fix $\theta$ and set $\bar{g}:=\nabla L(\theta)=\tfrac{1}{n}\sum_{i=1}^{n}\nabla\ell_{i}(\theta).$ For a single batch $B$ drawn without replacement
+$g_{B}:=\frac{1}{b}\sum_{i\in B}\nabla\ell_{i}(\theta)=\bar{g}+\xi_{B},\qquad\mathbb{E}[\xi_{B}\mid\theta]=0,$
+and the conditional covariance is [49, 25]
+$\operatorname{Var}[\xi_{B}\mid\theta]=\frac{n-b}{b(n-1)}\;\Sigma_{g}(\theta),\quad\Sigma_{g}(\theta):=\frac{1}{n}\sum_{i=1}^{n}\bigl{(}\nabla\ell_{i}-\bar{g}\bigr{)}\!\bigl{(}\nabla\ell_{i}-\bar{g}\bigr{)}^{\!\top}.$
+Key point: compared to i.i.d. sampling the variance is simply scaled by $\alpha_{b,n}:=\tfrac{n-b}{n-1}\in(0,1].$
+Updated Lyapunov solution.
+Denote $\Delta_{t}:=\theta_{t}-\theta^{\star}$ and keep the stepsize fixed. The error recursion for a batch update is
+$\Delta_{t+1}=(I-\eta\mathcal{H})\Delta_{t}-\eta\xi_{B_{t}},\qquad\xi_{B_{t}}\perp\!\!\!\perp\theta_{t}.$
+Exactly as in Appendix A, the discrete Lyapunov equation now reads
+$\Sigma_{x}=(I-\eta\mathcal{H})\Sigma_{x}(I-\eta\mathcal{H})+\eta^{2}\,\alpha_{b,n}\,\Sigma_{g}+\mathcal{O}(\eta^{3}),$
+so that
+$\boxed{\;\Sigma_{x}=\eta\,\alpha_{b,n}\,\mathcal{K}^{\dagger}(\Sigma_{g})+\mathcal{O}(\eta^{2}).\;}$
+Correlation between $\xi_{B_{t}}$ and $\Delta_{t}$ enters only at $\mathcal{O}\!\bigl{(}\tfrac{b}{n}\eta\bigr{)}$ and is absorbed into the $\mathcal{O}(\eta^{2})$ remainder [see 37, App. E].
+Gradient–Noise Interaction is mostly unchanged.
+Plugging the equation above into the ratio $\mathbb{E}_{B}[g^{\!\top}\mathcal{H}g]/\|\nabla L\|^{2}$ shows that the common factor $\alpha_{b,n}$ cancels:
+$\boxed{\;\mathrm{GNI}_{\text{w/o repl.}}=\frac{2}{\eta}\Bigl{[}1+\mathcal{O}(\eta)+\mathcal{O}\!\bigl{(}\tfrac{b}{n}\bigr{)}\Bigr{]}.\;}$
+Hence the plateau $2/\eta$ persists up to tiny $\mathcal{O}(b/n)$ corrections.
+Batch Sharpness and the instability edge.
+  1. 1.
+    Variable curvature ($H_{i}\!\not\equiv\!H$).
+The numerator and the denominator of BatchSharpness contain the same prefactor $\alpha_{b,n}$, so their ratio—like $\mathrm{GNI}$—is unchanged at first order:
+    $\textit{BatchSharpness}_{\text{w/o repl.}}=\frac{2}{\eta}\Bigl{[}1+\mathcal{O}(\eta)+\mathcal{O}\!\bigl{(}\tfrac{b}{n}\bigr{)}\Bigr{]}.$
+    Consequently EoSS still appears once the directional curvature meets $2/\eta$; the $\mathcal{O}(b/n)$ shift is negligible whenever $n\gg b$.
+  2. 2.
+    Constant curvature ($H_{i}\equiv H$).
+The bound $\textit{BatchSharpness}\leq\lambda_{\max}(H)$ of Appendix D is unaltered, so the classical quadratic threshold $\eta=2/\lambda_{\max}(H)$ remains the unique instability point.
+Appendix F Hardware & Compute Requirements
+All experiments were executed on a single NVIDIA A100 GPU (80 GB) with 256 GB of host RAM.
+The software stack comprises Python 3.12 and PyTorch 2.5.1 (built with the default CUDA tool-chain supplied by the wheel).
+Baseline MLP (2M parameters, Section M)
+Training for 200 epochs on the 8 k-image CIFAR-10 subset finishes in $\approx$ 5 min wall-clock while computing step sharpness every 8 updates and $\lambda_{\max}$ every 256 updates.
+Peak device memory is 14 GB during ordinary training and $\approx$ 70 GB while estimating $\lambda_{\max}$ on a 32k subset, comfortably fitting the 80 GB card.
+Algorithmic caveats.
+We rely on power iteration for $\lambda_{\max}$; while Lanczos would reduce the number of Hessian–vector products, the official PyTorch implementation remains CPU-only.
+To offset the extra memory incurred by double backward, we cache the first forward pass; batching $\lambda_{\max}$ is left to future work.
+Appendix G Comparison with Previous Empirical Work
+(a) Constant step size
+(b) Step size increased early
+(c) Step size increased late
+Figure 11: We demonstrate that the saturation of GNI doesn’t govern a sharpness-related regime of instability typical of Type-2 oscillations - and in particular, highlighting the difference in the two types of oscillations.
+When we double the learning rate after batch sharpness is at least half of $2/\eta$ threshold (so that it is beyond the new $2/\eta$ level), training exhibits a catapult surge in the loss (c).
+But if we make the same change before batch sharpness crosses that level—despite GNI already saturating—no catapult occurs.
+(b)
+(a) Constant batch size
+(b) Batch size decreased early
+(c) Batch size decreased late
+Figure 12: Similarly, reducing the batch size only triggers catapults if batch sharpness, not GNI, exceeds the threshold.
+Lee and Jang [17] introduce several quantities crucial for understanding neural network training dynamics.
+Below, we discuss the relationships among $\lambda_{\max}$, Batch Sharpness, and Interaction-Aware Sharpness (IAS, Lee and Jang [17]), emphasizing that a comprehensive theory of mini-batch dynamics should explain their distinct plateau timings and interconnected behaviors.
+We conjecture that a complete theory of stochastic gradient descent (SGD) dynamics would elucidate these metrics’ precise interrelations and their different plateau timings.
+Interaction-Aware Sharpness
+Lee and Jang [17] introduce Interaction-Aware Sharpness (IAS), denoted $\|\mathcal{H}\|_{S_{b}}$:
+$\|\mathcal{H}\|_{S_{b}}\quad:=\quad\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\nabla L_{B}(\mathbf{x})^{\top}\mathcal{H}\,\nabla L_{B}(\mathbf{x})\big{]}}{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\|\nabla L_{B}\|^{2}\bigr{]}}.$
+This quantity shares structural similarities with both Batch Sharpness (Definition 1) and the Gradient-Noise Interaction (Proposition 1), differing from the latter only in the denominator.
+The key distinction from Batch Sharpness lies in which Hessian is evaluated: IAS measures the directional curvature of the full-batch loss landscape $L$ along mini-batch gradient directions, while Batch Sharpness measures the directional curvature of mini-batch loss landscape $L_{B}$ along their corresponding gradients.
+This distinction is crucial, as mini-batch Hessians vary with batch selection while the full-batch Hessian remains fixed.
+Notably, IAS serves as a directional alternative to the maximal Hessian eigenvalue, $\lambda_{\max}$, introduced by Cohen et al.
+[1].
+IAS aligns closely with the $2/\eta$ threshold, unlike $\lambda_{\max}$, which often remains slightly above this threshold during EoS, especially at the beginning of it.
+Since IAS measures directional curvature, we have $\|\mathcal{H}\|_{S_{b}}\leq\lambda_{\max}$.
+Consequently, in the mini-batch setting, IAS stabilizes below $2/\eta$, consistent with empirical observations from Jastrzębski et al.
+[4, 5], Cohen et al.
+[1] and our Figure 2.
+Notably, when $B=n$, our Batch Sharpness coincides with IAS rather than $\lambda{\max}$, reinforcing the interpretation of Batch Sharpness as the relevant metric stabilizing at $2/\eta$ even under full-batch conditions.
+Relation to Gradient-Noise Interaction
+Another metric from Lee and Jang [17] is defined as:
+$\frac{\operatorname{tr}(HS_{b})}{\operatorname{tr}(S_{n})}=\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\nabla L_{B}(\mathbf{x})^{\top}\mathcal{H}\,\nabla L_{B}(\mathbf{x})\big{]}}{\lVert\nabla L\rVert^{2}}$
+which coincides exactly with our definition of GNI (Proposition 1).
+As detailed in Section 3 and Appendix A, the stabilization of GNI around $2/\eta$ signals the presence of oscillations, at least Type-1 oscillations.
+Lee and Jang [17] provide extensive empirical evidence demonstrating that neural networks spend much of their training within this oscillatory regime (see also Figures 11a and 12a).
+This contrasts traditional theoretical analyses (Bottou et al.
+[48], Mandt et al.
+[47]), which consider oscillations only near the manifold of minima.
+Distinguishing oscillation types
+It is crucial to note that GNI around $2/\eta$ does not inherently indicate instability.
+As clarified in Sections 3, 4 and Appendix A, not all oscillations are inherently unstable.
+Figures 4, 11b, 12b illustrate that altering hyperparameters when GNI is around $2/\eta$ typically does not trigger instability (catapult-like divergence), contrary to expectations if the system was in an EoS-like regime of instability.
+Instead, as shown in Figures 4, 11c, 12c, Batch Sharpness more reliably predicts a regime of instability.
+Additionally, Figure 13 highlights GNI’s independence from progressive sharpening, a necessary precursor to Type-2 (curvature-driven) oscillations and EoS-like instabilities, as detailed in Appendix A.5.
+Missing Progressive Sharpening.
+Extensively, both in our experiments and in the ones of Lee and Jang [17], GNI grows to $2/\eta$ in a few initial steps without ever being in subject to a phase of progressive sharpening unlike Batch Sharpness and $\lambda_{\max}$.
+The phase of growth of GNI is generally short and independent of the size, the behavior, and the phase in which Batch Sharpness and $\lambda_{\max}$ are.
+Figure 13: We construct a 32k-point "easy" CIFAR-10, where we "pull apart" all the 10 classes, so the classes become linearly separable.
+In this case, there is virtually no "learning" to be done, and therefore, there is barely any progressive sharpnening happening (as established Cohen et al.
+[1], progressive sharpening doesn’t happen if the dataset "is not complex enough").
+Yet, GNI still stabilizes at the initial level of $2/\eta$.
+More importantly, when we decrease and then increase the step size, the GNI measure restabilizes to the corresponding new thresholds, while $\lambda_{\max}$ doesn’t change.
+That means that GNI is independent of the curvature of the loss landscape and is unrelated to progressive sharpening, and thus Type-2 oscillations and EoS-like instability regimes.
+Appendix H The Geometry of the Landscape at EOSS: Level of stabilization of $\lambda_{\max}$
+So far we have determined that in the case of mini-batch training, it is the Batch Sharpness that stabilizes at $2/\eta$.
+Yet, there are both historical and generalization-related reasons to investigate what the spectrum of the full-batch Hessian is (see Section 2).
+Therefore, the question raised by the EoSS paradigm is whether we can predict the sharpness of the solution eventually found by SGD.
+H.1 Critical Batch Size
+We can characterize two regimes for the stabilization levels (see Figure 14):
+  • (i)
+    Small-batch regime ($b\leq b_{c}$): $\lambda_{\max}$ stabilizes well below the full-batch threshold $2/\eta$, signaling strong implicit regularization by SGD.
+The stabilization level rises steeply with batch size, so even modest changes in $b$ materially affect the final curvature of the loss landscape of the solution
+  • (ii)
+    Large-batch regime ($b\geq b_{c}$): the growth of $\lambda_{\max}$ with $b$ becomes much slower and the curve asymptotically approaches $2/\eta$ from below, mirroring full-batch gradient descent and reflecting weak implicit regularization.
+The critical batch size $b_{c}$ is therefore the point at which the training dynamics cross over into a full-batch-like regime.
+Works as Zhang et al.
+[63] study the following notion of critical batch size: "the point beyond which increasing batch size may result in computational efficiency degradation".
+Likewise, works focusing on generalization performance depending on the batch size [35] identify a cut-off batch sizes above which test performance degrades significantly.
+We conjecture there may be a relation between these quantities and leave a systematic investigation to future work.
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) gap to stability threshold (log–log)
+Figure 14: Baseline MLP: stabilization of $\lambda_{\max}$ as a function of batch size.
+Baseline MLP (2 hidden layers, width 512) trained on an 8k-subset of CIFAR-10 with step size 0.004 until convergence.
+(a) Final $\lambda_{\max}$ (linear axes).
+Smaller batches settle to flatter minima.
+For batch sizes below the critical batch size $b_{c}$ the level of stabilization is significantly below the $2/\eta$ level of full-batch, indicating strong implicit regularization.
+Moreover, the curve is steep, making the the final landscape sensitive to the choice of batch size.
+For larger batches ($b\!>\!b_{c}$) the slope flattens and $\lambda_{\max}$ plateaus close to $2/\eta$, so the dynamics resemble full-batch GD, implicit regularization is weak.
+(b) Log–log plot of the gap $2/\eta-\lambda_{\max}$, used to test for any power-law decay.
+The limited characterization of the level of stabilization of $\lambda_{\max}$ comes from analysis of $\lambda_{\max}^{b}$:
+$\lambda_{\max}^{b}:=\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\lambda_{\max}(\mathcal{H}(L_{\mathbf{B}}))\big{]}.$
+In particular, $\lambda_{\max}^{b}$ also stabilizes at a level $[2/\eta,4/\eta]$, see Appendix O.
+We analyze the level of stabilization $\lambda_{\max}$ through the dependence of the difference $\lambda_{\max}^{b}-\lambda_{\max}$ on the batch size (see Appendix I and J).
+H.2 Why $2/\eta-C/b^{\alpha}$ fails.
+From linear stability analyses near the manifold of minima [28, 29, 30, 32] or random matrix theory (see Appendix I) (together with the fact that we have Batch Sharpness stabilize at $2/\eta$) one would expect to have a law of the form $\lambda_{\max}\approx 2/\eta-O(1/b^{\alpha})$.
+Log-log plots of the gap $2/\eta-\lambda_{\max}$ in Figure 14b shows no robust power law (for the lack of any linear dependency), invalidating this prediction (see also Figures 15-21).
+Importantly, this doesn’t invalidate the findings of those theories, instead showcases the insufficiency of a static analysis.
+Indeed, those estimates are taken from changing the batch size statically, without making any training steps.
+In particular, linear stability analisys does accommodate virtually any law, as long as there is change in alignment between the mini-batch gradients.
+The fact that the static law does not apply means that there is a change to the alignment also happening.
+Therefore, as will be discussed further in detail, the fact that these estimates do not apply means that to give faithful description of the loss landscape at convergence one has to undertake an analysis that is path-dependent.
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) $\lambda_{\max}$ stabilization level (log–log)
+(c) gap to stability threshold (log–log)
+[…]
+H.3 Conclusion & Outlook: Why Path-Dependence Matters
+With all of the above, we arrive at a negative answer to the question posed at the start:
+There is no single, path-independent law that fixes the stabilization level of $\lambda_{\max}$ from basic hyper-parameters alone.
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) $\lambda_{\max}$ stabilization level (log–log)
+(c) gap to stability threshold (log–log)
+Figure 16: Varying dataset size alters the EOSS plateau of $\lambda_{\max}$ for a CNN.
+We use the same setup as Fig.
+15 but instead varying the number of training examples (2k, 8k, 32k).
+Larger datasets drive $\lambda_{\max}$ to lower plateaus—i.e.
+flatter minima—and push the critical batch size (the knee toward the full-batch limit $2/\eta$) to higher b, as expected from b/N scaling.
+Plateau heights also differ from the MLP results in Fig.
+15 or 14, highlighting architectural sensitivity.
+Panel order and axes mirror Fig.
+15; see that caption for sub-plot details.
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) gap to stability threshold (log–log)
+Figure 17: Level of stabilization of $\lambda_{\max}$.
+Same setup as Fig.
+14 but the initial weights are rescaled by $1/3$; see Fig.
+15 for the broader effect of initialization.
+(See Fig.
+14 for sub-plot explanations.)
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) gap to stability threshold (log–log)
+Figure 18: Level of stabilization of $\lambda_{\max}$. Identical to Fig. 14 except for a larger learning rate of $0.01$. (See Fig. 14 for sub-plot explanations.)
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) gap to stability threshold (log–log)
+Figure 19: Level of stabilization of $\lambda_{\max}$. Baseline network trained on a 32k-subset of CIFAR-10 subset with step size $0.002$. (See Fig. 14 for sub-plot explanations.)
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) gap to stability threshold (log–log)
+Figure 20: Level of stabilization of $\lambda_{\max}$. Deeper MLP (the mlp_l: $4$ hidden layers, width $512$) on the 8k-subset, learning rate $0.004$. (See Fig. 14 for sub-plot explanations.)
+(a) $\lambda_{\max}$ stabilization level (linear)
+(b) gap to stability threshold (log–log)
+Figure 21: Level of stabilization of $\lambda_{\max}$. Same deeper MLP as in Fig. 20 but trained on a 32k subset. (See Fig. 14 for sub-plot explanations.)
+Appendix I On Largest Eigenvalues of Sums of Matrices
 In this section we establish mathematically how the gap between Batch Sharpness and $\lambda_{\max}$ scales with the batch size.
@@
 In particular, the following linear algebra results collectively enhance our understanding of the stability and scaling properties of the largest eigenvalues in the context of matrix sums.
-F.1 Ordering the Largest Eigenvalues.
+I.1 Ordering the Largest Eigenvalues.
 The largest singular value of the Hessian matrix derived from single data points is positive. This observation is crucial in establishing the following well-known property of matrix eigenvalues.
@@
 $\lambda_{\max}^{1}\ <\ \lambda_{\max}^{4}\ <\ \lambda_{\max}^{16}\ <\ \lambda_{\max}^{64}\ <\ \lambda_{\max}^{256}\ <\ \ldots$ (47)
-F.2 Trends of $\lambda_{\max}$ given $b$
+I.2 Trends of $\lambda_{\max}$ given $b$
 […]
-As we discuss in Section 5.3, while the behavior we observe for very small batches is not surprising, for bigger batch sizes it is.
+As we discuss in Section J, while the behavior we observe for very small batches is not surprising, for bigger batch sizes it is.
 In computer vision tasks where there are way more parameters than datapoints, we observe that the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ decreases linearly with $1/b$ when evaluating the $\lambda_{\max}^{b}$ of any fixed model with different batch sizes, see Figure 22.
-This $1/b$ scaling is also what we expect from our mathematical analysis in Appendix F.
+This $1/b$ scaling is also what we expect from our mathematical analysis in Appendix I.
 We are indeed able to establish that for a fixed net, the gap scales as $1/b$ for small batch sizes, when the worst case $\lambda_{\max}^{b}$ is very different from $\lambda_{\max}$, and with $1/\sqrt{b}$ for big batch sizes [^13] Although big $b$ means such that the number of directions spanned in the parameter space by the vectors $\nabla_{\theta}f(\theta,x)$ are repeated multiple times, and that may be practically unrealistic with the current sizes of networks..
@@
 $\left\|\frac{1}{b}\sum_{i}M_{i}-M\right\|\quad=\quad O_{b}\left(\sqrt{\frac{\sigma^{2}\log m}{b}}+\frac{B\log m}{b}\right)$
-where $\sigma^{2}=\frac{1}{b}\max\{\|\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\|,\|\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\|\}$ is the expected second moment of the matrices and $B\geq\|M_{i}-M\|$, for all $i$, is a bound to the biggest random matrix $M_{i}$.
-F.3 Random Matrix Theory for Scaling Eigenvalues.
+where $\sigma^{2}=\frac{1}{b}\max\{\lVert\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\rVert,\lVert\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\rVert\}$ is the expected second moment of the matrices and $B\geq\lVert M_{i}-M\rVert$, for all $i$, is a bound to the biggest random matrix $M_{i}$.
+I.3 Random Matrix Theory for Scaling Eigenvalues.
 While Section 6 establishes mathematically the order of Batch Sharpness, it lacks of mathematical quantification of their magnitudes.
@@
 Lemma 7 (Matrix Bernstein Inequality).
-Let $n_{1},n_{2},b\in\mathbb{N}$, let $M_{1},M_{2},\ldots,M_{b}\in\mathbb{R}^{n_{1}\times n_{2}}$ be independent random matrices satisfying $\mathbb{E}[M_{i}]=M$ and $\|M_{i}-M\|\leq B$ for all $i$, let $v=\max\{\|\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\|,\|\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\|\}$ then for all $t>0$
+Let $n_{1},n_{2},b\in\mathbb{N}$, let $M_{1},M_{2},\ldots,M_{b}\in\mathbb{R}^{n_{1}\times n_{2}}$ be independent random matrices satisfying $\mathbb{E}[M_{i}]=M$ and $\lVert M_{i}-M\rVert\leq B$ for all $i$, let $v=\max\{\lVert\mathbb{E}[\sum_{i}M_{i}^{\top}M_{i}\rVert,\lVert\mathbb{E}[\sum_{i}M_{i}M_{i}^{\top}\rVert\}$ then for all $t>0$
 $\mathbb{P}\left({\textstyle\left\|\sum_{i}M_{i}-M\right\|\geq t}\right)\quad\leq\quad(n_{1}+n_{2})\cdot\exp\bigg{(}-\frac{b^{2}t^{2}/2}{v+Bbt/3}\bigg{)}.$
@@
 We now state a proposition that quantifies the expected spectral norm of the average of $b$ matrices $M_{i}$ based on this inequality.
-Proof of 1.
+Proof of 6.
 The Matrix Bernstein inequality bounds the probability of deviation of $|\sum_{i=1}^{b}(M_{i}-M)|$ by $t$.
@@
     Logarithmic Dimension Dependence: The $\log(n_{1}+n_{2})$ factor accounts for the high-dimensional nature of the problem.
-Appendix G Dependence of $\lambda_{\max}^{b}$-$\lambda_{\max}$ Gap on the Batch Size
+Appendix J Dependence of $\lambda_{\max}^{b}$-$\lambda_{\max}$ Gap on the Batch Size
+J.1 Highest Eigenvalue of Mini-Batch Hessian
+For a batch size $b$, denote by
+$\lambda_{\max}^{b}:=\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\lambda_{\max}(\mathcal{H}(L_{\mathbf{B}}))\big{]}.$
+{mdframed}
+[nobreak=true] Then, we establish here that:
+  • Also $\lambda_{\max}^{b}$ stabilizes.
+  • $\lambda_{\max}^{b}$ stabilizes at a level that ranges between $4/\eta$ and $2/\eta$. The level is lower for very small and very large batch sizes.
+  • $\lambda_{\max}^{b}$ is always greater than $\lambda_{\max}$.
+  • We find that the level at which $\lambda_{\max}$ stabilizes is characterized by two different regimes.
+The threshold is what we call critical batch size.
+This critical batch size depends on the complexity of the data-model.
+  • However, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ typically goes as $1/b^{\alpha}$ generally with $\alpha=0.70$.
+This is surprising, as when fixing the network, the gap goes as $1/b$ or $1/b^{1/2}$.
+See Appendices I, J
+Average $\lambda_{\max}^{b}$ for the mini-batch Hessians
+We establish that $\lambda_{\max}^{b}$–the average over the possible batches sampled of the highest eigenvalue of the mini-batch Hessian–generally stabilizes at a value just bigger than $2/\eta$.
+This happens on a wide range or models and datasets.
+Refer to Appendix O
+We observe that its stabilization levels is generally very close to $2/\eta$ for very small batch sizes, it increases until the critical batch size, then it decreases again.
 In this appendix, we further discuss the dependence of the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ from the batch size.
@@
 Figure 23: Log-log plots of (LHS) the gap between the maximum reached by the $\lambda_{\max}$ and the $\lambda_{\max}$ vs batch size; (RHS) $\lambda_{\max}^{b}$-$\lambda_{\max}$ gap vs batch size.
-G.1 The static case
+J.2 The static case
 […]
@@
 Figure 26: Log-log plot of the $\lambda_{\max}^{b}$-$\lambda_{\max}$ for fixed CNN model at convergence.
-G.2 The trained case
+J.3 The trained case
 […]
@@
 […]
-Appendix H The Hessian and the Fisher Information Matrix Overlap
+Appendix K The Hessian and the Fisher Information Matrix Overlap
 We show here empirically that at EoSS generally $\lambda_{\max}^{b}$ generally overlaps with the largest eigenvalue of the averaged mini-batch NTK and $\frac{1}{b}J_{B}^{\top}J_{B}$, which corresponds with the FIM in vision classification tasks.
@@
 Figure 30: Ranging different batch sizes, the $\lambda_{\max}^{b}$ corresponds to the largest eigenvalue of the averaged mini-batch NTK and $\frac{1}{b}J_{B}^{\top}J_{B}$, which corresponds with the FIM in vision classification tasks.
-Appendix I Exemplification Through a Simplified Model
+Appendix L Exemplification Through a Simplified Models
+A natural implicit notion of preceived curvature.
+We give here a qualitative heuristic for which Batch Sharpness is a sensible notion of curvature along the step of SGD.
+When the optimization dynamics occur on a fixed loss landscape, as in GD or noise-injected GD, the stability criterion trivially reduces to examining a single Hessian.
+Stability in such cases hinges directly on properties of this single Hessian, typically characterized by $\lambda_{\max}$, or other norms.
+In contrast, mini-batch SGD inherently samples different landscapes–and consequently different Hessians–at every iteration.
+Thus, it is non-trivial to identify which statistical properties of these sampled Hessians govern the stability dynamics.
+Contrary to initial intuition, the average Hessian does not adequately characterize stability [^14] In Appendices L.1 and L.2 we provide two toy examples where is clear that stability has to depend on higher moments or different statistics of the mini-batch Hessians and gradients and not on the full-batch (averaged) quantities..
+A natural first guess for what statistic reaches the value of $2/\eta$ would be a statistic of the maximum eigenvalue of the mini-batch Hessians, and not the maximum eigenvalue of the averaged Hessian.
+Indeed, if one has a few outlier Hessians, those may induce divergence even if the average Hessian is small.
+We also establish that is not the quantity to look at—and is generally bigger than Batch Sharpness—in Appendix O.
+While the landscape on the batch matter, the reason is that the (mis-)alignment of gradients and Hessians plays a bigger role.
+$\lambda_{\max}=2/\eta$ is the instability threshold if and only if the landscape is quadratic or the step is aligned with the eigenvector of the highest eigenvalue, here this is not the case.
+Batch Sharpness emerges naturally as the relevant measure because it explicitly captures the alignment between the gradient step direction and the curvature of the sampled mini-batch landscape at each iteration, effectively measuring the curvature perceived by SGD on each particular step.
+L.1 Stability cannot Depend on Full-Batch Quantities—Quadratics
+We show here with an example that in the mini-batch setting the stability thresholds can not depend only on full-batch Hessian or gradients, but it has to depend on the higher moments of them over batch sampling, as we can always construct a counterexample which diverges otherwise.
+Imagine on two data points we have $\mathbf{A}_{1}=\alpha\cdot I+M$ and $\mathbf{A}_{2}=\alpha\cdot I-M$, with $\alpha,\gamma>0$ and $M=\begin{pmatrix}0&0\\ 0&\gamma\end{pmatrix}$, and $\mathbf{b}_{1}=\mathbf{b}_{2}=0$.
+Here $\mathbf{A}$ is $\alpha I$.
+$\mathbf{A}_{1}$ has eigenvalues $\alpha,\alpha+\gamma$ and $\mathbf{A}_{2}$ has eigenvalues $\alpha,\alpha-\gamma$.
+Let us now look at possible full-batch stability results, as developed by Cohen et al.
+[1].
+If the right stability notion for mini-batch SGD depended on the full-batch Hessian or gradients, then it would be independent on $\gamma$, and this can not be the case.
+For instance, if a notion as the one in Equation (LABEL:eq:stab:2) would explain stability, then $\alpha$ could grow until $\alpha=2/\eta$.
+However, no matter the size of $\alpha$, if $\gamma>\max\big{\{}\alpha,\tfrac{2}{\eta}-\alpha\big{\}}$ a mini-batch iteration $x_{t+1}=(1-\eta\mathbf{A}_{i})x_{t}=$ increases the size of the second component of $x$ by a multiplicative factor bigger than 1.
+Thus a study of the stability of the system can not depend on the full-batch Hessian but has to depend on how big the oscillations due to the size of $\gamma$ are, i.e., on the higher moments of the distribution of the mini-batch Hessian.
+Note that this situation would be even more extreme if we had (as in the practice of deep learning) the top eigenvectors of the mini-batch Hessians to point in completely different directions, not just to have high variance.
+As a sanity check, any stability threshold dependent on the higher moments of the mini-batch Hessian or gradients–as Batch Sharpness or $\mathbb{E}_{B}[\lambda_{\max}(\mathcal{H}(L_{\mathbf{B}})]$–would induce $\eta_{\max}$ to depend on $\gamma$ too.
+L.2 Diagonal Linear Networks
 […]
@@
   • Smaller size: Thus setting $\lambda_{\max}$ equal to $\lambda$ means that the max between $\lambda_{1}$ and $\lambda_{2}$ is exactly $2\lambda$.
-Note that the fact that $a_{1}\cdot b_{1}=a_{2}\cdot b_{2}=1$ and Cauchy-Schwartz imply that $\lambda_{1},\lambda_{2}\geq 2$.
+Note that the fact that $a_{1}\cdot b_{1}=a_{2}\cdot b_{2}=1$ and Cauchy-Schwarz imply that $\lambda_{1},\lambda_{2}\geq 2$.
 Setting $\lambda_{\max}^{1}$ to $\lambda$ thus implies that the maximum between $\lambda_{1}$ and $\lambda_{2}$ is at most $22\lambda-$ 2, generally smaller.
@@
 This shows that in the same point of the gradient, SGD perceives the largest eigenvalue of the Hessian bigger and more relevant to the gradient then GD.
-Appendix J Illustration of EoSS in Variety of Settings: Batch Sharpness
+Appendix M Illustration of EoSS in Variety of Settings: Batch Sharpness
 In this appendix, we provide further empirical evidence that EoSS arises robustly across a variety of models, step sizes, and batch sizes.
@@
 We further confirm the EoSS regime in a five-layer CNN.
-As depicted in Figures 24 and 18, Batch Sharpness continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level.
+As depicted in Figures 41 and 32, Batch Sharpness continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level.
 Notably, as we vary the batch size, the gap between Batch Sharpness and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 6.
@@
 Figure 31: MLP: 2 hidden layers, hidden dimension 512; learning rate 0.01, 8k subset of CIFAR-10.
-Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
+Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
 Refer to caption Refer to caption
@@
 Figure 32: CNN: 5 layers (3 convolutional, 2 fully-connected), learning rate 0.03, 8k subset of CIFAR-10.
-Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
+Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
 Refer to caption Refer to caption
@@
 Figure 33: ResNet-10, learning rate 0.005, 8k subset of CIFAR-10.
-Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue dots).
-Appendix K Illustration of EoSS in Variety of Settings: $\lambda_{\max}^{b}$
+Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
+Appendix N Illustration of EoSS for the SVHN dataset
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 34: MLP: 2 hidden layers, hidden dimension 512; learning rate 0.01, 8k subset of SVHN.
+Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 35: CNN: 5 layers (3 convolutional, 2 fully-connected), learning rate 0.05, 8k subset of SVHN.
+Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 36: ResNet-10, learning rate 0.005, 8k subset of SVHN.
+Comparison between: the observed highest eigenvalue for the Hessian of the mini-batch loss (red dots), the empirical Batch Sharpness (green line), the $\lambda_{\max}$ (blue line).
+Appendix O Illustration of EoSS in Variety of Settings: $\lambda_{\max}^{b}$
 […]
@@
 To assess whether increased depth alters the phenomenon, we use a deeper MLP (MLP_L) with four hidden layers, training again on the same CIFAR-10 subset.
-Figures 22 and 23 show that $\lambda_{\max}^{b}$ exhibits the same EoSS behavior for two different step sizes, reinforcing that depth alone does not invalidate our findings.
+Figures 39 and 40 show that $\lambda_{\max}^{b}$ exhibits the same EoSS behavior for two different step sizes, reinforcing that depth alone does not invalidate our findings.
 5-Layer CNN.
 We further confirm the EoSS regime in a five-layer CNN.
-As depicted in Figures 24 and 25, $\lambda_{\max}^{b}$ continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level.
+As depicted in Figures 41 and 42, $\lambda_{\max}^{b}$ continues to plateau near the instability threshold for two distinct step sizes, while $\lambda_{\max}$ once again settles at a lower level.
 Notably, as we vary the batch size, the gap between $\lambda_{\max}^{b}$ and $\lambda_{\max}$ increases for smaller batches, mirroring the patterns described in Section 6.